### v2

Abstract
Since optimization on Riemannian manifolds relies on the chosen metric, it is appealing to know that how the performance of a Riemannian optimization method varies with different metrics and how to exquisitely construct a metric such that a method can be accelerated. To this end, we propose a general framework for optimization problems on product manifolds endowed with a preconditioned metric, and we develop the Riemannian gradient descent and Riemannian conjugate gradient methods under this metric. Generally, the metric is constructed by an operator that aims to approximate the diagonal blocks of the Riemannian Hessian of the cost function. We propose three specific approaches to design the operator: exact block diagonal preconditioning, left and right preconditioning, and Gauss–Newton type preconditioning. In addition, we tailor new preconditioned metrics and adapt the proposed Riemannian methods to the canonical correlation analysis and the truncated singular value decomposition problems, and we adopt the Gauss–Newton type preconditioning to solve the tensor ring completion problem. Numerical results among these applications verify that a delicate metric does accelerate the Riemannian optimization methods.
keywords:
Riemannian optimization; preconditioned metric; canonical correlation analysis; singular value decomposition; tensor completion
{MSCcodes}
53B21; 65K05; 65F30; 90C30
1 Introduction
We consider the optimization problems on product manifolds:
(1) $\min_{x\in\mathcal{M}}f(x),$
where $f$ is a smooth cost function and the search space $\mathcal{M}$ is a product manifold, i.e.,
$\mathcal{M}:=\mathcal{M}_{1}\times\mathcal{M}_{2}\times\cdots\times\mathcal{M}_{K},$
$\mathcal{M}_{k}$ is a smooth manifold for $k=1,2,\dots,K$ and $K$ is a positive integer. Optimization on product manifolds has a wide variety of applications, including singular value decomposition [23], joint approximate tensor diagonalization problem [30], dimensionality reduction of EEG covariance matrices [31], and canonical correlation analysis [26]. In addition, instead of working with full-size matrices or tensors, matrix and tensor decompositions—which decompose a matrix and tensor into smaller blocks—allow us to implement optimization methods on a product manifold in low-rank matrix and tensor completion [5, 15, 11, 8, 27, 12].
Related works and motivation
Recently, Riemannian optimization, designing algorithms based on the geometry of a Riemannian manifold $\mathcal{M}$, appears to be prosperous in many areas. Specifically, one can propose Riemannian optimization methods to solve problem Eq. 1, e.g., Riemannian gradient descent and Riemannian conjugate gradient methods. We refer to [2, 4] for a comprehensive overview.
Since different metrics result in different Riemannian gradients and thus distinct Riemannian methods, one is inquisitive about how the performance of a Riemannian method relies on the choice of a metric $g$. Moreover, the condition number of the Riemannian Hessian of the cost function at a local minimizer $x^{*}$, denoted by $\kappa:=\kappa_{g}(\operatorname{Hess}_{g}\!f(x^{*}))$, affects the local convergence of first-order methods in Riemannian optimization. For instance, in the Euclidean case, i.e., $\mathcal{M}=\mathbb{R}^{n}$, the asymptotic local linear convergence rates of the steepest gradient descent and the conjugate gradient methods for solving the symmetric positive-definite linear systems are $(\kappa-1)/(\kappa+1)$ and $(\sqrt{\kappa}-1)/(\sqrt{\kappa}+1)$ respectively [20, Theorems 3.3, Theorem 5.5]. In general, the asymptotic local linear convergence rate of a Riemannian gradient descent method was proved to be $1-1/\mathcal{O}(\kappa)$, see, e.g., [29, Chapter 7, Theorem 4.2][2, Theorem 4.5.6], and [4, Theorem 4.20]. Notice that an appropriate metric $g$ can lead to a smaller condition number. In view of these observations, it is natural to ask:
Can Riemannian optimization methods be accelerated
by choosing a metric “exquisitely”?
The following example presents a positive answer.
Example 1.1.
Consider the problem
$\min\ f(\mathbf{x}):=-\mathbf{b}^{\mathsf{T}}\mathbf{x},\quad\mathrm{s.\,t.}\ \mathbf{x}\in\mathcal{M}_{\mathbf{B}}:=\{\mathbf{x}\in\mathbb{R}^{n}:\ \mathbf{x}^{\mathsf{T}}\mathbf{B}\mathbf{x}=1\},$
where $\mathbf{B}\in\mathbb{R}^{n\times n}$ is symmetric positive definite and $\mathbf{b}\in\mathbb{R}^{n}$. The search space $\mathcal{M}_{\mathbf{B}}$ is an ellipsoid. The problem has a closed-form solution $\mathbf{x}^{*}=\mathbf{B}^{-1}\mathbf{b}/\|\mathbf{B}^{-1}\mathbf{b}\|_{\mathbf{B}}$ with $\|\mathbf{x}\|_{\mathbf{B}}^{2}:=\mathbf{x}^{\mathsf{T}}\mathbf{B}\mathbf{x}$. We explore the effect of a family of metrics,
$g_{\lambda,\mathbf{x}}:=\langle\xi,(\lambda\mathbf{I}_{n}+(1-\lambda)\mathbf{B})\eta\rangle\quad\text{for tangent vectors }\xi\text{ and }\eta,$
to the Riemannian gradient descent (RGD) method and the condition number of $\operatorname{Hess}_{g_{\lambda,\mathbf{x}}}\!f(\mathbf{x}^{*})$ in  Fig. 1, where $\lambda\in\mathbb{R}$ such that $\lambda\mathbf{I}_{n}+(1-\lambda)\mathbf{B}$ is positive definite. The left figure depicts the sequences generated by RGD under the Euclidean metric $g_{1,\mathbf{x}}(\xi,\eta)=\langle\xi,\eta\rangle$ and the scaled metric $g_{0,\mathbf{x}}(\xi,\eta)=\langle\xi,\mathbf{B}\eta\rangle$, and it shows that RGD under the metric $g_{0,\mathbf{x}}$ converges faster than the one under the Euclidean metric. Furthermore, the right figure confirms that the condition number varies with the metrics and $g_{0,\mathbf{x}}$ leads to the smallest condition number. The detailed computation can be found in Appendix A.
Figure 1: Left: sequences generated by the Riemannian gradient descent method under two metrics for $\mathbf{B}=\operatorname{diag}(2^{2},3^{2},1)$ and $\mathbf{b}=(1,1,1)$. Right: the condition number of $\operatorname{Hess}_{g_{\lambda,\mathbf{x}}}\!f(\mathbf{x}^{*})$ for $\lambda\in(-1/8,1]$. Blue marker: the Euclidean metric; green marker: the scaled metric.
Developing an appropriate metric to enhance the performance of Riemannian optimization methods was discussed in the existing works. For instance, Riemannian preconditioning was proposed by Mishra and Sepulchre [19] for solving equality-constrained optimization problems where the feasible set enjoys a manifold structure. The non-Euclidean metrics were derived from the Euclidean Hessian of the Lagrangian function, while the explicit construction of the Hessian can be expensive in practice. As a remedy, the block-diagonal approximation was considered to construct metrics in the matrix and tensor completion problems [18, 15, 11, 8, 12]. Specifically, in view of the block structure in tensor decompositions, the metric was developed by taking advantage of the diagonal blocks of the Hessian of the cost function, and the Riemannian optimization methods under those metrics were proved to be efficient. More recently, Shustin and Avron [25, 26] proposed a preconditioned metric for generalized Stiefel manifolds by exploiting the Riemannian Hessian of the cost function at the local minimizer.
In addition, there are other approaches that incorporate preconditioning techniques in Riemannian optimization. Boumal and Absil [5] developed a preconditioner to approximate the Riemannian Hessian in matrix completion. Kressner et al. [16] proposed preconditioned Richardson iteration and approximate Newton method to solve tensor equations by constructing a Laplacian-like operator. More recently, Tong et al. [28] introduced the scaled gradient descent method for low-rank matrix estimation. Bian et al. [3] presented a preconditioned Riemannian gradient descent algorithm for low-rank matrix recovery. Hamed and Hosseini [13] proposed a Riemannian coordinate descent method under a new Riemannian metric for low multilinear rank approximation.
Contributions
In this paper, we propose a general framework to construct a preconditioned metric on the product manifold $\mathcal{M}=\mathcal{M}_{1}\times\mathcal{M}_{2}\times\cdots\times\mathcal{M}_{K}$, which improves the performance of Riemannian optimization methods. Specifically, we consider a metric by designing a self-adjoint and positive-definite linear operator ${\bar{\mathcal{H}}}$ on the tangent bundle $\operatorname{T}\!\mathcal{E}$ such that
$g_{x}(\xi,\eta):=\langle\xi,{\bar{\mathcal{H}}}(x)[\eta]\rangle\approx\langle\xi,\operatorname{Hess}_{\mathrm{e}}\!f(x)[\eta]\rangle\quad\text{for }\xi,\eta\in\operatorname{T}_{x}\!\mathcal{M},$
[…]
Figure 2: A new metric on the product manifold $\mathcal{M}$.
We propose three types of preconditioning approaches. The first is exact block diagonal preconditioning. Instead of approximating the full Riemannian Hessian, which can be computationally unfavorable in practice, we benefit from the block structure of $\operatorname{Hess}_{\mathrm{e}}\!f(x)$ and construct a new metric by exploiting the diagonal blocks; see an illustration in Fig. 2. Specifically, given $x=(x_{1},x_{2},\dots,x_{K})\in\mathcal{M}$, we can take advantage of the diagonal blocks $H_{kk}(x)$ of $\operatorname{Hess}_{\mathrm{e}}\!f(x)$ and construct a metric
$g_{x_{k}}^{k}(\xi_{k},\eta_{k})=\langle\xi_{k},H_{kk}(x)[\eta_{k}]\rangle\quad\text{for }\xi_{k},\eta_{k}\in\operatorname{T}_{x_{k}}\!\mathcal{M}_{k}$
if $H_{kk}(x)$ is positive definite; see Section 3.1. The second type is left and right preconditioning. If at least one of the diagonal blocks is not positive definite, we construct positive-definite operators ¯${\bar{\mathcal{H}}}_{k}(x):\operatorname{T}_{x_{k}}\!\mathcal{E}_{k}\to\operatorname{T}_{x_{k}}\!\mathcal{E}_{k}$ for $k=1,2,\dots,K$ that aim to approximate the linear terms in $H_{kk}(x)$ and propose a metric
$g_{x_{k}}^{k}(\xi_{k},\eta_{k})=\langle\xi_{k},\mathbf{M}_{k,1}(x)\eta_{k}\mathbf{M}_{k,2}(x)\rangle\quad\text{for }\xi_{k},\eta_{k}\in\operatorname{T}_{x_{k}}\!\mathcal{M}_{k};$
see Section 3.2. The Riemannian metric on $\mathcal{M}$ for both diagonal types is defined by the sum of the Riemannian metric on each component, i.e., $g_{x}(\xi,\eta)=\sum_{k=1}^{K}g_{x_{k}}^{k}(\xi_{k},\eta_{k})$. Thirdly, we propose Gauss–Newton type preconditioning technique for minimization of $\frac{1}{2}\|F(x)\|_{2}^{2}$ by constructing the metric
$g_{x}(\xi,\eta)=\langle\xi,((\mathrm{D}F(x))^{*}\circ\mathrm{D}F(x))[\eta]\rangle\quad\text{for }\xi,\eta\in\operatorname{T}_{x}\!\mathcal{M};$
see Section 3.3.
By virtue of the new metric, we propose the Riemannian gradient descent and Riemannian conjugate gradient methods, and the condition number-related convergence results are presented. The preconditioned metric expands the scope of Riemannian preconditioning in [19] since it facilitates flexible choices of the operator ¯${\bar{\mathcal{H}}}(x)$. It is worth noting that exploiting more second-order information can improve the performance of Riemannian methods, but there is a trade-off between the increased cost brought by preconditioned metrics and the efficiency of preconditioned methods. The existing preconditioning methods that can be interpreted by the proposed framework are listed in Table 1.
Table 1: Existing and our works interpreted by preconditioned metrics. MC: matrix completion; TC: tensor completion; CP: canonical polyadic; TT: tensor train; TR: tensor ring.
Problem Methods Search space $\mathcal{M}$ Category
MC [18] RGD, RCG, RTR $\mathbb{R}_{*}^{m\times r}\times\mathbb{R}_{*}^{n\times r}$ Exact block diagonal
Matrix sensing [28] ScaledGD $\mathbb{R}_{*}^{m\times r}\times\mathbb{R}_{*}^{n\times r}$ Exact block diagonal
Tucker TC [19] RCG ×$\times_{k=1}^{3}\mathrm{St}(r_{k},n_{k})\times\mathbb{R}^{r_{1}\times r_{2}\times r_{3}}$ Exact block diagonal
CP TC [11] RGD, RCG ×$\times_{k=1}^{K}\mathbb{R}^{n_{k}\times r}$ Exact block diagonal
TT TC [8] RGD, RCG, RGN ×$\times_{k=1}^{K}\mathbb{R}_{*}^{r_{k-1}\times n_{k}\times r_{k}}$ Exact block diagonal
TR TC [12] RGD, RCG ×$\times_{k=1}^{K}\mathbb{R}^{n_{k}\times r_{k-1}r_{k}}$ Exact block diagonal
CCA [32, 26] RCG $\mathrm{St}_{\Sigma_{xx}}(m,d_{x})\times\mathrm{St}_{\Sigma_{yy}}(m,d_{y})$ Left and right
CCA (this work) RGD, RCG $\mathrm{St}_{\Sigma_{xx}}(m,d_{x})\times\mathrm{St}_{\Sigma_{yy}}(m,d_{y})$ Left and right
SVD (this work) RGD, RCG $\mathrm{St}(p,m)\times\mathrm{St}(p,n)$ Left and right
TR TC (this work) Gauss–Newton ×$\times_{k=1}^{K}\mathbb{R}^{n_{k}\times r_{k-1}r_{k}}$ Gauss–Newton type
Furthermore, we construct novel preconditioned metrics and apply RGD and RCG to canonical correlation analysis (CCA) and truncated singular value decomposition (SVD). We compute the condition numbers of the Riemannian Hessian of the cost function at a local minimizer for these problems. We show that the the proposed metrics indeed improve the condition numbers of Hessian and thus are able to accelerate the Riemannian methods. In addition, we propose the Gauss–Newton method for tensor ring completion. Numerical results among three applications validate the effectiveness of the proposed preconditioning framework, and these methods remain a comparable computational cost with the existing Riemannian methods.
Organization
We introduce the concepts in Riemannian optimization on product manifolds and present the convergence properties in Section 2. We propose a general framework and three specific approaches for developing a preconditioned metric on product manifolds in Section 3. We apply the proposed framework to solve the canonical correlation analysis and truncated singular value decomposition in Sections 4 and 5. The Gauss–Newton method for tensor ring completion is proposed in Section 6. Finally, we draw the conclusion in Section 7.
2 Optimization on product manifolds
In this section, we provide basic tools in Riemannian geometry on product manifolds and develop the Riemannian gradient descent and Riemannian conjugate gradient methods for optimization on product manifolds; see [2, 4] for an overview. Metric-based and condition number-related convergence properties are presented.
2.1 Riemannian optimization on product manifolds
A product manifold $\mathcal{M}$ is defined by the Cartesian product of manifolds, i.e.,
$\mathcal{M}=\mathcal{M}_{1}\times\mathcal{M}_{2}\times\cdots\times\mathcal{M}_{K}.$
Assume that $\mathcal{M}$ is embedded in a Euclidean space $\mathcal{E}=\mathcal{E}_{1}\times\mathcal{E}_{2}\times\cdots\times\mathcal{E}_{K}$, which is called the ambient space. The tangent space of $\mathcal{M}$ at $x=(x_{1},x_{2},\dots,x_{K})$ is denoted by $\operatorname{T}_{x}\!\mathcal{M}=\operatorname{T}_{x_{1}}\!\mathcal{M}_{1}\times\operatorname{T}_{x_{2}}\!\mathcal{M}_{2}\times\cdots\times\operatorname{T}_{x_{K}}\!\mathcal{M}_{K}$, and a tangent vector is denoted by $\eta=(\eta_{1},\eta_{2},\dots,\eta_{K})$. Let each manifold $\mathcal{M}_{k}$ be endowed with a Riemannian metric $g^{k}$, and hence the Riemannian metric on the product manifold $\mathcal{M}$ can be defined by
$g_{x}(\xi,\eta):=g_{x_{1}}^{1}(\xi_{1},\eta_{1})+g_{x_{2}}^{2}(\xi_{2},\eta_{2})+\cdots+g_{x_{K}}^{K}(\xi_{K},\eta_{K})$
for $\xi,\eta\in\operatorname{T}_{x}\!\mathcal{M}$, which induces a norm $\|\eta\|_{x}:=\sqrt{g_{x}(\eta,\eta)}$. Given a vector ${\bar{\eta}}=({\bar{\eta}}_{1},{\bar{\eta}}_{2},\dots,{\bar{\eta}}_{K})\in\operatorname{T}_{x}\!\mathcal{E}\simeq\mathcal{E}$, the orthogonal projection operator onto $\operatorname{T}_{x}\!\mathcal{M}$ with respect to the metric $g$ is $\Pi_{g,x}({\bar{\eta}}):=(\Pi_{g^{1},x_{1}}({\bar{\eta}}_{1}),\Pi_{g^{2},x_{2}}({\bar{\eta}}_{2}),\dots,\Pi_{g^{K},x_{K}}({\bar{\eta}}_{K}))$, where each $\Pi_{g^{k},x_{k}}$ refers to orthogonal projection operator with respect to the metric $g^{k}$ onto $\operatorname{T}_{x_{k}}\!\mathcal{M}_{k}$ for $k=1,2,\dots,K$. Let $\operatorname{T}\!\mathcal{M}:=\cup_{x\in\mathcal{M}}\operatorname{T}_{x}\!\mathcal{M}$ be the tangent bundle. A smooth mapping $\operatorname{R}:\operatorname{T}\!\mathcal{M}\to\mathcal{M}$ satisfying $\operatorname{R}_{x}(0_{x})=x$ and $\mathrm{D}\!\operatorname{R}_{x}(0_{x})=\mathrm{I}_{x}$ is called a retraction, where $0_{x}\in\operatorname{T}_{x}\!\mathcal{M}$ is the zero element, $\mathrm{D}\!\operatorname{R}_{x}(0_{x})$ is the differential of $\operatorname{R}_{x}$ at $0_{x}$, and $\mathrm{I}_{x}:\ \operatorname{T}_{x}\!\mathcal{M}\to\operatorname{T}_{x}\!\mathcal{M}$ is the identity operator on $\operatorname{T}_{x}\!\mathcal{M}$. A retraction on a product manifold $\mathcal{M}$ is defined by $\operatorname{R}_{x}(\eta):=(\operatorname{R}_{x_{1}}^{1}(\eta_{1}),\operatorname{R}_{x_{2}}^{2}(\eta_{2}),\dots,\operatorname{R}_{x_{K}}^{K}(\eta_{K}))$, where $\operatorname{R}^{k}$ is a retraction on $\mathcal{M}_{k}$. The vector transport operator is denoted by $\mathcal{T}_{y\leftarrow x}:\operatorname{T}_{x}\!\mathcal{M}\to\operatorname{T}_{y}\!\mathcal{M}$ for $x,y\in\mathcal{M}$.
Consider a smooth function $f:\mathcal{M}\to\mathbb{R}$. The Riemannian gradient under the metric $g$ is denoted by $\mathrm{grad}_{g}f(x)$, which is the unique tangent vector satisfying $g_{x}(\mathrm{grad}_{g}f(x),\eta)=\mathrm{D}f(x)[\eta]$ for all $\eta\in\operatorname{T}\!_{x}\mathcal{M}$, where $\mathrm{D}f(x)[\eta]$ refers to the differential of $f$ at $x$ along $\eta$. The Riemannian Hessian operator of $f$ at $x$ with respect to $g$ is defined by $\operatorname{Hess}_{g}\!f(x)[\eta]:=\boldsymbol{\nabla}_{\eta}\mathrm{grad}_{g}f$, where $\boldsymbol{\nabla}$ refers to Levi–Civita connection on $\mathcal{M}$. If $\mathcal{M}$ is a Riemannian submanifold of the Euclidean space $\mathcal{E}$, it follows from [4, Corollary 5.1.6] that
(1) $\operatorname{Hess}_{\mathrm{e}}\!f(x)[\eta]=\Pi_{\mathrm{e},x}(\mathrm{D}\bar{G}(x)[\eta]),$
where $\bar{G}$ is a smooth extension of $\mathrm{grad}_{\mathrm{e}}f(x)$ on a neighborhood of $\mathcal{M}$, $\mathrm{grad}_{\mathrm{e}}f(x)$ and $\operatorname{Hess}_{\mathrm{e}}\!f(x)$ are the Riemannian gradient and Riemannian Hessian of $f$ under the Euclidean metric.
By assembling the required ingredients, we present the Riemannian gradient descent and Riemannian conjugate gradient methods in Algorithms 1 and 2. We refer to [2, 21] for the global convergence of RGD and RCG.
Algorithm 1 Riemannian gradient descent method (RGD)
0:  Riemannian manifold $(\mathcal{M},g)$, initial guess $x^{(0)}\in\mathcal{M}$, $t=0$.
1:  while the stopping criteria are not satisfied do
2:     Compute $\eta^{(t)}=-\mathrm{grad}_{g}f(x^{(t)})$.
3:     Compute a stepsize $s^{(t)}$.
4:     Update $x^{(t+1)}=\operatorname{R}_{x^{(t)}}(s^{(t)}\eta^{(t)})$; $t=t+1$.
5:  end while
5:  $x^{(t)}\in\mathcal{M}$.
Algorithm 2 Riemannian conjugate gradient method (RCG)
0:  Riemannian manifold $(\mathcal{M},g)$, initial guess $x^{(0)}\in\mathcal{M}$, $t=0$, $\beta^{(0)}=0$.
1:  while the stopping criteria are not satisfied do
2:     Compute $\eta^{(t)}=-\mathrm{grad}_{g}f(x^{(t)})+\beta^{(t)}\mathcal{T}_{x^{(t)}\leftarrow x^{(t-1)}}\eta^{(t-1)}$, where $\beta^{(t)}$ is a conjugate gradient parameter.
3:     Compute a stepsize $s^{(t)}$.
4:     Update $x^{(t+1)}=\operatorname{R}_{x^{(t)}}(s^{(t)}\eta^{(t)})$; $t=t+1$.
5:  end while
5:  $x^{(t)}\in\mathcal{M}$.
Observe that the Riemannian gradients in RGD and RCG depend on the chosen metric $g$. In other words, the Riemannian methods are metric-dependent. Moreover, the computational cost in the updates of RGD and RCG varies with different metrics. Therefore, choosing an appropriate metric is apt to improve the performance of Riemannian methods.
Definition 2.1 (critical points).
Given a smooth function $f$ defined on a manifold $\mathcal{M}$ endowed with a metric $g$, a point $x^{*}\in\mathcal{M}$ is called a critical point of $f$ if $\mathrm{grad}_{g}f(x^{*})=0$.
Note that the definition of Riemannian gradient relies on the metric $g$, whereas the set of critical points of $f$ is invariant to the choice of metrics; see the following proposition.
Proposition 2.2.
Given a smooth function $f$ defined on a manifold $\mathcal{M}$. Consider two Riemannian manifolds $(\mathcal{M},g)$ and $(\mathcal{M},\tilde{g})$, it holds that
$g_{x}(\mathrm{grad}_{g}f(x),\mathrm{grad}_{\tilde{g}}f(x))\geq 0\quad\text{and}\quad\tilde{g}_{x}(\mathrm{grad}_{g}f(x),\mathrm{grad}_{\tilde{g}}f(x))\geq 0$
for $x\in\mathcal{M}$. The equality holds if and only if $x$ is a critical point. Moreover, $\mathrm{grad}_{g}f(x)=0$ if and only if $\mathrm{grad}_{\tilde{g}}f(x)=0$.
Proof 2.3.
See Appendix B.
The second-order critical point of $f$ is defined as follows.
Definition 2.4 (second-order critical points).
Given a smooth function $f$ defined on a manifold $\mathcal{M}$ endowed with a metric $g$, a critical point $x^{*}\in\mathcal{M}$ of $f$ is called a second-order critical point of $f$ if $\operatorname{Hess}_{g}\!f(x^{*})$ is positive semidefinite. Furthermore, if $\operatorname{Hess}_{g}\!f(x^{*})$ is positive definite, then $x^{*}$ is a local minimizer for Eq. 1.
Note that the set of second-order critical points is also invariant in terms of metrics; see [4, Proposition 6.3]. Specifically, if $x^{*}$ is a second-order critical point of $f$, it holds that $\operatorname{Hess}_{g}\!f(x^{*})$ is positive semidefinite if and only if $\operatorname{Hess}_{\tilde{g}}\!f(x^{*})$ is positive semidefinite for different metrics $g$ and $\tilde{g}$.
2.2 Local convergence properties
We present the local convergence properties of the Riemannian gradient descent method in terms of condition numbers. Specifically, the Armijo backtracking line search is applied to computing the stepsize in Algorithm 1.
Definition 2.5 (Armijo backtracking line search).
Given a smooth function $f$ defined on a manifold $\mathcal{M}$ endowed with a metric $g$, a point $x\in\mathcal{M}$, a vector $\eta\in\operatorname{T}_{x}\!\mathcal{M}$, an initial stepsize $s_{0}>0$, and constants $\rho,a\in(0,1)$. The Armijo backtracking line search aims to find the smallest non-negative integer $\ell$, such that for $s=\rho^{\ell}s_{0}$, the condition
(2) $f(x)-f(\operatorname{R}_{x}(s\eta))\geq-sag_{x}(\mathrm{grad}_{g}f(x),\eta)$
holds.
In Riemannian optimization, the condition number of the Riemannian Hessian at the local minimizer $x^{*}$ is crucial to the local converge rate of Riemannian methods; see, e.g., [2, Theorem 4.5.6] and [4, Theorem 4.20]. The condition number of the Riemannian Hessian $\operatorname{Hess}_{g}\!f(x^{*})$ is defined by
(3) $\kappa_{g}(\operatorname{Hess}_{g}\!f(x^{*})):=\frac{\lambda_{\max}(\operatorname{Hess}_{g}\!f(x^{*}))}{\lambda_{\min}(\operatorname{Hess}_{g}\!f(x^{*}))}=\frac{\sup_{\eta\in\operatorname{T}_{x^{*}}\!\mathcal{M}}q_{x^{*}}(\eta)}{\inf_{\eta\in\operatorname{T}_{x^{*}}\!\mathcal{M}}q_{x^{*}}(\eta)},$
where $\lambda_{\min}(\operatorname{Hess}_{g}\!f(x^{*}))$ and $\lambda_{\max}(\operatorname{Hess}_{g}\!f(x^{*}))$ denote the smallest and largest eigenvalue of $\operatorname{Hess}_{g}\!f(x^{*})$, and
(4) $q_{x^{*}}(\eta):=\frac{g_{x^{*}}(\eta,\operatorname{Hess}_{g}\!f(x^{*})[\eta])}{g_{x^{*}}(\eta,\eta)}$
refers to the Rayleigh quotient, which depends on the metric $g$. Then, the local convergence rate of RGD with Armijo backtracking line search Eq. 2 for optimization on product manifolds can be proved by following [2, Theorem 4.5.6].
Theorem 2.6.
Let $\{x^{(t)}\}_{t=0}^{\infty}$ be an infinite sequence generated by Algorithm 1 with backtracking line search Eq. 2 converging to a local minimizer $x^{*}$. There exists $T>0$, such that for all $t>T$, it holds that
$\frac{f(x^{(t)})-f(x^{*})}{f(x^{(t-1)})-f(x^{*})}\leq 1-\min\{2as_{0}\lambda_{\min}(\operatorname{Hess}_{g}\!f(x^{*})),\frac{4a(1-a)\rho}{\kappa_{g}(\operatorname{Hess}_{g}\!f(x^{*}))}\}.$
It is worth noting that different metrics can lead to different $\lambda_{\min}(\operatorname{Hess}_{g}\!f(x^{*}))$ and $\kappa_{g}(\operatorname{Hess}_{g}\!f(x^{*}))$ by Eq. 3, which affect the local convergence rate. More precisely, a lower condition number indicates a faster convergence in RGD.
3 Developing preconditioned metric on product manifolds
We first propose a general framework for developing a preconditioned metric on a product manifold $\mathcal{M}$ by constructing an operator ¯${\bar{\mathcal{H}}}(x)$ that aims to approximate the diagonal blocks of Riemannian Hessian. Next, we develop three specific approaches to construct the operator ¯${\bar{\mathcal{H}}}(x)$.
Generally, we first propose to endow $\mathcal{E}$ with a metric ${\bar{g}}$ by designing a self-adjoint and positive-definite linear operator ${\bar{\mathcal{H}}}$ on the tangent bundle $\operatorname{T}\!\mathcal{E}$ such that
(1) ${\bar{g}}_{x}(\xi,\eta)=\langle\xi,{\bar{\mathcal{H}}}(x)[\eta]\rangle\approx\langle\xi,\operatorname{Hess}_{\mathrm{e}}\!f(x)[\eta]\rangle\quad\text{for }\xi,\eta\in\operatorname{T}_{x}\!\mathcal{M}.$
The rationale behind Eq. 1 is that the metric ${\bar{g}}$, aiming to approximate the second-order information of $\operatorname{Hess}_{\mathrm{e}}\!f(x)$ on $\operatorname{T}_{x}\!\mathcal{M}$, is able to facilitate the computation of the Riemannian gradient. Then, the Riemannian metric $g$ on $\mathcal{M}$ is inherited from $\mathcal{E}$ in view of the Riemannian submanifold. Since $\mathcal{M}=\mathcal{M}_{1}\times\mathcal{M}_{2}\times\cdots\times\mathcal{M}_{K}$ is a product manifold, it follows from [4, Example 5.19] that the Riemannian Hessian of $f$ at $x=(x_{1},x_{2},\dots,x_{K})$ along $\eta=(\eta_{1},\eta_{2},\dots,\eta_{K})$ enjoys a block structure, i.e.,
(2) $\operatorname{Hess}_{\mathrm{e}}\!f(x)[\eta]=($ ${H_{11}(x)}[\eta_{1}]~{}+H_{12}(x)[\eta_{2}]~{}+\cdots+H_{1K}(x)[\eta_{K}],$
$H_{21}(x)[\eta_{1}]~{}+H_{22}(x)[\eta_{2}]~{}+\cdots+H_{2K}(x)[\eta_{K}],$
[…]
$H_{K1}(x)[\eta_{1}]+H_{K2}(x)[\eta_{2}]+\cdots+{H_{KK}(x)}[\eta_{K}]),$
where $H_{ij}(x)[\eta_{j}]:=\operatorname{Hess}_{\mathrm{e}}\!f(x_{1},\dots,x_{i-1},\cdot,x_{i+1},\dots,x_{K})(x_{i})[\eta_{i}]$ if $i=j$, $H_{ij}(x)[\eta_{j}]:=\mathrm{D}G_{i}(x_{1},\dots,x_{j-1},\cdot,x_{j+1},\dots,x_{K})(x_{j})[\eta_{j}]$ if $i\neq j$ and $f(x_{1},\dots,x_{i-1},\cdot,x_{i+1},\dots,x_{K})$ denotes the function that $f$ is restricted on $\mathcal{M}_{i}$. The operator $G_{i}:\mathcal{M}\to\operatorname{T}_{x_{i}}\!\mathcal{M}_{i}$, $G_{i}(x):=\mathrm{grad}_{\mathrm{e}}f(x_{1},\dots,x_{i-1},\cdot,x_{i+1},\dots,x_{K})(x_{i})$, gives the Riemannian gradient of the above function. The restriction of $G_{i}$ on $\mathcal{M}_{j}$ is $G_{i}(x_{1},\dots,x_{j-1},\cdot,x_{j+1},\dots,x_{K}):\mathcal{M}_{j}\to\operatorname{T}_{x_{i}}\!\mathcal{M}_{i}$. In the light of the Riemannian submanifold, the Riemannian gradient of $f$ at $x\in\mathcal{M}$ with respect to $g$ can be computed by following [2, (3.37)].
Proposition 3.1.
Let $(\mathcal{M},g)$ be a Riemannian submanifold of $(\mathcal{E},{\bar{g}})$. Given a function $f:\mathcal{M}\to\mathbb{R}$ and its smooth extension ${\bar{f}}:\mathcal{E}\to\mathbb{R}$, the Riemannian gradient of $f$ at $x\in\mathcal{M}$ can be computed by
$\mathrm{grad}_{g}f(x)=\Pi_{g,x}({\bar{\mathcal{H}}}(x)^{-1}[\nabla{\bar{f}}(x)]),$
where $\Pi_{g,x}:\operatorname{T}_{x}\!\mathcal{E}\simeq\mathcal{E}\to\operatorname{T}_{x}\!\mathcal{M}$ is the orthogonal projection operator with respect to the metric $g$ onto $\operatorname{T}_{x}\!\mathcal{M}$, and $\nabla{\bar{f}}(x)$ denotes the Euclidean gradient of ${\bar{f}}$.
In view of Eqs. 1 and 3.1, the operator ¯${\bar{\mathcal{H}}}(x)$ has a preconditioning effect. Hence, we refer to the metric $g$ as a preconditioned metric, and ${\bar{\mathcal{H}}}$ is the preconditioner. The methodology of using a preconditioned metric can be deemed a general framework to accelerate the Riemannian methods. Subsequently, we design specific approaches for constructing the operator ¯${\bar{\mathcal{H}}}(x)$.
3.1 Exact block diagonal preconditioning
Instead of acquiring the full Riemannian Hessian $\operatorname{Hess}_{\mathrm{e}}\!f(x)$, which involves the computation of all blocks $H_{ij}(x)$ in Eq. 2, we develop a metric in a more economical manner by using the diagonal blocks $H_{11},H_{22},\dots,H_{KK}$, as a trade-off between the efficiency and the computational cost.
Recall that a Riemannian metric on the product manifold $\mathcal{M}$ is defined by the sum of the metrics on each component, i.e., $g_{x}(\xi,\eta)=\sum_{k=1}^{K}g_{x_{k}}^{k}(\xi_{k},\eta_{k})$ for $\xi=(\xi_{1},\xi_{2},\dots,\xi_{K}),\eta=(\eta_{1},\eta_{2},\dots,\eta_{K})\in\operatorname{T}_{x}\!\mathcal{M}$. Consider the smooth extension $\bar{H}_{kk}(x):\operatorname{T}_{x_{k}}\!\mathcal{E}_{k}\to\operatorname{T}_{x_{k}}\!\mathcal{E}_{k}$ of the diagonal block $H_{kk}$ for $k=1,\dots,K$. If $\bar{H}_{11},\bar{H}_{22},\dots,\bar{H}_{KK}$ are positive definite on the ambient space $\mathcal{E}$, one can immediately adopt these blocks to construct the operator ${\bar{\mathcal{H}}}$ by
(3) ¯${\bar{\mathcal{H}}}(x)[\eta]=({\bar{\mathcal{H}}}_{1}(x)[\eta_{1}],\dots,{\bar{\mathcal{H}}}_{K}(x)[\eta_{K}])=(\bar{H}_{11}(x)[\eta_{1}],\dots,\bar{H}_{KK}(x)[\eta_{K}]).$
Hence, the metric
$g_{x}(\xi,\eta)$ $=g_{x_{1}}^{1}(\xi_{1},\eta_{1})+g_{x_{2}}^{2}(\xi_{2},\eta_{2})+\cdots+g_{x_{K}}^{K}(\xi_{K},\eta_{K})$
$=\langle\xi_{1},\bar{H}_{11}(x)[\eta_{1}]\rangle+\langle\xi_{2},\bar{H}_{22}(x)[\eta_{2}]\rangle+\cdots+\langle\xi_{K},\bar{H}_{KK}(x)[\eta_{K}]\rangle$
is a well-defined Riemannian metric on $\mathcal{M}$, which leads to exact block diagonal preconditioning. Note that the proposed Exact block diagonal preconditioning is not applicable when at least one of the blocks $\bar{H}_{11},\bar{H}_{22},\dots,\bar{H}_{KK}$ is not positive definite. In practice, one can consider a regularization term $\delta_{k}\mathcal{I}_{k}(x)$ with the identity operator $\mathcal{I}_{k}(x):\operatorname{T}_{x_{k}}\!\mathcal{E}_{k}\to\operatorname{T}_{x_{k}}\!\mathcal{E}_{k}$ and some $\delta_{k}>0$ to ensure that the operator $\bar{H}_{kk}(x)+\delta_{k}\mathcal{I}_{k}(x)$ is positive definite.
In contrast with the Riemannian Hessian Eq. 2, which contains cross terms among $\eta_{1},\eta_{2},\dots,\eta_{K}$, the operator ¯${\bar{\mathcal{H}}}(x)$ in Eq. 3 enjoys a block structure. Therefore, the Riemannian gradient of a smooth function $f:\mathcal{M}\to\mathbb{R}$ at $x\in\mathcal{M}$ can be computed on each block in view of Proposition 3.1.
Proposition 3.2.
Let $\mathcal{M}=\mathcal{M}_{1}\times\mathcal{M}_{2}\times\cdots\times\mathcal{M}_{K}$ be a product manifold endowed with the metric $g$. Given a function $f:\mathcal{M}\to\mathbb{R}$ and its smooth extension ${\bar{f}}:\mathcal{E}\to\mathbb{R}$, the Riemannian gradient of $f$ at $x$ is
(4) $\mathrm{grad}_{g}f(x)=\big{(}\Pi_{g^{1},x_{1}}({\bar{\mathcal{H}}}_{1}(x)^{-1}[\partial_{1}{\bar{f}}(x)]),\dots,\Pi_{g^{K},x_{K}}({\bar{\mathcal{H}}}_{K}(x)^{-1}[\partial_{K}{\bar{f}}(x)])\big{)},$
where $\Pi_{g^{k},x_{k}}$ is the orthogonal projection operator with respect to the metric $g^{k}$ onto $\operatorname{T}_{x_{k}}\!\mathcal{M}_{k}$ for $k=1,2,\dots,K$ and $\partial_{k}{\bar{f}}(x)$ is the partial derivative of $f$ with respect to $x_{k}$.
It is worth noting that developing an appropriate metric by exploiting the diagonal blocks is closely related to the block-Jacobi preconditioning [10] in numerical linear algebra. Specifically, given a symmetric positive definite matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$, the goal of block-Jacobi preconditioning is to construct an invertible block-diagonal matrix
$\mathbf{D}=\begin{bmatrix}\mathbf{D}_{11}&&&\\ &\mathbf{D}_{22}&&\\ &&\ddots&\\ &&&\mathbf{D}_{KK}\\ \end{bmatrix}\in\mathbb{R}^{n\times n}\ \text{with}\ \mathbf{D}_{kk}\in\mathbb{R}^{n_{k}\times n_{k}},\ k=1,2,\dots,K$
such that $\kappa_{2}(\mathbf{D}\mathbf{A}\mathbf{D}^{\mathsf{T}}):=\lambda_{\max}(\mathbf{D}\mathbf{A}\mathbf{D}^{\mathsf{T}})/\lambda_{\min}(\mathbf{D}\mathbf{A}\mathbf{D}^{\mathsf{T}})$ is reduced, where $n_{1}+n_{2}+\cdots+n_{K}=n$. Alternatively, consider the minimization problem of a quadratic function $\min_{\mathbf{x}\in\mathbb{R}^{n}}f(\mathbf{x}):=\frac{1}{2}\mathbf{x}^{\mathsf{T}}\mathbf{A}\mathbf{x}$. We can construct a preconditioned metric on the product manifold $\mathbb{R}^{n}=\mathbb{R}^{n_{1}}\times\mathbb{R}^{n_{2}}\times\cdots\times\mathbb{R}^{n_{K}}$ by
$g_{\mathbf{x}}(\xi,\eta):=\sum_{k=1}^{K}\langle\xi_{k},(\mathbf{D}_{kk}^{\mathsf{T}}\mathbf{D}_{kk})^{-1}\eta_{k}\rangle=\langle\xi,(\mathbf{D}^{\mathsf{T}}\mathbf{D})^{-1}\eta\rangle.$
Given $\mathbf{x}\in\mathbb{R}^{n}$, it follows from Proposition 3.2 and the definition of Riemannian Hessian that $\mathrm{grad}_{g}f(\mathbf{x})=(\mathbf{D}^{\mathsf{T}}\mathbf{D})\mathbf{A}\mathbf{x}$ and $\operatorname{Hess}_{g}\!f(\mathbf{x})=(\mathbf{D}^{\mathsf{T}}\mathbf{D})\mathbf{A}$. Therefore, the Rayleigh quotient is given by
[…]
for $\eta\in\operatorname{T}_{\mathbf{x}}\!\mathbb{R}^{n}\simeq\mathbb{R}^{n}$. It follows from Eq. 3 that
$\kappa_{g}(\operatorname{Hess}_{g}\!f(\mathbf{x}))=\frac{\sup_{\eta\in\operatorname{T}_{\mathbf{x}}\!\mathcal{M}}q_{\mathbf{x}}(\eta)}{\inf_{\eta\in\operatorname{T}_{\mathbf{x}}\!\mathcal{M}}q_{\mathbf{x}}(\eta)}=\frac{\lambda_{\max}(\mathbf{D}\mathbf{A}\mathbf{D}^{\mathsf{T}})}{\lambda_{\min}(\mathbf{D}\mathbf{A}\mathbf{D}^{\mathsf{T}})}=\kappa_{2}(\mathbf{D}\mathbf{A}\mathbf{D}^{\mathsf{T}}).$
As a result, the block-Jacobi preconditioning that aims to reduce $\kappa_{2}(\mathbf{D}\mathbf{A}\mathbf{D}^{\mathsf{T}})$ is equivalent to selecting a specific preconditioned metric on $\mathbb{R}^{n}$ to reduce the condition number of the Riemannian Hessian of $f$, i.e., $\kappa_{g}(\operatorname{Hess}_{g}\!f(\mathbf{x}))$. Additionally, the preconditioning in matrix and tensor completion can be interpreted as the proposed exact block diagonal preconditioning; see Section 6 for details.
3.2 Left and right preconditioning
Generally, in block diagonal preconditioning, $\bar{H}_{11},\bar{H}_{22},\dots,\bar{H}_{KK}$ are not positive definite on the ambient space $\mathcal{E}$. Therefore, we seek an appropriate approximation of these terms.
Specifically, we assume that $\mathcal{E}$ is a product space of matrices and ${H}_{kk}(x)[\eta_{k}]$ contains a linear term $\bar{\mathbf{M}}_{k,1}(x)\eta_{k}\bar{\mathbf{M}}_{k,2}(x)$, where $\bar{\mathbf{M}}_{k,1}(x)$ and $\bar{\mathbf{M}}_{k,2}(x)$ are square matrices for fixed $x\in\mathcal{E}$. In view of the Riemannian Hessian Eq. 2, we approximate the diagonal block ${H}_{kk}(x)$ via $\bar{\mathbf{M}}_{k,1}(x)\eta_{k}\bar{\mathbf{M}}_{k,2}(x)$ by constructing an operator ¯${\bar{\mathcal{H}}}_{k}(x):\operatorname{T}_{x_{k}}\!\mathcal{E}_{k}\to\operatorname{T}_{x_{k}}\!\mathcal{E}_{k}$ such that
$\langle\xi_{k},{\bar{\mathcal{H}}}_{k}(x)[\eta_{k}]\rangle=\langle\xi_{k},\mathbf{M}_{k,1}(x)\eta_{k}\mathbf{M}_{k,2}(x)\rangle\approx\langle\xi_{k},{H}_{kk}(x)[\eta_{k}]\rangle\quad\text{for }\xi_{k},\eta_{k}\in\operatorname{T}_{x_{k}}\!\mathcal{M}_{k},$
where $\mathbf{M}_{k,j}(x)=(\operatorname{sym}(\bar{\mathbf{M}}_{k,j}(x))^{2}+\delta\mathbf{I})^{1/2}$ for $j=1,2$, $\operatorname{sym}(\mathbf{A}):=(\mathbf{A}+\mathbf{A}^{\mathsf{T}})/2$ and $\delta>0$ to ensure the positive definiteness. Subsequently, a left and right preconditioned metric on $\operatorname{T}_{x}\!\mathcal{M}$ is given by
$g_{x}(\xi,\eta)$ $=g_{x_{1}}^{1}(\xi_{1},\eta_{1})+\cdots+g_{x_{K}}^{K}(\xi_{K},\eta_{K})$
(5) $=\langle\xi_{1},\mathbf{M}_{1,1}(x)\eta_{1}\mathbf{M}_{1,2}(x)\rangle+\cdots+\langle\xi_{K},\mathbf{M}_{K,1}(x)\eta_{K}\mathbf{M}_{K,2}(x)\rangle.$
Note that $\mathbf{M}_{k,j}(x)$ are smooth and positive definite for all $x\in\mathcal{E}$ and thus Eq. 5 is a well-defined Riemannian metric. The corresponding Riemannian gradient can be also computed by following Proposition 3.2 since the operator ${\bar{\mathcal{H}}}$ is defined on each block.
We adopt the proposed left and right preconditioning to accelerate the Riemannian methods for canonical correlation analysis (CCA); see Section 4 for details. In practice, we can consider only the left or right preconditioning to save the computational cost. Section 5 presents how we develop right preconditioning for truncated singular value decomposition (SVD). It is worth noting that if the operators $\mathbf{M}_{k,j}$ in Eq. 5 are not chosen appropriately, Riemannian methods under the metric Eq. 5 can even perform worse than those under the Euclidean metric; see Section 4.4. Nevertheless, the proposed metrics tailored for CCA and SVD improve the condition number of $\operatorname{Hess}\!f(x)$, thereby accelerating the Riemannian methods indeed; see Proposition 4.8.
3.3 Gauss–Newton type preconditioning
If the cost function in Eq. 1 satisfies that $f(x):=\frac{1}{2}\|F(x)\|_{\mathrm{F}}^{2}$ for some smooth function $F:\mathcal{M}\to\mathbb{R}^{n}$ with injective $\mathrm{D}F(x)$, one can consider the operator ¯${\bar{\mathcal{H}}}(x)=(\mathrm{D}F(x))^{*}\circ\mathrm{D}F(x)$ to approximate $\operatorname{Hess}_{\mathrm{e}}\!f(x)$, and construct the preconditioned metric as follows,
(6) $g_{x}(\xi,\eta)=\langle\xi,{\bar{\mathcal{H}}}(x)[\eta]\rangle=\langle\xi,((\mathrm{D}F(x))^{*}\circ\mathrm{D}F(x))[\eta]\rangle\approx\langle\xi,\operatorname{Hess}_{\mathrm{e}}\!f(x)[\eta]\rangle,$
where $(\mathrm{D}F(x))^{*}$ is the adjoint operator of $\mathrm{D}F(x)$. The Gauss–Newton type preconditioning is no longer a block diagonal preconditioning since ¯${\bar{\mathcal{H}}}(x)$ contains cross terms among $\eta_{1},\eta_{2},\dots,\eta_{K}$. As a result, the Riemannian gradient directly follows from Proposition 3.1.
Note that the Riemannian gradient descent method with the metric $g$ is exactly the Riemannian Gauss–Newton method [2, §8.4.1], where the search direction $\eta^{(t)}\in\operatorname{T}_{x^{(t)}}\!\mathcal{M}$ at $x^{(t)}\in\mathcal{M}$ is computed by the following Gauss–Newton equation
$\langle\mathrm{D}F(x^{(t)})[\xi],\mathrm{D}F(x^{(t)})[\eta^{(t)}]\rangle+\langle\mathrm{D}F(x^{(t)})[\xi],F(x^{(t)})\rangle=0\quad\text{for }\xi\in\operatorname{T}_{x^{(t)}}\!\mathcal{M},$
or $((\mathrm{D}F(x^{(t)}))^{*}\circ\mathrm{D}F(x^{(t)}))[\eta^{(t)}]=-(\mathrm{D}F(x^{(t)}))^{*}[F(x^{(t)})]$. It follows from the injectivity of $\mathrm{D}F(x^{(t)})$ that
$\eta^{(t)}=-((\mathrm{D}F(x^{(t)}))^{*}\circ\mathrm{D}F(x^{(t)}))^{-1}[(\mathrm{D}F(x^{(t)}))^{*}[F(x^{(t)})]],$
which is also the solution of the following least-squares problem
(7) $\min_{\eta\in\operatorname{T}_{x^{(t)}}\!\mathcal{M}}\frac{1}{2}\langle\mathrm{D}F(x^{(t)})[\eta],\mathrm{D}F(x^{(t)})[\eta]\rangle+\langle\mathrm{D}F(x^{(t)})[\eta],F(x^{(t)})\rangle.$
Since $\langle\mathrm{D}F(x^{(t)})[\eta^{(t)}],F(x^{(t)})\rangle=\mathrm{D}f(x^{(t)})[\eta^{(t)}]=\mathrm{D}{\bar{f}}(x^{(t)})[\eta^{(t)}]=\langle\nabla{\bar{f}}(x^{(t)}),\eta^{(t)}\rangle$, where ${\bar{f}}:\mathcal{E}\to\mathbb{R}$ is any smooth extension of $f$Eq. 7 is equivalent to
(8) $\min_{\eta\in\operatorname{T}_{x^{(t)}}\!\mathcal{M}}\frac{1}{2}\langle{\bar{\mathcal{H}}}(x^{(t)})[\eta],\eta\rangle+\langle\nabla{\bar{f}}(x^{(t)}),\eta\rangle.$
It follows from [19] that the solution of Eq. 8 is $\eta^{(t)}=-\mathrm{grad}_{g}f(x^{(t)})$. In other words, the Riemannian Gauss–Newton method can be interpreted by the Riemannian gradient descent method with the metric $g$. Therefore, we refer to the proposed framework as the Gauss–Newton type preconditioning, which can be adopted to tensor completion; see Section 6 for details.
Remark 3.3.
Let the operator ¯${\bar{\mathcal{H}}}(x)$ be the Riemannian Hessian $\operatorname{Hess}_{\mathrm{e}}\!f(x)$ of $f$ at $x\in\mathcal{M}$ under the Euclidean metric. If $\operatorname{Hess}_{\mathrm{e}}\!f(x)$ is positive definite, the metric $g_{x}(\xi,\eta)=\langle\xi,\operatorname{Hess}_{\mathrm{e}}\!f(x)[\eta]\rangle$ is referred to Hessian metric in [24]. The negative Riemannian gradient $-\mathrm{grad}_{g}f(x)$ echoes the Riemannian Newton direction under the Euclidean metric; see [1, Proposition 4.1] and [19, Proposition 2.1]. Note that the proposed Gauss–Newton type preconditioning is different from the Hessian metric since Eq. 6 adopts partial information of the Riemannian Hessian.
4 Application to canonical correlation analysis
In this section, we apply the proposed framework to solve the canonical correlation analysis (CCA) problem. A new left and right preconditioned metric is proposed. Then, we prove that the the proposed metric improves the condition number of the Riemannian Hessian. Numerical experiments verify that the proposed metric accelerates the Riemannian methods.
Consider two data matrices $\mathbf{X}\in\mathbb{R}^{n\times d_{x}}$ and $\mathbf{Y}\in\mathbb{R}^{n\times d_{y}}$ with $n$ samples and $d_{x}$, $d_{y}$ variables respectively. The goal of CCA is to choose $m$ weights $\mathbf{u}_{1},\dots,\mathbf{u}_{m}\in\mathbb{R}^{d_{x}}$ and $\mathbf{v}_{1},\dots,\mathbf{v}_{m}\in\mathbb{R}^{d_{y}}$ such that the data matrices $\mathbf{X}\mathbf{U}$ and $\mathbf{Y}\mathbf{V}$ have the highest correlation, where $\mathbf{U}=[\mathbf{u}_{1},\dots,\mathbf{u}_{m}]$ and $\mathbf{V}=[\mathbf{v}_{1},\dots,\mathbf{v}_{m}]$. CCA can be written as an optimization problem on the product manifold of two generalized Stiefel manifolds, i.e,
(1) $\min_{\mathbf{U},\mathbf{V}}\ f(\mathbf{U},\mathbf{V}):=-\mathrm{tr}(\mathbf{U}^{\mathsf{T}}\Sigma_{xy}\mathbf{V}\mathbf{N}),\ \mathrm{s.\,t.}\ (\mathbf{U},\mathbf{V})\in\mathcal{M}=\mathcal{M}_{1}\times\mathcal{M}_{2},$
where $\Sigma_{xx}:=\mathbf{X}^{\mathsf{T}}\mathbf{X}+\lambda_{x}\mathbf{I}_{d_{x}}$, $\Sigma_{yy}:=\mathbf{Y}^{\mathsf{T}}\mathbf{Y}+\lambda_{y}\mathbf{I}_{d_{y}}$, $\lambda_{x},\lambda_{y}\geq 0$ are regularization parameters, $\Sigma_{xy}:=\mathbf{X}^{\mathsf{T}}\mathbf{Y}$, $\mathcal{M}_{1}:=\mathrm{St}_{\Sigma_{xx}}(m,d_{x})=\{\mathbf{U}\in\mathbb{R}^{d_{x}\times m}:\mathbf{U}^{\mathsf{T}}\Sigma_{xx}\mathbf{U}=\mathbf{I}_{m}\}$ and $\mathcal{M}_{2}:=\mathrm{St}_{\Sigma_{yy}}(m,d_{y})$ refer to the generalized Stiefel manifolds, and $\mathbf{N}:=\operatorname{diag}(\mu_{1},\mu_{2},\dots,\mu_{m})$ satisfies $\mu_{1}>\mu_{2}>\cdots>\mu_{m}>0$. The cost function $f$ in Eq. 1 is also known as the von Neumann cost function [33]. The problem Eq. 1 has a closed-form solution
(2) $(\mathbf{U}^{*},\mathbf{V}^{*})=(\Sigma_{xx}^{-1/2}\bar{\mathbf{U}},\Sigma_{yy}^{-1/2}\bar{\mathbf{V}}),$
where $\bar{\mathbf{U}}:=[\bar{\mathbf{u}}_{1},\dots,\bar{\mathbf{u}}_{m}]$ and $\bar{\mathbf{V}}:=[\bar{\mathbf{v}}_{1},\dots,\bar{\mathbf{v}}_{m}]$ are the $m$ leading left and right singular vectors of the matrix $\Sigma_{xx}^{-1/2}\Sigma_{xy}\Sigma_{yy}^{-1/2}$ respectively. The $m$ largest singular values $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{m}>0$ of $\Sigma_{xx}^{-1/2}\Sigma_{xy}\Sigma_{yy}^{-1/2}$ are referred to as the canonical correlations. We intend to propose preconditioned metrics on $\mathcal{M}$ and adapt the Riemannian methods to solve Eq. 1.
4.1 Left preconditioner
Shustin and Avron [26, §4.2] proposed to endow $\mathcal{M}$ with the following metric
(3) $g_{(\mathbf{U},\mathbf{V})}(\xi,\eta):=\langle\xi_{1},\Sigma_{xx}\eta_{1}\rangle+\langle\xi_{2},\Sigma_{yy}\eta_{2}\rangle\quad\text{for }\xi,\eta\in\operatorname{T}_{(\mathbf{U},\mathbf{V})}\!\mathcal{M},$
where the tangent space $\operatorname{T}_{(\mathbf{U},\mathbf{V})}\!\mathcal{M}$ is defined by $\operatorname{T}_{(\mathbf{U},\mathbf{V})}\!\mathcal{M}\simeq\operatorname{T}_{\mathbf{U}}\!\mathcal{M}_{1}\times\operatorname{T}_{\mathbf{V}}\!\mathcal{M}_{2}$ and
(4) $\operatorname{T}_{\mathbf{U}}\!\mathcal{M}_{1}=\{\mathbf{U}\boldsymbol{\Omega}_{1}+\mathbf{U}_{\Sigma_{xx}\perp}\mathbf{K}_{1}:\boldsymbol{\Omega}_{1}\in\mathbb{R}^{m\times m},\boldsymbol{\Omega}_{1}^{\mathsf{T}}=-\boldsymbol{\Omega}_{1},\mathbf{K}_{1}\in\mathbb{R}^{(d_{x}-m)\times m}\}$
is the tangent space of the generalized Stiefel manifold $\mathcal{M}_{1}$ with dimension $md_{x}-m(m+1)/2$, the matrix $\mathbf{U}_{\Sigma_{xx}\perp}\in\mathbb{R}^{d_{x}\times(d_{x}-m)}$ satisfies that $(\mathbf{U}_{\Sigma_{xx}\perp})^{\mathsf{T}}\Sigma_{xx}\mathbf{U}_{\Sigma_{xx}\perp}=\mathbf{I}_{d_{x}-m}$ and $\mathbf{U}^{\mathsf{T}}\Sigma_{xx}\mathbf{U}_{\Sigma_{xx}\perp}=0$. The $\operatorname{T}_{\mathbf{V}}\!\mathcal{M}_{2}$ is defined in a same fashion.
In our framework, it is equivalent that the operators in Eq. 5 are defined by ¯${\bar{\mathcal{H}}}_{1}(\mathbf{U},\mathbf{V})[\eta_{1}]=\Sigma_{xx}\eta_{1}$ and ¯${\bar{\mathcal{H}}}_{2}(\mathbf{U},\mathbf{V})[\eta_{2}]=\Sigma_{yy}\eta_{2}$, which have left preconditioning effect. The orthogonal projection with respect to $g$ of a vector ${\bar{\eta}}\in\operatorname{T}_{(\mathbf{U},\mathbf{V})}\!\mathcal{E}\simeq\mathcal{E}$ onto $\operatorname{T}_{(\mathbf{U},\mathbf{V})}\!\mathcal{M}$ is given by $\Pi_{g,(\mathbf{U},\mathbf{V})}({\bar{\eta}})=\left({\bar{\eta}}_{1}-\mathbf{U}\operatorname{sym}(\mathbf{U}^{\mathsf{T}}\Sigma_{xx}{\bar{\eta}}_{1}),{\bar{\eta}}_{2}-\mathbf{V}\operatorname{sym}(\mathbf{V}^{\mathsf{T}}\Sigma_{yy}{\bar{\eta}}_{2})\right)$, where $\mathcal{E}=\mathbb{R}^{d_{x}\times m}\times\mathbb{R}^{d_{y}\times m}$ is the ambient space of $\mathcal{M}$. Therefore, it follows from Eq. 4 that the Riemannian gradient reads
(5) $\mathrm{grad}_{g}f(\mathbf{U},\mathbf{V})=($ $-\Sigma_{xx}^{-1}\Sigma_{xy}\mathbf{V}\mathbf{N}+\mathbf{U}\operatorname{sym}(\mathbf{U}^{\mathsf{T}}\Sigma_{xy}\mathbf{V}\mathbf{N}),$
$-\Sigma_{yy}^{-1}\Sigma_{xy}^{\mathsf{T}}\mathbf{U}\mathbf{N}+\mathbf{V}\operatorname{sym}(\mathbf{V}^{\mathsf{T}}\Sigma_{xy}^{\mathsf{T}}\mathbf{U}\mathbf{N})).$
Since the local convergence rate of Riemannian optimization methods is closely related to the condition number $\kappa_{g}(\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$ (see Section 2.2), we first compute the Riemannian Hessian of $f$ at $(\mathbf{U},\mathbf{V})$ along $\eta=(\eta_{1},\eta_{2})\in\operatorname{T}_{(\mathbf{U},\mathbf{V})}\!\mathcal{M}$ by using Eq. 1
(6) $\operatorname{Hess}_{g}\!f(\mathbf{U},\mathbf{V})[\eta]=\Pi_{g,(\mathbf{U},\mathbf{V})}\big{(}$ $\eta_{1}\operatorname{sym}(\mathbf{U}^{\mathsf{T}}\Sigma_{xy}\mathbf{V}\mathbf{N})+\mathbf{U}\operatorname{sym}(\eta_{1}^{\mathsf{T}}\Sigma_{xy}\mathbf{V}\mathbf{N})$
$+\mathbf{U}\operatorname{sym}(\mathbf{U}^{\mathsf{T}}\Sigma_{xy}\eta_{2}\mathbf{N})-\Sigma_{xx}^{-1}\Sigma_{xy}\eta_{2}\mathbf{N},$
$\eta_{2}\operatorname{sym}(\mathbf{V}^{\mathsf{T}}\Sigma_{xy}^{\mathsf{T}}\mathbf{U}\mathbf{N})+\mathbf{V}\operatorname{sym}(\eta_{2}^{\mathsf{T}}\Sigma_{xy}^{\mathsf{T}}\mathbf{U}\mathbf{N})$
$+\mathbf{V}\operatorname{sym}(\mathbf{V}^{\mathsf{T}}\Sigma_{xy}^{\mathsf{T}}\eta_{1}\mathbf{N})-\Sigma_{yy}^{-1}\Sigma_{xy}^{\mathsf{T}}\eta_{1}\mathbf{N}\big{)},$
Then, the condition number of $\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*})$ can be computed as follows.
Proposition 4.1.
Let $\sigma_{1}>\sigma_{2}>\cdots>\sigma_{m+1}\geq\cdots\geq\sigma_{\min\{d_{x},d_{y}\}}$ be singular values of the matrix $\Sigma_{xx}^{-1/2}\Sigma_{xy}\Sigma_{yy}^{-1/2}$. It holds that
$\kappa_{g}(\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))=\frac{\max\left\{\frac{1}{2}(\mu_{1}+\mu_{2})(\sigma_{1}+\sigma_{2}),\mu_{1}(\sigma_{1}+\sigma_{m+1})\right\}}{\min\{\min_{i,j\in[m],i\neq j}\frac{1}{2}(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j}),\mu_{m}(\sigma_{m}-\sigma_{m+1})\}},$
where $[m]:=\{1,2,\dots,m\}$.
A proof is given in [25, Theorem 5]. Nevertheless, we provide a slightly different proof for Proposition 4.1 in Appendix C for the sake of brevity, since the proofs of Propositions 4.4 and 5.1 stem from the details of the proof of Proposition 4.1. Note that Proposition 4.1 boils down to $\kappa_{g}(\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))=(\sigma_{1}+\sigma_{2})/(\sigma_{1}-\sigma_{2})$ for $m=1$, which coincides with the result in [26, Lemma 4.1].
4.2 New left and right preconditioner
Observing from the second-order information in Eq. 6, we aim to approximate the diagonal blocks of Eq. 6 and propose a new metric where the operators in Eq. 5 have both left and right preconditioning effect. To this end, we adopt the left and right preconditioning in Section 3.2 and propose a new Riemannian metric
(7) $g_{\mathrm{new},(\mathbf{U},\mathbf{V})}(\xi,\eta):=\langle\xi_{1},\Sigma_{xx}\eta_{1}\mathbf{M}_{1,2}\rangle+\langle\xi_{2},\Sigma_{yy}\eta_{2}\mathbf{M}_{2,2}\rangle,$
where $\mathbf{M}_{1,2}:=(\operatorname{sym}(\mathbf{U}^{\mathsf{T}}\Sigma_{xy}\mathbf{V}\mathbf{N})^{2}+\delta\mathbf{I}_{m})^{1/2}$, $\mathbf{M}_{2,2}:=(\operatorname{sym}(\mathbf{V}^{\mathsf{T}}\Sigma_{xy}^{\mathsf{T}}\mathbf{U}\mathbf{N})^{2}+\delta\mathbf{I}_{m})^{1/2}$, and $\delta>0$. The projection operator $\Pi_{\mathrm{new},(\mathbf{U},\mathbf{V})}$ is given by the following proposition.
Proposition 4.2.
Given the new metric Eq. 7, the orthogonal projection operator on $\operatorname{T}_{(\mathbf{U},\mathbf{V})}\!\mathcal{M}$ is given by
(8) $\Pi_{\mathrm{new},(\mathbf{U},\mathbf{V})}({\bar{\eta}})=(\Pi_{\mathrm{new},\mathbf{U}}({\bar{\eta}}_{1}),\Pi_{\mathrm{new},\mathbf{V}}({\bar{\eta}}_{2}))=({\bar{\eta}}_{1}-\mathbf{U}\mathbf{S}_{1}\mathbf{M}_{1,2}^{-1},{\bar{\eta}}_{2}-\mathbf{V}\mathbf{S}_{2}\mathbf{M}_{2,2}^{-1})$
for ${\bar{\eta}}\in\operatorname{T}_{(\mathbf{U},\mathbf{V})}\!\mathcal{E}\simeq\mathcal{E}$, where $\mathbf{S}_{1}$, $\mathbf{S}_{2}$ are the unique solutions of the Lyapunov equations $\mathbf{M}_{1,2}^{-1}\mathbf{S}_{1}+\mathbf{S}_{1}\mathbf{M}_{1,2}^{-1}=2\operatorname{sym}(\mathbf{U}^{\mathsf{T}}\Sigma_{xx}{\bar{\eta}}_{1})$ and $\mathbf{M}_{2,2}^{-1}\mathbf{S}_{2}+\mathbf{S}_{2}\mathbf{M}_{2,2}^{-1}=2\operatorname{sym}(\mathbf{V}^{\mathsf{T}}\Sigma_{yy}{\bar{\eta}}_{2})$.
Proof 4.3.
See Appendix D.
It follows from Propositions 3.2 and 4.2 that the Riemannian gradient of $f$ at $(\mathbf{U},\mathbf{V})\in\mathcal{M}$ is
(9) $\mathrm{grad}_{\mathrm{new}}f(\mathbf{U},\mathbf{V})=-((\Sigma_{xx}^{-1}\Sigma_{xy}\mathbf{V}\mathbf{N}+\mathbf{U}\mathbf{S}_{1})\mathbf{M}_{1,2}^{-1},(\Sigma_{yy}^{-1}\Sigma_{xy}^{\mathsf{T}}\mathbf{U}\mathbf{N}+\mathbf{V}\mathbf{S}_{2})\mathbf{M}_{2,2}^{-1}).$
Since $\mathbf{M}_{1,2},\mathbf{M}_{2,2}\in\mathbb{R}^{m\times m}$ and $m\ll\min\{d_{x},d_{y}\}$, the computational cost of the Riemannian gradient under the new metric Eq. 7 is comparable to one under Eq. 3. Subsequently, the Riemannian Hessian of $f$ at $(\mathbf{U}^{*},\mathbf{V}^{*})$ along $\eta$ is given by
$\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*})[\eta]$ $=\Pi_{\mathrm{new},(\mathbf{U}^{*},\mathbf{V}^{*})}(\mathrm{D}\bar{G}_{\mathrm{new}}(\mathbf{U}^{*},\mathbf{V}^{*})[\eta])$
$=\Pi_{\mathrm{new},(\mathbf{U}^{*},\mathbf{V}^{*})}($ $-\Sigma_{xx}^{-1}\Sigma_{xy}\eta_{2}\mathbf{N}\mathbf{M}_{1,2}^{-1}+\Sigma_{xx}^{-1}\Sigma_{xy}\mathbf{V}^{*}\mathbf{N}\mathbf{M}_{1,2}^{-1}\dot{\mathbf{M}}_{1,2}\mathbf{M}_{1,2}^{-1}$
$-\eta_{1}\mathbf{S}_{1}\mathbf{M}_{1,2}^{-1}-\mathbf{U}^{*}\dot{\mathbf{S}}_{1}\mathbf{M}_{1,2}^{-1}+\mathbf{U}^{*}\mathbf{S}_{1}\mathbf{M}_{1,2}^{-1}\dot{\mathbf{M}}_{1,2}\mathbf{M}_{1,2}^{-1},$
$-\Sigma_{yy}^{-1}\Sigma_{xy}^{\mathsf{T}}\eta_{1}\mathbf{N}\mathbf{M}_{2,2}^{-1}+\Sigma_{yy}^{-1}\Sigma_{xy}^{\mathsf{T}}\mathbf{U}^{*}\mathbf{N}\mathbf{M}_{2,2}^{-1}\dot{\mathbf{M}}_{2,2}\mathbf{M}_{2,2}^{-1}$
$-\eta_{2}\mathbf{S}_{2}\mathbf{M}_{2,2}^{-1}-\mathbf{V}^{*}\dot{\mathbf{S}}_{2}\mathbf{M}_{2,2}^{-1}+\mathbf{V}^{*}\mathbf{S}_{2}\mathbf{M}_{2,2}^{-1}\dot{\mathbf{M}}_{2,2}\mathbf{M}_{2,2}^{-1})$
where $\bar{G}_{\mathrm{new}}:\mathcal{E}\to\mathbb{R}$ is a smooth extension of $\mathrm{grad}_{\mathrm{new}}f$, $\dot{\mathbf{M}}_{1,2}:=\mathrm{D}\mathbf{M}_{1,2}(\mathbf{U}^{*},\mathbf{V}^{*})[\eta]$, $\dot{\mathbf{M}}_{2,2}:=\mathrm{D}\mathbf{M}_{2,2}(\mathbf{U}^{*},\mathbf{V}^{*})[\eta]$, and the symmetric matrices $\dot{\mathbf{S}}_{1}$ and $\dot{\mathbf{S}}_{2}$ satisfy the Lyapunov equations
$\operatorname{sym}(\dot{\mathbf{M}}_{1,2}\mathbf{S}_{1}+{\mathbf{M}}_{1,2}\dot{\mathbf{S}}_{1}+\dot{\mathbf{M}}_{1,2}\Sigma\mathbf{N}+{\mathbf{M}}_{1,2}(\eta_{1}^{\mathsf{T}}\Sigma_{xy}\mathbf{V}^{*}+(\mathbf{U}^{*})^{\mathsf{T}}\Sigma_{xy}\eta_{2}))$ $=0,$
$\operatorname{sym}(\dot{\mathbf{M}}_{2,2}\mathbf{S}_{2}+{\mathbf{M}}_{2,2}\dot{\mathbf{S}}_{2}+\dot{\mathbf{M}}_{2,2}\Sigma\mathbf{N}+{\mathbf{M}}_{2,2}(\eta_{2}^{\mathsf{T}}\Sigma_{xy}^{\mathsf{T}}\mathbf{U}^{*}+(\mathbf{V}^{*})^{\mathsf{T}}\Sigma_{xy}^{\mathsf{T}}\eta_{1}))$ $=0.$
Finally, we illustrate the effect of the metric Eq. 7 by computing the condition number of $\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*})$ in the following proposition, which can be proved in a similar fashion as in Proposition 4.1.
Proposition 4.4.
Let $\sigma_{1}>\sigma_{2}>\cdots>\sigma_{m+1}\geq\cdots\geq\sigma_{\min\{d_{x},d_{y}\}}$ be the singular values of the matrix $\Sigma_{xx}^{-1/2}\Sigma_{xy}\Sigma_{yy}^{-1/2}$. Then, the condition number at the local minimizer $(\mathbf{U}^{*},\mathbf{V}^{*})$ is computed by
(10) $\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))=\frac{\max\{\underset{i,j\in[m],i\neq j}{\max}\frac{(\mu_{i}+\mu_{j})(\sigma_{i}+\sigma_{j})}{\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta}},\frac{\mu_{1}(\sigma_{1}+\sigma_{m+1})}{\sqrt{\mu_{1}^{2}\sigma_{1}^{2}+\delta}}\}}{\min\{\underset{i,j\in[m],i\neq j}{\min}\frac{(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j})}{\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta}},\frac{\mu_{m}(\sigma_{m}-\sigma_{m+1})}{\sqrt{\mu_{m}^{2}\sigma_{m}^{2}+\delta}}\}}.$
Proof 4.5.
We aim to compute the Rayleigh quotient Eq. 4 under the proposed metric Eq. 7, and evaluate its upper and lower bounds. To this end, we firstly simplify the Riemannian Hessian $\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*})[\eta]$ as follows.
By solving the Lyapunov equations in Proposition 4.2, we obtain that $\mathbf{S}_{1}=\mathbf{S}_{2}=-\Sigma\mathbf{N}$, where $\Sigma=\operatorname{diag}(\sigma_{1},\sigma_{2},\dots,\sigma_{m})$. Since $(\mathbf{U}^{*},\mathbf{V}^{*})$ is a critical point of $f$, it follows from Proposition 2.2 that $\mathrm{grad}_{\mathrm{new}}f(\mathbf{U}^{*},\mathbf{V}^{*})=0$ and thus $\Sigma_{xx}^{-1}\Sigma_{xy}\mathbf{V}^{*}=\mathbf{U}^{*}\Sigma$ and $\Sigma_{yy}^{-1}\Sigma_{xy}^{\mathsf{T}}\mathbf{U}^{*}=\mathbf{V}^{*}\Sigma$. Hence, we can simplify the Riemannian Hessian $\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*})[\eta]$ to
$\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*})[\eta]$ $=\Pi_{\mathrm{new},(\mathbf{U}^{*},\mathbf{V}^{*})}(\mathrm{D}\bar{G}_{\mathrm{new}}(\mathbf{U}^{*},\mathbf{V}^{*})[\eta])$
$=\Pi_{\mathrm{new},(\mathbf{U}^{*},\mathbf{V}^{*})}($ $-\Sigma_{xx}^{-1}\Sigma_{xy}\eta_{2}\mathbf{N}(\mathbf{M}_{1,2}^{*})^{-1}+\eta_{1}\Sigma\mathbf{N}(\mathbf{M}_{1,2}^{*})^{-1}-\mathbf{U}^{*}\dot{\mathbf{S}}_{1}(\mathbf{M}_{1,2}^{*})^{-1},$
$-\Sigma_{yy}^{-1}\Sigma_{xy}^{\mathsf{T}}\eta_{1}\mathbf{N}(\mathbf{M}_{2,2}^{*})^{-1}+\eta_{2}\Sigma\mathbf{N}(\mathbf{M}_{2,2}^{*})^{-1}-\mathbf{V}^{*}\dot{\mathbf{S}}_{2}(\mathbf{M}_{2,2}^{*})^{-1}),$
where $\mathbf{M}_{1,2}^{*}=\mathbf{M}_{2,2}^{*}=(\Sigma^{2}\mathbf{N}^{2}+\delta\mathbf{I}_{m})^{1/2}$.
Subsequently, we compute the Rayleigh quotient. It follows from Eq. 4 that
$q(\eta)$ $=\frac{g_{\mathrm{new},(\mathbf{U}^{*},\mathbf{V}^{*})}(\eta,\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*})[\eta])}{g_{\mathrm{new},(\mathbf{U}^{*},\mathbf{V}^{*})}(\eta,\eta)}$
$=\frac{\langle\eta_{1},\Sigma_{xx}\eta_{1}\Sigma\mathbf{N}\rangle-2\langle\eta_{1},\Sigma_{xy}\eta_{2}\mathbf{N}\rangle+\langle\eta_{2},\Sigma_{yy}\eta_{2}\Sigma\mathbf{N}\rangle}{\langle\eta_{1},\Sigma_{xx}\eta_{1}\mathbf{M}_{1,2}^{*}\rangle+\langle\eta_{2},\Sigma_{yy}\eta_{2}\mathbf{M}_{2,2}^{*}\rangle}$
for $\eta=(\eta_{1},\eta_{2})\in\operatorname{T}_{(\mathbf{U}^{*},\mathbf{V}^{*})}\!\mathcal{M}$, where we use the facts that $\dot{\mathbf{S}}_{1}$ and $\dot{\mathbf{S}}_{2}$ are symmetric and $\langle\eta_{1},\Sigma_{xx}\mathbf{U}^{*}\dot{\mathbf{S}}_{1}\rangle=\langle\eta_{2},\Sigma_{yy}\mathbf{V}^{*}\dot{\mathbf{S}}_{2}\rangle=0$ by Eq. 4.
Consequently, we can evaluate the upper and lower bounds of $q(\eta)$ in a similar fashion as Appendix C.
Improved local convergence rate
We illustrate the effect of the proposed Riemannian metric Eq. 7 through the improved condition number, which is able to accelerate the Riemannian methods in the sense of Theorem 2.6. To this end, we first provide Lemma 4.6 to simplify Eq. 9. Next, we prove that $\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))\leq\kappa_{g}(\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$ in Proposition 4.8.
Lemma 4.6.
Denote $\bar{v}_{ij}(\delta):=(\mu_{i}+\mu_{j})(\sigma_{i}+\sigma_{j})/(\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta})$ and $\underline{v}_{ij}(\delta):=(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j})/(\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta})$. It holds that
$\bar{v}_{ij}(0)>\bar{v}_{ik}(0)>\bar{v}_{jk}(0)\qquad\text{and}\qquad\underline{v}_{ij}(0)<\underline{v}_{ik}(0)<\underline{v}_{jk}(0)$
for all $1\leq i<j<k\leq m$ with $m\geq 3$.
Proof 4.7.
First, we prove that $\bar{v}_{ij}(0)>\bar{v}_{ik}(0)$. Since $\mu_{i}>\mu_{j}>\mu_{k}$ and $\sigma_{i}>\sigma_{j}>\sigma_{k}$, it holds that
$\bar{v}_{ij}(0)-\bar{v}_{ik}(0)$ $=\frac{(\mu_{i}\sigma_{j}+\mu_{j}\sigma_{i})(\mu_{i}\sigma_{i}+\mu_{k}\sigma_{k})-(\mu_{i}\sigma_{k}+\mu_{k}\sigma_{i})(\mu_{i}\sigma_{i}+\mu_{j}\sigma_{j})}{(\mu_{i}\sigma_{i}+\mu_{j}\sigma_{j})(\mu_{i}\sigma_{i}+\mu_{k}\sigma_{k})}$
$=\frac{(\mu_{i}^{2}-\mu_{j}\mu_{k})\sigma_{i}(\sigma_{j}-\sigma_{k})+\mu_{i}(\mu_{j}-\mu_{k})(\sigma_{i}^{2}-\sigma_{j}\sigma_{k})}{(\mu_{i}\sigma_{i}+\mu_{j}\sigma_{j})(\mu_{i}\sigma_{i}+\mu_{k}\sigma_{k})}>0.$
Therefore, $\bar{v}_{ij}(0)>\bar{v}_{ik}(0)$ holds and $\bar{v}_{ik}(0)>\bar{v}_{jk}(0)$ is also true. The result $\underline{v}_{ij}(0)<\underline{v}_{ik}(0)<\underline{v}_{jk}(0)$ is ready by using $\bar{v}_{ij}(0)+\underline{v}_{ij}(0)=2$, $\bar{v}_{ik}(0)+\underline{v}_{ik}(0)=2$ and $\bar{v}_{jk}(0)+\underline{v}_{jk}(0)=2$.
Then, we can prove that $\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))\leq\kappa_{g}(\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$ for $m\geq 2$. Note that $\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))=\kappa_{g}(\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$ for $m=1$ since the right preconditioners in Eq. 7 boil down to scalars that no longer have preconditioning effect.
Proposition 4.8.
Assume that $m\geq 2$. There exists a constant $\bar{\delta}>0$, such that
$\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))\leq\kappa_{g}(\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$
holds for all $\delta\in(0,\bar{\delta})$ in Eq. 7.
Proof 4.9.
First, it follows from Lemma 4.6 and the continuity of $\bar{v}_{ij}$ and $\underline{v}_{ij}$ with respect to $\delta\in[0,\infty)$ that there exists a constant $\bar{\delta}>0$, such that
$\bar{v}_{ij}(\delta)>\bar{v}_{ik}(\delta)>\bar{v}_{jk}(\delta)\qquad\text{and}\qquad\underline{v}_{ij}(\delta)<\underline{v}_{ik}(\delta)<\underline{v}_{jk}(\delta)$
hold for all $1\leq i<j<k\leq m$ and $\delta\in(0,\bar{\delta})$. Therefore, we have $\bar{v}_{12}(\delta)=\max_{i,j\in[m],i\neq j}\bar{v}_{ij}(\delta)$ and $\underline{v}_{12}(\delta)=\min_{i,j\in[m],i\neq j}\underline{v}_{ij}(\delta)$. Subsequently, we observe that
$~{}~{}~{}~{}\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$
$=\frac{\max\{\underset{i,j\in[m],i\neq j}{\max}\bar{v}_{ij}(\delta),\frac{\mu_{1}(\sigma_{1}+\sigma_{m+1})}{\sqrt{\mu_{1}^{2}\sigma_{1}^{2}+\delta}}\}}{\min\{\underset{i,j\in[m],i\neq j}{\min}\underline{v}_{ij}(\delta),\frac{\mu_{m}(\sigma_{m}-\sigma_{m+1})}{\sqrt{\mu_{m}^{2}\sigma_{m}^{2}+\delta}}\}}$
$=\frac{\max\{\bar{v}_{12}(\delta),\frac{\mu_{1}(\sigma_{1}+\sigma_{m+1})}{\sqrt{\mu_{1}^{2}\sigma_{1}^{2}+\delta}}\}}{\min\{\underline{v}_{12}(\delta),\frac{\mu_{m}(\sigma_{m}-\sigma_{m+1})}{\sqrt{\mu_{m}^{2}\sigma_{m}^{2}+\delta}}\}}$
(11) $=\frac{\max\{\frac{1}{2}(\mu_{1}+\mu_{2})(\sigma_{1}+\sigma_{2}),\frac{\mu_{1}(\sigma_{1}+\sigma_{m+1})}{2\sqrt{\mu_{1}^{2}\sigma_{1}^{2}+\delta}}(\sqrt{\mu_{1}^{2}\sigma_{1}^{2}+\delta}+\sqrt{\mu_{2}^{2}\sigma_{2}^{2}+\delta})\}}{\min\{\frac{1}{2}(\mu_{1}-\mu_{2})(\sigma_{1}-\sigma_{2}),\frac{\mu_{m}(\sigma_{m}-\sigma_{m+1})}{2\sqrt{\mu_{m}^{2}\sigma_{m}^{2}+\delta}}(\sqrt{\mu_{1}^{2}\sigma_{1}^{2}+\delta}+\sqrt{\mu_{2}^{2}\sigma_{2}^{2}+\delta})\}}$
(12) $\leq\frac{\max\{\frac{1}{2}(\mu_{1}+\mu_{2})(\sigma_{1}+\sigma_{2}),\mu_{1}(\sigma_{1}+\sigma_{m+1})\}}{\min\{\underset{i,j\in[m],i\neq j}{\min}\frac{1}{2}(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j}),\mu_{m}(\sigma_{m}-\sigma_{m+1})\}}$
$=\kappa_{g}(\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*})),$
where we multiply both the numerator and the denominator by $\frac{1}{2}(\sqrt{\mu_{1}^{2}\sigma_{1}^{2}+\delta}+\sqrt{\mu_{2}^{2}\sigma_{2}^{2}+\delta})$ to yield Eq. 11, and we use the fact that $2\sqrt{\mu_{m}^{2}\sigma_{m}^{2}+\delta}<\sqrt{\mu_{1}^{2}\sigma_{1}^{2}+\delta}+\sqrt{\mu_{2}^{2}\sigma_{2}^{2}+\delta}<2\sqrt{\mu_{1}^{2}\sigma_{1}^{2}+\delta}$ to yield Eq. 12.
It is worth noting that the parameter $\delta>0$ theoretically ensures that Eq. 7 is a Riemannian metric. In practice, one can choose a sufficiently small $\delta$, e.g., $\delta=10^{-15}$.
4.3 RGD and RCG for canonical correlation analysis
By using the Riemannian metric Eq. 7 and required ingredients, we adapt the Riemannian gradient descent (Algorithm 1) and Riemannian conjugate gradient (Algorithm 2) methods to solve the CCA problem in Algorithms 3 and 4.
Algorithm 3 RGD for CCA
0:  $\mathcal{M}$ endowed with a metric Eq. 7, initial guess $(\mathbf{U}^{(0)},\mathbf{V}^{(0)})\in\mathcal{M}$, $t=0$.
1:  while the stopping criteria are not satisfied do
2:     Compute $\eta^{(t)}=-\mathrm{grad}_{g}f(\mathbf{U}^{(t)},\mathbf{V}^{(t)})$ by Eq. 9.
3:     Compute the stepsize $s^{(t)}$ by Armijo backtracking Eq. 2.
4:     Update $\mathbf{U}^{(t+1)}=\Sigma_{xx}^{-1/2}\operatorname{qf}(\Sigma_{xx}^{1/2}(\mathbf{U}^{(t)}+\eta_{1}^{(t)}))$, $\mathbf{V}^{(t+1)}=\Sigma_{yy}^{-1/2}\operatorname{qf}(\Sigma_{yy}^{1/2}(\mathbf{V}^{(t)}+\eta_{2}^{(t)}))$; $t=t+1$.
5:  end while
5:  $(\mathbf{U}^{(t)},\mathbf{V}^{(t)})\in\mathcal{M}$.
Algorithm 4 RCG for CCA
0:  $\mathcal{M}$ endowed with a metric Eq. 7, initial guess $(\mathbf{U}^{(0)},\mathbf{V}^{(0)})\in\mathcal{M}$, $t=0$, $\beta^{(0)}=0$.
1:  while the stopping criteria are not satisfied do
2:     Compute $\eta^{(t)}=-\mathrm{grad}_{g}f(\mathbf{U}^{(t)},\mathbf{U}^{(t)})+\beta^{(t)}\Pi_{g,(\mathbf{U}^{(t)},\mathbf{V}^{(t)})}(\eta^{(t-1)})$ by Eq. 9.
3:     Compute the stepsize $s^{(t)}$ by Armijo backtracking Eq. 2.
4:     Update $\mathbf{U}^{(t+1)}=\Sigma_{xx}^{-1/2}\operatorname{qf}(\Sigma_{xx}^{1/2}(\mathbf{U}^{(t)}+\eta_{1}^{(t)}))$, $\mathbf{V}^{(t+1)}=\Sigma_{yy}^{-1/2}\operatorname{qf}(\Sigma_{yy}^{1/2}(\mathbf{V}^{(t)}+\eta_{2}^{(t)}))$; $t=t+1$.
5:  end while
5:  $(\mathbf{U}^{(t)},\mathbf{V}^{(t)})\in\mathcal{M}$.
Note that 1) the retraction mapping is the generalized QR decomposition [22] with respect to $\Sigma_{xx}$ and $\Sigma_{yy}$, i.e.,
$\operatorname{R}_{(\mathbf{U},\mathbf{V})}(\eta):=(\Sigma_{xx}^{-\frac{1}{2}}\operatorname{qf}(\Sigma_{xx}^{\frac{1}{2}}(\mathbf{U}+\eta_{1})),\Sigma_{yy}^{-\frac{1}{2}}\operatorname{qf}(\Sigma_{yy}^{\frac{1}{2}}(\mathbf{V}+\eta_{2})))\quad\text{for }\eta\in\operatorname{T}_{(\mathbf{U},\mathbf{V})}\!\mathcal{M},$
where $\operatorname{qf}(\mathbf{X})$ refers to the $\mathbf{Q}$ factor in the QR decomposition $\mathbf{Q}\mathbf{R}=\mathbf{X}$. In practice, the retraction can be efficiently computed [22] by $\operatorname{R}_{(\mathbf{U},\mathbf{V})}(\eta)=((\mathbf{U}+\eta_{1})\mathbf{R}_{1}^{-1},(\mathbf{V}+\eta_{2})\mathbf{R}_{2}^{-1})$ instead, where $\mathbf{R}_{1}^{\mathsf{T}}\mathbf{R}_{1}=(\mathbf{U}+\eta_{1})^{\mathsf{T}}\Sigma_{xx}(\mathbf{U}+\eta_{1})$ and $\mathbf{R}_{2}^{\mathsf{T}}\mathbf{R}_{2}=(\mathbf{V}+\eta_{2})^{\mathsf{T}}\Sigma_{yy}(\mathbf{V}+\eta_{2})$ are Cholesky decompositions; 2) the vector transport in Algorithm 4 is defined by the projection operator in Eq. 8, i.e., $\mathcal{T}_{t\leftarrow t-1}(\eta)=\Pi_{g,(\mathbf{U}^{(t)},\mathbf{V}^{(t)})}(\eta)$ for $\eta\in\operatorname{T}_{(\mathbf{U}^{(t-1)},\mathbf{V}^{(t-1)})}\!\mathcal{M}$.
4.4 Numerical validation
Algorithms 3 and 4 are implemented in toolbox Manopt v7.1.0 [6], a Matlab library for Riemannian methods. The stopping criteria are the same as default settings in Manopt. All experiments are performed on a MacBook Pro 2019 with MacOS Ventura 13.3, 2.4 GHz 8 core Intel Core i9 processor, 32GB memory, and Matlab R2020b. The codes are available at https://github.com/JimmyPeng1998/popman.
Table 2: Compared metrics in CCA.
(E) (L1) (L2) (L12) (LR12)
¯${\bar{\mathcal{H}}}_{1}(\mathbf{U},\mathbf{V})[\eta_{1}]$ $\eta_{1}$ $\Sigma_{xx}\eta_{1}$ $\eta_{1}$ $\Sigma_{xx}\eta_{1}$ $\Sigma_{xx}\eta_{1}\mathbf{M}_{1,2}$
¯${\bar{\mathcal{H}}}_{2}(\mathbf{U},\mathbf{V})[\eta_{2}]$ $\eta_{2}$ $\eta_{2}$ $\Sigma_{yy}\eta_{2}$ $\Sigma_{yy}\eta_{2}$ $\Sigma_{yy}\eta_{2}\mathbf{M}_{2,2}$
We test the performance of RGD and RCG under different metrics, i.e., five different choice of ¯${\bar{\mathcal{H}}}_{1},{\bar{\mathcal{H}}}_{2}$ in $g_{(\mathbf{U},\mathbf{V})}(\xi,\eta)=\langle\xi_{1},{\bar{\mathcal{H}}}_{1}(\mathbf{U},\mathbf{V})[\eta_{1}]\rangle+\langle\xi_{2},{\bar{\mathcal{H}}}_{2}(\mathbf{U},\mathbf{V})[\eta_{2}]\rangle$; see Table 2. The Euclidean metric is denoted by “(E)”. “(L1)” and “(L2)” are the metrics wherein only one component of $\mathcal{M}=\mathcal{M}_{1}\times\mathcal{M}_{2}$ is endowed with a preconditioned metric. The metric Eq. 3 proposed by [26] is called “(L12)”. The metric Eq. 7, denoted by “(LR12)”, has the effect of preconditioning both on the left and right. We set $d_{x}=800$, $d_{y}=400$, $n=30000$, $m=5$, $\delta=10^{-15}$, $\lambda_{x}=\lambda_{y}=10^{-6}$, and $\mathbf{N}=\operatorname{diag}(m,m-1,\dots,1)$. Elements of the data matrices $\mathbf{X}$ and $\mathbf{Y}$ are sampled i.i.d. from the uniform distribution on $[0,1]$. The performance of a method is evaluated by the residual $(f(\mathbf{U},\mathbf{V})-f_{\min})$, gradient norm “gnorm”, and the subspace distances $D(\mathbf{U},\mathbf{U}^{*}):=\|\mathbf{U}\mathbf{U}^{\mathsf{T}}-\mathbf{U}^{*}(\mathbf{U}^{*})^{\mathsf{T}}\|_{\mathrm{F}}$ and $D(\mathbf{V},\mathbf{V}^{*}):=\|\mathbf{V}\mathbf{V}^{\mathsf{T}}-\mathbf{V}^{*}(\mathbf{V}^{*})^{\mathsf{T}}\|_{\mathrm{F}}$, where $f_{\min}=f(\mathbf{U}^{*},\mathbf{V}^{*})$ and $(\mathbf{U}^{*},\mathbf{V}^{*})$ is defined in Eq. 2.
Numerical results are reported in Figs. 3, 4, and 3. We have following observations: 1) the proposed metric Eq. 7 improves the performance of RGD and RCG since it benefits more from the second-order information; 2) Fig. 4 shows that the computation time per iteration of Algorithms 3 and 4 is comparable to RGD(L12) and RCG(L12); 3) Table 3 illustrates that RGD(LR12) and RCG(LR12) require fewer iterations and less time to reach the stopping criteria than the others. The subspace distances are smaller than $10^{-8}$, and hence the sequences generated by proposed methods converge to the correct subspace.
Figure 3: Numerical results for CCA problem for $d_{x}=800$, $d_{y}=400$, and $m=5$. Left: RGD. Right: RCG. Each method is tested for 10 runs.
Figure 4: Computation time per iteration for RGD (left) and RCG (right) under different metrics for CCA problem for $d_{x}=800$, $m=5$, and $d_{y}=200,400,\dots,1000$.
Table 3: Convergence results of the CCA problem for $d_{x}=800$, $d_{y}=400$, and $m=5$.
metric method #iter time (s) gnorm $D(\mathbf{U},\mathbf{U}^{*})$ $D(\mathbf{V},\mathbf{V}^{*})$ $\kappa_{g}$
(E) RGD 10000 249.11 5.95e-02 2.69e-05 2.66e-05 2.10e+04
RCG 1745 31.03 1.70e-05 4.01e-10 3.89e-10
(L1) RGD 10000 255.33 1.02e+00 4.12e-04 4.07e-04 1.43e+07
RCG 2500 74.13 4.94e-02 2.85e-04 2.79e-04
(L2) RGD 10000 245.81 8.20e-01 4.13e-04 4.05e-04 1.52e+07
RCG 2500 56.16 6.90e-02 2.93e-04 2.90e-04
(L12) RGD 10000 274.91 4.67e-04 9.68e-07 9.57e-07 1.12e+04
RCG 937 30.39 8.82e-07 1.68e-09 1.65e-09
(LR12) RGD 6607 195.03 1.34e-06 7.47e-16 7.46e-16 2.38e+03
RCG 410 15.38 8.49e-07 4.63e-09 4.53e-09
Moreover, the condition number of the Riemannian Hessian is numerically computed by the Manopt function hessianspectrum: $\kappa_{g}(\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$ of five metrics are $2.10\cdot 10^{4}$ (E), $1.43\cdot 10^{7}$ (L1), $1.52\cdot 10^{7}$ (L2), $1.12\cdot 10^{4}$ (L12), and $2.38\cdot 10^{3}$ (LR12). It is direct to verify that these numbers coincide with the theoretical results in Propositions 4.1, 4.4, and 4.8. We observe that the Riemannian Hessian under the proposed metric (LR12) has the smallest condition number among all choices, which is reflected in the observation that RGD(LR12) and RCG(LR12) outperform the others.
5 Appplication to truncated singular value decomposition
In this section, the truncated singular value decomposition (SVD) problem is considered. Given a matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$, the $p<\min\{m,n\}$ largest singular vectors $(\mathbf{U}^{*},\mathbf{V}^{*})$ is the global minimizer of the following problem,
(1) $\min\limits_{\mathbf{U},\mathbf{V}}f(\mathbf{U},\mathbf{V}):=-\mathrm{tr}(\mathbf{U}^{\mathsf{T}}\mathbf{A}\mathbf{V}\mathbf{N}),\ \mathrm{s.\,t.}\ (\mathbf{U},\mathbf{V})\in\mathcal{M}:=\mathrm{St}(p,m)\times\mathrm{St}(p,n),$
where $\mathrm{St}(p,m):=\{\mathbf{U}\in\mathbb{R}^{m\times p}:\mathbf{U}^{\mathsf{T}}\mathbf{U}=\mathbf{I}_{p}\}$ is the Stiefel manifold and $\mathbf{N}:=\operatorname{diag}\{\mu_{1},\dots,\mu_{p}\}$ with $\mu_{1}>\mu_{2}>\cdots>\mu_{p}>0$. Sato and Iwai [23] proposed RGD and RCG methods to solve problem Eq. 1, where the search space is endowed with the Euclidean metric. We apply the proposed framework to solve Eq. 1 by endowing $\mathcal{M}$ with a non-Euclidean metric to accelerate the Riemannian optimization methods.
5.1 A new preconditioned metric
Observe that the Riemannian Hessian of $f$ at $(\mathbf{U},\mathbf{V})$ along $\eta=(\eta_{1},\eta_{2})\in\operatorname{T}_{(\mathbf{U},\mathbf{V})}\!\mathcal{M}$ is given by
$\operatorname{Hess}_{\mathrm{e}}\!f(\mathbf{U},\mathbf{V})[\eta]=($ $\eta_{1}\mathbf{M}_{1}-\mathbf{A}\eta_{2}\mathbf{N}-\mathbf{U}\operatorname{sym}(\mathbf{U}^{\mathsf{T}}(\eta_{1}\mathbf{M}_{1}-\mathbf{A}\eta_{2}\mathbf{N})),$
$\eta_{2}\mathbf{M}_{2}-\mathbf{A}^{\mathsf{T}}\eta_{1}\mathbf{N}-\mathbf{V}\operatorname{sym}(\mathbf{V}^{\mathsf{T}}(\eta_{2}\mathbf{M}_{2}-\mathbf{A}^{\mathsf{T}}\eta_{1}\mathbf{N})))$
in [23, Proposition 3.5], where $\mathbf{M}_{1}:=\operatorname{sym}(\mathbf{U}^{\mathsf{T}}\mathbf{A}\mathbf{V}\mathbf{N})$ and $\mathbf{M}_{2}:=\operatorname{sym}(\mathbf{V}^{\mathsf{T}}\mathbf{A}^{\mathsf{T}}\mathbf{U}\mathbf{N})$. Taking advantage of the diagonal blocks of the Riemannian Hessian and the left and right preconditioning in Section 3.1, we define a new preconditioned metric on $\mathcal{M}$:
(2) $g_{\mathrm{new},(\mathbf{U},\mathbf{V})}(\xi,\eta):=\langle\xi_{1},\eta_{1}\mathbf{M}_{1,2}\rangle+\langle\xi_{2},\eta_{2}\mathbf{M}_{2,2}\rangle\quad\text{for}\ \xi,\eta\in\operatorname{T}_{(\mathbf{U},\mathbf{V})}\!\mathcal{M},$
where $\mathbf{M}_{1,2}=(\operatorname{sym}(\mathbf{U}^{\mathsf{T}}\mathbf{A}\mathbf{V}\mathbf{N})^{2}+\delta\mathbf{I}_{p})^{1/2}$, $\mathbf{M}_{2,2}=(\operatorname{sym}(\mathbf{V}^{\mathsf{T}}\mathbf{A}^{\mathsf{T}}\mathbf{U}\mathbf{N})^{2}+\delta\mathbf{I}_{p})^{1/2}$, and $\delta>0$. Note that the left preconditioners are chosen as the identity matrix. The projection operator with respect to Eq. 2 is given by
(3) $\Pi_{\mathrm{new},(\mathbf{U},\mathbf{V})}({\bar{\eta}})=({\bar{\eta}}_{1}-\mathbf{U}\mathbf{S}_{1}\mathbf{M}_{1,2}^{-1},{\bar{\eta}}_{2}-\mathbf{V}\mathbf{S}_{2}\mathbf{M}_{2,2}^{-1})$
for ${\bar{\eta}}\in\operatorname{T}_{(\mathbf{U},\mathbf{V})}\!(\mathbb{R}^{m\times p}\times\mathbb{R}^{n\times p})\simeq\mathbb{R}^{m\times p}\times\mathbb{R}^{n\times p}$, where $\mathbf{S}_{1}$, $\mathbf{S}_{2}$ are the unique solutions of the Lyapunov equations $\mathbf{M}_{1,2}^{-1}\mathbf{S}_{1}+\mathbf{S}_{1}\mathbf{M}_{1,2}^{-1}=2\operatorname{sym}(\mathbf{U}^{\mathsf{T}}{\bar{\eta}}_{1})$, $\mathbf{M}_{2,2}^{-1}\mathbf{S}_{2}+\mathbf{S}_{2}\mathbf{M}_{2,2}^{-1}=2\operatorname{sym}(\mathbf{V}^{\mathsf{T}}{\bar{\eta}}_{2})$. Then, it follows from Propositions 3.2 and 3 that
(4) $\mathrm{grad}_{\mathrm{new}}f(\mathbf{U},\mathbf{V})=(\mathbf{A}\mathbf{V}\mathbf{N}\mathbf{M}_{1,2}^{-1}-\mathbf{U}\mathbf{S}_{1}\mathbf{M}_{1,2}^{-1},\mathbf{A}^{\mathsf{T}}\mathbf{U}\mathbf{N}\mathbf{M}_{2,2}^{-1}-\mathbf{V}\mathbf{S}_{2}\mathbf{M}_{2,2}^{-1}).$
These results can be obtained in a same fashion as in Proposition 4.2 for CCA. Note that the computational cost of the Riemannian gradient Eq. 4 is comparable to one under the Euclidean metric since $\mathbf{M}_{1,2},\mathbf{M}_{2,2}\in\mathbb{R}^{p\times p}$ and $p\ll\min\{m,n\}$.
The effect of the new metric Eq. 2 is illustrated by the following proposition, which can be proved in a same fashion as Propositions 4.1, 4.4, and 4.8.
Proposition 5.1.
Let $\sigma_{1}>\sigma_{2}>\cdots>\sigma_{p}>\sigma_{p+1}\geq\cdots\geq\sigma_{\min\{m,n\}}$ be the singular values of $\mathbf{A}$, $\mathbf{U}^{*}$ and $\mathbf{V}^{*}$ be the $p$ largest left and right singular vectors of $\mathbf{A}$ respectively. It holds that
$\kappa_{\mathrm{e}}(\operatorname{Hess}_{\mathrm{e}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$ $=\frac{\max\left\{\frac{1}{2}(\mu_{1}+\mu_{2})(\sigma_{1}+\sigma_{2}),\mu_{1}(\sigma_{1}+\sigma_{p+1})\right\}}{\min\{\min_{i,j\in[p],i\neq j}\frac{1}{2}(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j}),\mu_{p}(\sigma_{p}-\sigma_{p+1})\}},$
$\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$ $=\frac{\max\{\underset{i,j\in[p],i\neq j}{\max}\frac{(\mu_{i}+\mu_{j})(\sigma_{i}+\sigma_{j})}{\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta}},\frac{\mu_{1}(\sigma_{1}+\sigma_{p+1})}{\sqrt{\sigma_{1}^{2}\mu_{1}^{2}+\delta}}\}}{\min\{\underset{i,j\in[p],i\neq j}{\min}\frac{(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j})}{\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta}},\frac{\mu_{p}(\sigma_{p}-\sigma_{p+1})}{\sqrt{\sigma_{p}^{2}\mu_{p}^{2}+\delta}}\}}.$
Moreover, the new metric Eq. 2 indeed improves the condition number of the Riemannian Hessian, i.e., $\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))\leq\kappa_{\mathrm{e}}(\operatorname{Hess}_{\mathrm{e}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$.
5.2 RGD and RCG for truncated singular value decomposition
Let $\mathcal{M}$ be endowed with the Riemannian metric Eq. 2. We apply the Riemannian gradient descent (Algorithm 1) and Riemannian conjugate gradient (Algorithm 2) methods to solve the SVD problem Eq. 1 in Algorithms 5 and 6. Note that the retraction mapping is based on the QR decomposition, i.e., $\operatorname{R}_{(\mathbf{U},\mathbf{V})}(\eta):=(\operatorname{qf}(\mathbf{U}+\eta_{1}),\operatorname{qf}(\mathbf{V}+\eta_{2}))$ for $\eta\in\operatorname{T}_{(\mathbf{U},\mathbf{V})}\!\mathcal{M}$. The vector transport in Algorithm 6 is defined by the projection operator Eq. 3.
Algorithm 5 RGD for SVD
0:  $\mathcal{M}$ endowed with a metric Eq. 2, initial guess $(\mathbf{U}^{(0)},\mathbf{V}^{(0)})\in\mathcal{M}$, $t=0$.
1:  while the stopping criteria are not satisfied do
2:     Compute $\eta^{(t)}=-\mathrm{grad}_{g}f(\mathbf{U}^{(t)},\mathbf{U}^{(t)})$ by Eq. 4.
3:     Compute the stepsize $s^{(t)}$ by Armijo backtracking line search Eq. 2.
4:     Update $\mathbf{U}^{(t+1)}=\operatorname{qf}(\mathbf{U}^{(t)}+s^{(t)}\eta_{1}^{(t)})$, $\mathbf{V}^{(t+1)}=\operatorname{qf}(\mathbf{V}^{(t)}+s^{(t)}\eta_{2}^{(t)})$; $t=t+1$.
5:  end while
5:  $(\mathbf{U}^{(t)},\mathbf{V}^{(t)})\in\mathcal{M}$.
Algorithm 6 RCG for SVD
0:  $\mathcal{M}$ endowed with a metric Eq. 2, initial guess $(\mathbf{U}^{(0)},\mathbf{V}^{(0)})\in\mathcal{M}$, $t=0$, $\beta^{(0)}=0$.
1:  while the stopping criteria are not satisfied do
2:     Compute $\eta^{(t)}=-\mathrm{grad}_{g}f(\mathbf{U}^{(t)},\mathbf{U}^{(t)})+\beta^{(t)}\Pi_{g,(\mathbf{U}^{(t)},\mathbf{V}^{(t)})}(\eta^{(t-1)})$ by Eq. 4.
3:     Compute the stepsize $s^{(t)}$ by Armijo backtracking line search Eq. 2.
4:     Update $\mathbf{U}^{(t+1)}=\operatorname{qf}(\mathbf{U}^{(t)}+s^{(t)}\eta_{1}^{(t)})$, $\mathbf{V}^{(t+1)}=\operatorname{qf}(\mathbf{V}^{(t)}+s^{(t)}\eta_{2}^{(t)})$; $t=t+1$.
5:  end while
5:  $(\mathbf{U}^{(t)},\mathbf{V}^{(t)})\in\mathcal{M}$.
5.3 Numerical validation
We compare the performance of Algorithms 5 and 6 with RGD and RCG under the Euclidean metric in [23]. The proposed preconditioned metric Eq. 2, which has a right preconditioning effect, is denoted by “(R12)”. We set $m=1000$, $n=500$, $p=10$, and $\mathbf{N}=\operatorname{diag}(p,p-1,\dots,1)$. The matrix $\mathbf{A}$ is constructed by $\mathbf{A}=\mathbf{U}^{*}\Sigma(\mathbf{V}^{*})^{\mathsf{T}}$, where the entries of $\mathbf{U}^{*}\in\mathbb{R}^{m\times p}$ and $\mathbf{V}^{*}\in\mathbb{R}^{n\times p}$ are firstly sampled i.i.d. from the uniform distribution on $[0,1]$, and $\mathbf{U}^{*}$ and $\mathbf{V}^{*}$ are orthogonalized by QR decomposition. We set $\Sigma:=\operatorname{diag}(1,\gamma,\gamma^{2},\dots,\gamma^{p-1})$ and $\gamma=1/1.5$. The implementation of RGD and RCG is the same as Section 4.
Numerical results are shown in Figs. 5, 6, and 4. We have similar observations as the previous experiments in Section 4. First, the proposed methods significantly outperform RGD(E) and RCG(E) with fewer iterations since the proposed metric benefits from the second-order information. Second, computational cost per iteration of Algorithms 5 and 6 is comparable to RGD(E) and RCG(E) respectively. Third, Table 4 shows that the subspace distances are smaller than $10^{-6}$ in RGD(R12) and RCG(R12), which indicates that the sequences generated by proposed methods converge to the correct subspace.
Figure 5: Numerical results for the SVD problem for $m=1000$, $n=500$, and $p=10$. Left: RGD. Right: RCG. Each method is tested for 10 runs
Figure 6: Average computation time per iteration for RGD (left) and RCG (right) under the Euclidean and proposed metric for $m=1000$, $p=10$, and $n=200,400,\dots,1000$
In addition, we compute the condition numbers of $\operatorname{Hess}\!f(\mathbf{U}^{*},\mathbf{V}^{*})$ under two metrics. It follows from the construction of $\mathbf{A}$ and Proposition 5.1 that
$\kappa(\operatorname{Hess}_{\mathrm{e}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$ $=\frac{(\mu_{1}+\mu_{2})(\gamma+1)}{(\mu_{p-1}-\mu_{p})(\gamma^{p-2}-\gamma^{p-1})}=\frac{153389}{63}\approx 2.43\times 10^{3},$
$\kappa(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$ $=\frac{(\mu_{1}+\mu_{2})(1+\gamma)}{(\mu_{1}-\mu_{2})(1-\gamma)}=95,$
which exactly coincide with the numerical results in Table 4. Therefore, the lower condition number suggests faster convergence of the proposed methods.
Table 4: Convergence results of the SVD problem for $m=1000$, $n=500$, and $p=10$
metric method #iter time (s) gnorm $D(\mathbf{U},\mathbf{U}^{*})$ $D(\mathbf{V},\mathbf{V}^{*})$ $\kappa_{g}$
(E) RGD 7781 117.29 9.64e-07 4.53e-05 4.53e-05 2.43e+03
RCG 478 5.44 8.54e-07 2.00e-05 2.00e-05
(R12) RGD 387 3.41 8.72e-07 2.38e-15 1.38e-15 9.50e+01
RCG 105 1.45 7.88e-07 3.26e-07 3.83e-07
6 Application to matrix and tensor completion
In this section, we investigate the matrix and tensor completion problem. Given a partially observed tensor $\mathcal{A}\in\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{d}}$ on an index set $\Omega\subseteq[n_{1}]\times[n_{2}]\times\cdots\times[n_{d}]$, the goal of tensor completion is to recover the tensor $\mathcal{A}$ from its entries on $\Omega$. Note that it boils down to matrix completion for $d=2$.
There are several different formulations in tensor completion. One type is based on the nuclear norm minimization, e.g., [9, 17]. These methods require working with full-size tensors. Instead, tensor decompositions—which take advantage of the low-rank structure of a tensor—reduce the number of parameters in search space. Therefore, it is economical to formulate the tensor completion problem based on a tensor decomposition, which leads to an optimization problem on a product manifold
(1) $\min f(x):=\frac{1}{2p}\left\|\mathrm{Proj}_{\Omega}(\tau(x)-\mathcal{A})\right\|_{\mathrm{F}}^{2},\ \mathrm{s.\,t.}\ x\in\mathcal{M}=\mathcal{M}_{1}\times\mathcal{M}_{2}\times\cdots\times\mathcal{M}_{K},$
where $p:=|\Omega|/(n_{1}n_{2}\cdots n_{d})$ is the sampling rate, $\mathrm{Proj}_{\Omega}$ refers to the projection operator onto $\Omega$, i.e, $\mathrm{Proj}_{\Omega}(\mathcal{X})(i_{1},\dots,i_{d})=\mathcal{X}(i_{1},\dots,i_{d})$ if $(i_{1},\dots,i_{d})\in\Omega$, otherwise $\mathrm{Proj}_{\Omega}(\mathcal{X})(i_{1},\dots,i_{d})=0$ for $\mathcal{X}\in\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$, and $\tau(x)$ denotes the tensor decomposition with components $x_{k}\in\mathcal{M}_{k}$ for $k\in[K]$ and $x=(x_{1},x_{2},\dots,x_{K})$.
Since computing the Euclidean Hessian $\nabla^{2}f(x)$ can be complicated, Kasai and Mishra [15] introduced a preconditioned metric based on the block diagonal approximation of $\nabla^{2}f(x)$ for tensor completion in Tucker decomposition. More recently, the idea became prosperous in low-rank tensor approximation and completion for other tensor formats, e.g., [7, 11, 8, 12], see Table 1 for details. In summary, the metric was developed by constructing an operator ¯${\bar{\mathcal{H}}}(x)$ based on the diagonal blocks of the Hessian of the cost function $\phi(x):=\frac{1}{2}\|\tau(x)-\mathcal{A}\|_{\mathrm{F}}^{2}$, i.e.,
¯${\bar{\mathcal{H}}}(x)[\eta]:=(\partial_{11}^{2}\phi(x)[\eta_{1}],\dots,\partial_{KK}^{2}\phi(x)[\eta_{K}])\quad\text{for }\eta=(\eta_{1},\eta_{2},\dots,\eta_{K})\in\operatorname{T}_{x}\!\mathcal{M},$
where $\partial_{kk}^{2}\phi(x)[\eta_{k}]:=\lim_{h\to 0}(\partial_{k}\phi(x_{1},\dots,x_{k-1},x_{k}+h\eta_{k},x_{k+1},\dots,x_{K})-\partial_{k}\phi(x))/h$ for $k\in[K]$. Note that this preconditioning approach coincides with the exact block diagonal preconditioning in Section 3.1. Alternatively, observing that the cost function $f$ in Eq. 1 enjoys a least-squares structure, we can also adopt the Gauss–Newton type preconditioning in Section 3.3 to solve Eq. 1.
6.1 Gauss–Newton method for tensor ring completion
Since tensor ring decomposition has been shown effective for the tensor completion problem, e.g., [12], we consider the following tensor ring completion problem
(2) $\begin{array}[]{cc}\min\limits_{\mathcal{U}_{k}\in\mathbb{R}^{r_{k-1}\times n_{k}\times r_{k}}}&\ f(\mathcal{U}_{1},\dots,\mathcal{U}_{d}):=\frac{1}{2p}\left\|\mathrm{Proj}_{\Omega}(\llbracket\mathcal{U}_{1},\dots,\mathcal{U}_{d}\rrbracket)-\mathrm{Proj}_{\Omega}(\mathcal{A})\right\|_{\mathrm{F}}^{2},\end{array}$
where $\llbracket\mathcal{U}_{1},\dots,\mathcal{U}_{d}\rrbracket$ denotes the tensor ring decomposition [34]. Specifically, given $\mathcal{X}=\llbracket\mathcal{U}_{1},\dots,\mathcal{U}_{d}\rrbracket$ with $\mathcal{U}_{k}\in\mathbb{R}^{r_{k-1}\times n_{k}\times r_{k}}$ for $k\in[d]$ and $r_{0}=r_{d}$, the $(i_{1},i_{2},\dots,i_{d})$-th element of $\mathcal{X}$ is defined by
$\mathcal{X}(i_{1},i_{2},\dots,i_{d}):=\mathrm{tr}(\mathbf{U}_{1}(i_{1})\mathbf{U}_{2}(i_{2})\cdots\mathbf{U}_{d}(i_{d})),$
where $\mathbf{U}_{k}(i_{k}):=\mathcal{U}_{k}(:,i_{k},:)\in\mathbb{R}^{r_{k-1}\times r_{k}}$ refers the $i_{k}$-th lateral slice of the tensor $\mathcal{U}$ for $i_{k}\in[n_{k}]$. Since the $k$-th unfolding matrix of $\mathcal{X}$ satisfies $\mathbf{X}_{(k)}=(\mathcal{U}_{k})_{(2)}(\mathcal{U}_{\neq k})_{(2)}$, problem Eq. 2 can be reformulated by introducing [12] $\mathbf{W}_{k}:=(\mathcal{U}_{k})_{(2)}$ and $\mathbf{W}_{\neq_{k}}:=(\mathcal{U}_{\neq k})_{(2)}$, where $(\mathcal{U}_{k})_{(2)}$ and $(\mathcal{U}_{\neq k})_{(2)}$ are the 2-nd unfolding matrix of the tensor $\mathcal{U}_{k}$ and $\mathcal{U}_{\neq k}$ respectively, and $\mathcal{U}_{\neq k}\in\mathbb{R}^{r_{k-1}\times\prod_{j\neq k}n_{j}\times r_{k}}$ is defined by its lateral slice matrices, i.e., $\mathbf{U}_{\neq k}(1+\sum_{\ell\neq k,\ell=1}^{d}(i_{\ell}-1)J_{\ell}){:=}(\prod_{j=k+1}^{d}\mathbf{U}_{j}(i_{j})\prod_{j=1}^{k-1}\mathbf{U}_{j}(i_{j}))^{\mathsf{T}}$ with $J_{\ell}:=\prod_{m=1,m\neq k}^{\ell-1}n_{m}$. Consequently, a reformulation of Eq. 2 is given by
(3) $\min_{\vec{\mathbf{W}}}$ $\ \ f({\vec{\mathbf{W}}}):=\frac{1}{2p}\|\mathrm{Proj}_{\Omega}(\tau({\vec{\mathbf{W}}})-\mathcal{A})\|_{\mathrm{F}}^{2}$
$\mathrm{s.\,t.}$ $\ \ \vec{\mathbf{W}}\in\mathcal{M}=\mathbb{R}^{n_{1}\times r_{0}r_{1}}\times\mathbb{R}^{n_{2}\times r_{1}r_{2}}\times\cdots\times\mathbb{R}^{n_{d}\times r_{d-1}r_{d}},$
where the mapping $\tau$ is defined by $\tau:{\vec{\mathbf{W}}}\mapsto\llbracket\operatorname{ten}_{(2)}(\mathbf{W}_{1}),\operatorname{ten}_{(2)}(\mathbf{W}_{2}),\dots,\operatorname{ten}_{(2)}(\mathbf{W}_{d})\rrbracket$ and $\operatorname{ten}_{(2)}(\cdot)$ is the second tensorization operator.
Noticing that the cost function $f$ in Eq. 3 enjoys a least-squares structure: $f({\vec{\mathbf{W}}})=\frac{1}{2}\|F({\vec{\mathbf{W}}})\|_{\mathrm{F}}^{2}$, where $F({\vec{\mathbf{W}}})=\mathrm{Proj}_{\Omega}(\tau({\vec{\mathbf{W}}})-\mathcal{A})/\sqrt{p}$ is a smooth function, we adopt the Gauss–Newton type preconditioning to solve Eq. 3. Since the search space $\mathcal{M}$ is flat, the RGD method under the metric Eq. 6 is essentially a Euclidean Gauss–Newton method (see [20, §10.3]). We list the Gauss–Newton method in Algorithm 7; see Appendix E for implementation details.
Algorithm 7 Gauss–Newton method for tensor ring completion (TR-GN)
0:  $\mathcal{M}$ endowed with a metric $g$, initial guess ${\vec{\mathbf{W}}}^{(0)}\in\mathcal{M}$, $t=0$.
1:  while the stopping criteria are not satisfied do
2:     Compute $\eta^{(t)}$ by solving Eq. 7.
3:     Update ${\vec{\mathbf{W}}}^{(t+1)}={\vec{\mathbf{W}}}^{(t)}+\eta^{(t)}$; $t=t+1$.
4:  end while
4:  ${\vec{\mathbf{W}}}^{(t)}\in\mathcal{M}$.
Since the tensor ring decomposition is complicated, we leave the condition number results analogous to Propositions 4.1, 4.4, and 5.1 for future work. Nevertheless, if the sequence generated by the Gauss–Newton method (Algorithm 7) converges to ${\vec{\mathbf{W}}}^{*}\in\mathcal{M}$ with $F({\vec{\mathbf{W}}}^{*})=0$, the Gauss–Newton method enjoys superlinear convergence, see [2, §8.4.1] for Riemannian Gauss–Newton method and [20, §10.3] for the Euclidean Gauss–Newton method.
6.2 Numerical validation
We compare Algorithm 7 with the Riemannian gradient descent (TR-RGD) and the Riemannian conjugate gradient (TR-RCG) methods in [12] under the metric $g_{{\vec{\mathbf{W}}}}(\xi,\eta):=\sum_{k=1}^{d}\langle\xi_{k},\eta_{k}(\mathbf{W}_{\neq k}^{\mathsf{T}}\mathbf{W}_{\neq k}+\delta\mathbf{I}_{r_{k-1}r_{k}})\rangle$ for $\xi,\eta\in\operatorname{T}_{{\vec{\mathbf{W}}}}\!\mathcal{M}$, where $\delta>0$ is a constant. The codes for TR-RGD, TR-RCG, and TR-GN methods are available at https://github.com/JimmyPeng1998/LRTCTR.
The tensor $\mathcal{A}\in\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{d}}$ is constructed by $\mathcal{A}=\tau({\vec{\mathbf{W}}}^{*})$ and each entry of ${\vec{\mathbf{W}}}^{*}\in\mathcal{M}$ is uniformly sampled from $[0,1]$. The initial guess ${\vec{\mathbf{W}}}^{(0)}\in\mathcal{M}$ is generated in a same fashion. Given the sampling rate $p$, we randomly select $pn_{1}n_{2}\cdots n_{d}$ samples from $[n_{1}]\times[n_{2}]\times\cdots\times[n_{d}]$ to formulate the sampling set $\Omega$. We set $d=3$, $n_{1}=n_{2}=n_{3}=100$, $p=0.05$, TR ranks $\mathbf{r}^{*}=(1,1,1),(2,2,2),\dots,(8,8,8)$, and $\delta=10^{-15}$.
Figure 7: Training and test errors for TR rank $\mathbf{r}^{*}=(5,5,5)$. Each method is tested for 10 runs
We specify the default settings of all methods. The stepsize rule for TR-RGD method and the TR-RCG method is the Armijo backtracking line search Eq. 2. The conjugate gradient parameter is set to be the Riemannian version [6] of the modified Hestenes–Stiefel rule. The parameters in Eq. 2 are $\rho=0.3$, $a=2^{-13}$, and $s_{0}=1$. The performance of each method is evaluated by the training error $\varepsilon_{\Omega}({\vec{\mathbf{W}}}^{(t)}):=\|\mathrm{Proj}_{\Omega}(\tau({\vec{\mathbf{W}}}^{(t)}))-\mathrm{Proj}_{\Omega}(\mathcal{A})\|_{\mathrm{F}}/\|\mathrm{Proj}_{\Omega}(\mathcal{A})\|_{\mathrm{F}}$ and the test error $\varepsilon_{\Gamma}({\vec{\mathbf{W}}}^{(t)})$, where $\Gamma$ is a test set different from $\Omega$ and we set $|\Gamma|=100$. A method is terminated if one of the following stopping criteria is achieved: 1) training error $\varepsilon_{\Omega}({\vec{\mathbf{W}}}^{(t)})<10^{-14}$; 2) the maximum iteration $1000$; 3) the relative change $\lvert{(\varepsilon_{\Omega}({\vec{\mathbf{W}}}^{(t)})-\varepsilon_{\Omega}({\vec{\mathbf{W}}}^{(t-1)}))}/{\varepsilon_{\Omega}({\vec{\mathbf{W}}}^{(t-1)})}\rvert<\varepsilon$; 4) the stepsize $s^{(t)}<10^{-10}$.
Figure 8: Computation time required for each method to reach the stopping criteria for TR rank $\mathbf{r}^{*}=(1,1,1),(2,2,2),\dots,(8,8,8)$
Numerical results are illustrated in Figs. 7 and 8. On the one hand, we observe that the TR-GN method has faster convergence than TR-RGD and TR-RCG since TR-GN exploits more second-order information of $\nabla^{2}f({\vec{\mathbf{W}}})$, while the preconditioned metric in TR-RGD and TR-RCG only takes advantage of its diagonal blocks. On the other hand, Fig. 8 suggests that the computation time for TR-GN to reach the stopping criteria grows faster than TR-RGD and TR-RCG as TR rank $\mathbf{r}^{*}$ increases. In other words, there is a trade-off between exploiting second-order information and the computational efficiency.
7 Conclusions and future works
The performance of the Riemannian methods varies with different metrics. We have shown that an exquisitely constructed metric is indeed helpful to accelerate the Riemannian methods. Specifically, we have proposed a general framework for optimization on product manifolds endowed with a preconditioned metric and three specific approaches to construct an operator that aims to approximate the Riemannian Hessian. Conceptually, various existing methods including the Riemannian Gauss–Newton method and the block-Jacobi preconditioning in numerical linear algebra can be interpreted by the proposed framework with specific metrics. We have tailored novel preconditioned metrics based on the proposed framework for canonical correlation analysis and truncated singular value decomposition and have shown the effect of the proposed metric by computing the condition number of the Riemannian Hessian at the local minimizer, which indeed improves the condition numbers. Numerical results verify that a delicate metric does improve the performance of the Riemannian optimization methods.
In the future, we intend to specify the proposed framework to other problems, e.g., the distributed PCA. Moreover, taking into account the block structure of product manifolds, parallel computing is capable of further accelerating the Riemannian optimization methods.
Acknowledgments
We would like to thank the editor and two anonymous reviewers for insightful comments.
Appendix A Computational details in Example 1.1
In Example 1.1, we investigate a class of preconditioned metrics defined by
$g_{\lambda,\mathbf{x}}=\langle\xi,\mathbf{B}_{\lambda}\eta\rangle\quad\text{for }\xi,\eta\in\operatorname{T}_{\mathbf{x}}\!\mathcal{M}_{\mathbf{B}},$
where $\lambda\in\mathbb{R}$ such that the matrix $\mathbf{B}_{\lambda}:=\lambda\mathbf{I}_{n}+(1-\lambda)\mathbf{B}$ is positive definite. The Riemannian gradients at $x\in\mathcal{M}_{\mathbf{B}}$ under these metrics are
$\mathrm{grad}_{g_{\lambda,\mathbf{x}}}f(x)=-(\mathbf{I}_{n}-\frac{\mathbf{B}_{\lambda}^{-1}\mathbf{B}\mathbf{x}\mathbf{x}^{\mathsf{T}}\mathbf{B}}{\mathbf{x}^{\mathsf{T}}\mathbf{B}\mathbf{B}_{\lambda}^{-1}\mathbf{B}\mathbf{x}})\mathbf{B}_{\lambda}^{-1}\mathbf{b}=-\mathbf{B}_{\lambda}^{-1}\mathbf{b}+\frac{\mathbf{x}^{\mathsf{T}}\mathbf{B}\mathbf{B}_{\lambda}^{-1}\mathbf{b}}{\mathbf{x}^{\mathsf{T}}\mathbf{B}\mathbf{B}_{\lambda}^{-1}\mathbf{B}\mathbf{x}}\mathbf{B}_{\lambda}^{-1}\mathbf{B}\mathbf{x},$
by following Proposition 4.2. Subsequently, the update rule for RGD under the metric $g_{\lambda,\mathbf{x}}$ is $\mathbf{x}^{(t+1)}={\bar{\mathbf{x}}^{(t)}}/{\|\bar{\mathbf{x}}^{(t)}\|_{\mathbf{B}}}$ with $\bar{\mathbf{x}}^{(t)}=\mathbf{x}^{(t)}-s^{(t)}\mathrm{grad}_{g_{\lambda,\mathbf{x}}}f(\mathbf{x}^{(t)})$, where we adopt the polar retraction [26, (3.3)]. The Riemannian Hessian of $f$ at $\mathbf{x}^{*}$ along $\eta\in\operatorname{T}_{\mathbf{x}^{*}}\!\mathcal{M}_{\mathbf{B}}$ can be computed by
$\operatorname{Hess}_{g_{\lambda,\mathbf{x}}}\!f(\mathbf{x}^{*})[\eta]$ $=\Pi_{g_{\lambda,\mathbf{x}},\mathbf{x}^{*}}(\mathrm{D}\mathrm{grad}_{g_{\lambda,\mathbf{x}}}f(\mathbf{x}^{*})[\eta])=\Pi_{g_{\lambda,\mathbf{x}},\mathbf{x}^{*}}(\|\mathbf{B}^{-1}\mathbf{b}\|_{\mathbf{B}}{\mathbf{B}_{\lambda}}^{-1}\mathbf{B}\eta),$
since $\mathrm{grad}_{g_{\lambda,\mathbf{x}}}f(\mathbf{x}^{*})=\Pi_{g_{\lambda,\mathbf{x}},\mathbf{x}^{*}}(-\mathbf{B}_{\lambda}^{-1}\mathbf{b})=0$. The Rayleigh quotient Eq. 4 is given by
$q(\eta)=\|\mathbf{B}^{-1}\mathbf{b}\|_{\mathbf{B}}\cdot\frac{\langle\eta,\mathbf{B}\eta\rangle}{\langle\eta,{\mathbf{B}_{\lambda}}\eta\rangle}\quad\text{for }\eta\in\operatorname{T}_{\mathbf{x}^{*}}\!\mathcal{M}_{\mathbf{B}}.$
Consequently, we can compute the condition number of $\operatorname{Hess}_{g_{\lambda,\mathbf{x}}}\!f(\mathbf{x}^{*})$ in a same fashion as Proposition 4.4. Note that if $\lambda=0$, the Rayleigh quotient boils down to a constant $\|\mathbf{B}^{-1}\mathbf{b}\|_{\mathbf{B}}$ and thus $\kappa_{g_{0,\mathbf{x}}}(\operatorname{Hess}_{g_{0,\mathbf{x}}}\!f(\mathbf{x}^{*}))=1$.
Appendix B Proof of Proposition 2.2
Proof B.1.
It suffices to prove the first inequality, and the other can be proved in a same fashion. Let $(\mathcal{U},\varphi)$ be a chart of the manifold $\mathcal{M}$, and $E_{i}$ be the $i$-th coordinate vector field. For vector fields $\zeta=\sum_{i}\alpha_{i}E_{i}$ and $\chi=\sum_{i}\beta_{i}E_{i}$, it follows the definition of the Riemannian metric $g$ that $g_{x}(\zeta_{x},\chi_{x})=\sum_{i,j}g_{ij}\alpha_{i}\beta_{j}=\hat{\zeta}_{\hat{x}}^{\mathsf{T}}\mathbf{G}_{\hat{x}}\hat{\chi}_{\hat{x}}$, where $\hat{x}:=\varphi(x)$, $\hat{\zeta}_{\hat{x}}:=\mathrm{D}\varphi(\varphi^{-1}(\hat{x}))[\zeta_{x}]$, $\hat{\chi}_{\hat{x}}:=\mathrm{D}\varphi(\varphi^{-1}(\hat{x}))[\chi_{x}]$, and the $(i,j)$-th element of $\mathbf{G}_{\hat{x}}$ is $g_{ij}:=g(E_{i},E_{j})$. Denote $\zeta_{x}:=\mathrm{grad}_{g}f(x)$ and $\chi_{x}:=\mathrm{grad}_{\tilde{g}}f(x)$. It follows from the coordinate expression [2, §3.6] that $\hat{\zeta}_{\hat{x}}=\mathbf{G}^{-1}_{\hat{x}}\nabla\hat{f}(\hat{x})$ and $\hat{\chi}_{\hat{x}}=\tilde{\mathbf{G}}_{\hat{x}}^{-1}\nabla\hat{f}(\hat{x})$, where $\hat{f}(\hat{x}):=f\circ\varphi^{-1}(\hat{x})$ and $\nabla\hat{f}$ refers to the Euclidean gradient of $\hat{f}$. We obtain that
$g_{x}(\mathrm{grad}_{g}f(x),\mathrm{grad}_{\tilde{g}}f(x))=\hat{\zeta}_{\hat{x}}^{\mathsf{T}}\mathbf{G}_{\hat{x}}\hat{\chi}_{\hat{x}}=(\nabla\hat{f}(\hat{x}))^{\mathsf{T}}\tilde{\mathbf{G}}_{\hat{x}}^{-1}\nabla\hat{f}(\hat{x})\geq 0.$
The equality holds if and only if $\nabla\hat{f}(\hat{x})=0$, i.e., $\mathrm{grad}_{g}f(x)=\mathrm{grad}_{\tilde{g}}f(x)=0$. Moreover, if $\mathrm{grad}_{g}f(x)=0$, it follows that $\hat{\zeta}_{\hat{x}}=0$, and hence $\hat{\chi}_{\hat{x}}=\tilde{\mathbf{G}}_{\hat{x}}^{-1}\nabla\hat{f}(\hat{x})=\tilde{\mathbf{G}}_{\hat{x}}^{-1}\mathbf{G}_{\hat{x}}\mathbf{G}^{-1}_{\hat{x}}\nabla\hat{f}(\hat{x})=\tilde{\mathbf{G}}_{\hat{x}}^{-1}\mathbf{G}_{\hat{x}}\hat{\zeta}_{\hat{x}}=0$, i.e., $\mathrm{grad}_{\tilde{g}}f(x)=0$.
Appendix C Proof of Proposition 4.1
Proposition 4.1 gives the condition number of the Riemannian Hessian of $f$ at $(\mathbf{U}^{*},\mathbf{V}^{*})$ under the metric Eq. 3. A concrete proof of Proposition 4.1 is given as follows.
Proof C.1.
Since $(\mathbf{U}^{*},\mathbf{V}^{*})$ is a critical point of $f$, it follows from $(\mathbf{U}^{*})^{\mathsf{T}}\Sigma_{xy}\mathbf{V}^{*}=\Sigma$ and $\mathrm{grad}_{g}f(\mathbf{U}^{*},\mathbf{V}^{*})=0$ that
(1) $\Sigma_{xx}^{-1}\Sigma_{xy}\mathbf{V}^{*}=\mathbf{U}^{*}\Sigma\quad\text{and}\quad\Sigma_{yy}^{-1}\Sigma_{xy}^{\mathsf{T}}\mathbf{U}^{*}=\mathbf{V}^{*}\Sigma,$
where $\Sigma=\operatorname{diag}(\sigma_{1},\sigma_{2},\dots,\sigma_{m})$ and $\sigma_{1}>\sigma_{2}>\cdots>\sigma_{m}$ are the $m$ largest singular values of $\Sigma_{xx}^{-1/2}\Sigma_{xy}\Sigma_{yy}^{-1/2}$. Therefore, it follows from Eqs. 4, 6, and 1 that
$g_{(\mathbf{U}^{*},\mathbf{V}^{*})}(\eta,\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*})[\eta])=\langle\eta_{1},\Sigma_{xx}\eta_{1}\Sigma\mathbf{N}\rangle-2\langle\eta_{1},\Sigma_{xy}\eta_{2}\mathbf{N}\rangle+\langle\eta_{2},\Sigma_{yy}\eta_{2}\Sigma\mathbf{N}\rangle$
for $\eta=(\eta_{1},\eta_{2})\in\operatorname{T}_{(\mathbf{U}^{*},\mathbf{V}^{*})}\!\mathcal{M}$.
The goal of the proof is to compute the condition number of $\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*})$. To this end, we calculate the Rayleigh quotient of $\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*})$ by Eq. 4 and evaluate its upper and lower bounds. First, the Rayleigh quotient reads
$q(\eta)$ $=\frac{\langle\eta_{1},\Sigma_{xx}\eta_{1}\Sigma\mathbf{N}\rangle-2\langle\eta_{1},\Sigma_{xy}\eta_{2}\mathbf{N}\rangle+\langle\eta_{2},\Sigma_{yy}\eta_{2}\Sigma\mathbf{N}\rangle}{\langle\eta_{1},\Sigma_{xx}\eta_{1}\rangle+\langle\eta_{2},\Sigma_{yy}\eta_{2}\rangle}$
$=\frac{\langle\tilde{\eta}_{1},\tilde{\eta}_{1}\Sigma\mathbf{N}\rangle-2\langle\tilde{\eta}_{1},\Sigma_{xx}^{-1/2}\Sigma_{xy}\Sigma_{yy}^{-1/2}\tilde{\eta}_{2}\mathbf{N}\rangle+\langle\tilde{\eta}_{2},\tilde{\eta}_{2}\Sigma\mathbf{N}\rangle}{\langle\tilde{\eta}_{1},\tilde{\eta}_{1}\rangle+\langle\tilde{\eta}_{2},\tilde{\eta}_{2}\rangle}$
$=\frac{\begin{bmatrix}\mathrm{vec}(\tilde{\eta}_{1})^{\mathsf{T}}&\mathrm{vec}(\tilde{\eta}_{2})^{\mathsf{T}}\end{bmatrix}\begin{bmatrix}\Sigma\mathbf{N}\otimes\mathbf{I}_{d_{x}}&-\mathbf{N}\otimes\mathbf{M}\\ -\mathbf{N}\otimes\mathbf{M}^{\mathsf{T}}&\Sigma\mathbf{N}\otimes\mathbf{I}_{d_{y}}\end{bmatrix}\begin{bmatrix}\mathrm{vec}(\tilde{\eta}_{1})\\ \mathrm{vec}(\tilde{\eta}_{2})\end{bmatrix}}{\langle\tilde{\eta}_{1},\tilde{\eta}_{1}\rangle+\langle\tilde{\eta}_{2},\tilde{\eta}_{2}\rangle}$
(2) $=\frac{\sum\limits_{i=1}^{m}\mu_{i}\begin{bmatrix}(\tilde{\eta}_{1}(:,i))^{\mathsf{T}}&(\tilde{\eta}_{2}(:,i))^{\mathsf{T}}\end{bmatrix}\begin{bmatrix}\sigma_{i}\mathbf{I}_{d_{x}}&-\mathbf{M}\\ -\mathbf{M}^{\mathsf{T}}&\sigma_{i}\mathbf{I}_{d_{y}}\end{bmatrix}\begin{bmatrix}\tilde{\eta}_{1}(:,i)\\ \tilde{\eta}_{2}(:,i)\end{bmatrix}}{\langle\tilde{\eta}_{1},\tilde{\eta}_{1}\rangle+\langle\tilde{\eta}_{2},\tilde{\eta}_{2}\rangle},$
where $\mathbf{M}=\Sigma_{xx}^{-1/2}\Sigma_{xy}\Sigma_{yy}^{-1/2}$, $\tilde{\eta}_{1}=\Sigma_{xx}^{1/2}\eta_{1}$ and $\tilde{\eta}_{2}=\Sigma_{yy}^{1/2}\eta_{2}$. By using Eq. 4, $\mathbf{U}^{*}=\Sigma_{xx}^{-1/2}\bar{\mathbf{U}}$ and $\mathbf{V}^{*}=\Sigma_{yy}^{-1/2}\bar{\mathbf{V}}$, we can represent $\tilde{\eta}$ by
(3) $\tilde{\eta}=(\Sigma_{xx}^{\frac{1}{2}}\eta_{1},\Sigma_{yy}^{\frac{1}{2}}\eta_{2})=(\bar{\mathbf{U}}\boldsymbol{\Omega}_{1}+\bar{\mathbf{U}}_{\perp}\mathbf{K}_{1},\bar{\mathbf{V}}\boldsymbol{\Omega}_{2}+\bar{\mathbf{V}}_{\perp}\mathbf{K}_{2}),$
where $\bar{\mathbf{U}}_{\perp}\in\mathrm{St}(d_{x}-m,d_{x})$ and $\bar{\mathbf{V}}_{\perp}\in\mathrm{St}(d_{y}-m,d_{y})$ satisfy $\bar{\mathbf{U}}^{\mathsf{T}}\bar{\mathbf{U}}_{\perp}=0$, $\bar{\mathbf{V}}^{\mathsf{T}}\bar{\mathbf{V}}_{\perp}=0$ and $\mathbf{M}=[\bar{\mathbf{U}}\ \bar{\mathbf{U}}_{\perp}]\operatorname{diag}(\sigma_{1},\dots,\sigma_{r},0,\dots,0)[\bar{\mathbf{V}}\ \bar{\mathbf{V}}_{\perp}]^{\mathsf{T}}$ with $r=\mathrm{rank}(\mathbf{M})$. Taking Eq. 3 into Eq. 2, we obtain that
$q(\eta)$ $=\dfrac{\sum\limits_{i=1}^{m}\mu_{i}\begin{bmatrix}[\bar{\mathbf{U}}\ \bar{\mathbf{U}}_{\perp}]\bar{\boldsymbol{\Omega}}_{1}(:,i)\\ [\bar{\mathbf{V}}\ \bar{\mathbf{V}}_{\perp}]\bar{\boldsymbol{\Omega}}_{2}(:,i)\end{bmatrix}^{\mathsf{T}}\begin{bmatrix}\sigma_{i}\mathbf{I}_{d_{x}}\ -\mathbf{M}\\ -\mathbf{M}^{\mathsf{T}}\ \sigma_{i}\mathbf{I}_{d_{y}}\end{bmatrix}\begin{bmatrix}[\bar{\mathbf{U}}\ \bar{\mathbf{U}}_{\perp}]\bar{\boldsymbol{\Omega}}_{1}(:,i)\\ [\bar{\mathbf{V}}\ \bar{\mathbf{V}}_{\perp}]\bar{\boldsymbol{\Omega}}_{2}(:,i)\end{bmatrix}}{\langle\tilde{\eta}_{1},\tilde{\eta}_{1}\rangle+\langle\tilde{\eta}_{2},\tilde{\eta}_{2}\rangle}$
(4) $=\dfrac{\sum\limits_{i=1}^{m}\mu_{i}\Big{(}-\sum\limits_{j=1}^{r}2\sigma_{j}\bar{\boldsymbol{\Omega}}_{1}(j,i)\bar{\boldsymbol{\Omega}}_{2}(j,i)+\sum\limits_{j=1}^{d_{x}}\sigma_{i}\bar{\boldsymbol{\Omega}}_{1}(j,i)^{2}+\sum\limits_{j=1}^{d_{y}}\sigma_{i}\bar{\boldsymbol{\Omega}}_{2}(j,i)^{2}\Big{)}}{\|\bar{\boldsymbol{\Omega}}_{1}\|_{\mathrm{F}}^{2}+\|\bar{\boldsymbol{\Omega}}_{2}\|_{\mathrm{F}}^{2}},$
where $\bar{\boldsymbol{\Omega}}_{\ell}:=\begin{bmatrix}\boldsymbol{\Omega}_{\ell}\\ \mathbf{K}_{\ell}\end{bmatrix}$ for $\ell=1,2$.
Subsequently, by using $\bar{\boldsymbol{\Omega}}_{\ell}(j,i)^{2}=\bar{\boldsymbol{\Omega}}_{\ell}(i,j)^{2}$ for $\ell=1,2$, $i,j\in[m]$, we regroup the terms $\bar{\boldsymbol{\Omega}}_{\ell}(j,i)^{2}$ in Eq. 4 and yield
(5) $q(\eta)$ $\leq\dfrac{\sum\limits_{i=1}^{m}\mu_{i}\Big{(}\sum\limits_{j=1}^{r}\sigma_{j}(\bar{\boldsymbol{\Omega}}_{1}(j,i)^{2}+\bar{\boldsymbol{\Omega}}_{2}(j,i)^{2})+\sum\limits_{j=1}^{d_{x}}\sigma_{i}\bar{\boldsymbol{\Omega}}_{1}(j,i)^{2}+\sum\limits_{j=1}^{d_{y}}\sigma_{i}\bar{\boldsymbol{\Omega}}_{2}(j,i)^{2}\Big{)}}{\|\bar{\boldsymbol{\Omega}}_{1}\|_{\mathrm{F}}^{2}+\|\bar{\boldsymbol{\Omega}}_{2}\|_{\mathrm{F}}^{2}}$
$=\dfrac{\sum\limits_{i=1}^{m}\Big{(}\sum\limits_{j=1}^{r}\bar{s}_{ij}(\bar{\boldsymbol{\Omega}}_{1}(j,i)^{2}+\bar{\boldsymbol{\Omega}}_{2}(j,i)^{2})+\sum\limits_{j=r+1}^{d_{x}}\mu_{i}\sigma_{i}\bar{\boldsymbol{\Omega}}_{1}(j,i)^{2}+\sum\limits_{j=r+1}^{d_{y}}\mu_{i}\sigma_{i}\bar{\boldsymbol{\Omega}}_{2}(j,i)^{2}\Big{)}}{\|\bar{\boldsymbol{\Omega}}_{1}\|_{\mathrm{F}}^{2}+\|\bar{\boldsymbol{\Omega}}_{2}\|_{\mathrm{F}}^{2}}$
$\leq\max\left\{(\mu_{1}+\mu_{2})(\sigma_{1}+\sigma_{2})/2,\mu_{1}(\sigma_{1}+\sigma_{m+1})\right\},$
where $\bar{s}_{ij}:=\begin{cases}(\mu_{i}+\mu_{j})(\sigma_{i}+\sigma_{j})/2,\ j=1,2,\dots,m;\\ \mu_{i}(\sigma_{i}+\sigma_{j}),\ j=m+1,m+2,\dots,r\end{cases}$ for $i=1,2,\dots,m$. The equality holds if and only if: 1) $\bar{\boldsymbol{\Omega}}_{1}(j,i)=-\bar{\boldsymbol{\Omega}}_{2}(j,i)$ for all $i\in[m],j\in[r]$; 2) $\bar{\boldsymbol{\Omega}}_{1}(j,i)^{2}=0$ for all $i\in[m],j=r+1,r+2,\dots,d_{x}$; 3) $\bar{\boldsymbol{\Omega}}_{2}(j,i)^{2}=0$ for all $i\in[m],j=r+1,r+2,\dots,d_{y}$; 4) $\bar{\boldsymbol{\Omega}}_{1}(j,i)=\bar{\boldsymbol{\Omega}}_{2}(j,i)=0$ for $(i,j)\neq(i^{*},j^{*})$, where $(i^{*},j^{*})\in\operatorname{arg\,max}_{i\in[m],j\in[r],i\neq j}\bar{s}_{ij}\subseteq\{(1,2),(2,1),(1,m+1)\}$.
Additionally, we compute the lower bound of Rayleigh quotient in a same fashion as Eq. 5 and yield
(6) $q(\eta)$ $\geq\dfrac{\sum\limits_{i=1}^{m}\mu_{i}\Big{(}-\sum\limits_{j=1}^{r}\sigma_{j}(\bar{\boldsymbol{\Omega}}_{1}(j,i)^{2}+\bar{\boldsymbol{\Omega}}_{2}(j,i)^{2})+\sum\limits_{j=1}^{d_{x}}\sigma_{i}\bar{\boldsymbol{\Omega}}_{1}(j,i)^{2}+\sum\limits_{j=1}^{d_{y}}\sigma_{i}\bar{\boldsymbol{\Omega}}_{2}(j,i)^{2}\Big{)}}{\|\bar{\boldsymbol{\Omega}}_{1}\|_{\mathrm{F}}^{2}+\|\bar{\boldsymbol{\Omega}}_{2}\|_{\mathrm{F}}^{2}}$
$=\dfrac{\sum\limits_{i=1}^{m}\Big{(}\sum\limits_{j=1}^{r}\underline{s}_{ij}(\bar{\boldsymbol{\Omega}}_{1}(j,i)^{2}+\bar{\boldsymbol{\Omega}}_{2}(j,i)^{2})+\sum\limits_{j=r+1}^{d_{x}}\mu_{i}\sigma_{i}\bar{\boldsymbol{\Omega}}_{1}(j,i)^{2}+\sum\limits_{j=r+1}^{d_{y}}\mu_{i}\sigma_{i}\bar{\boldsymbol{\Omega}}_{2}(j,i)^{2}\Big{)}}{\|\bar{\boldsymbol{\Omega}}_{1}\|_{\mathrm{F}}^{2}+\|\bar{\boldsymbol{\Omega}}_{2}\|_{\mathrm{F}}^{2}}$
$\geq\min\{\min_{i,j\in[m],i\neq j}(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j})/2,\mu_{m}(\sigma_{m}-\sigma_{m+1})\},$
where $\underline{s}_{ij}:=\begin{cases}(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j})/2,\ j=1,2,\dots,m;\\ \mu_{i}(\sigma_{i}-\sigma_{j}),\ j=m+1,m+2,\dots,r\end{cases}$ for $i=1,2,\dots,m$. The equality holds if and only if: 1) $\bar{\boldsymbol{\Omega}}_{1}(j,i)=\bar{\boldsymbol{\Omega}}_{2}(j,i)$ for all $i\in[m],j\in[r]$; 2) $\bar{\boldsymbol{\Omega}}_{1}(j,i)^{2}=0$ for all $i\in[m],j=r+1,\dots,d_{x}$; 3) $\bar{\boldsymbol{\Omega}}_{2}(j,i)^{2}=0$ for all $i\in[m],j=r+1,\dots,d_{y}$; 4) $\bar{\boldsymbol{\Omega}}_{1}(j,i)=\bar{\boldsymbol{\Omega}}_{2}(j,i)=0$ for $(i,j)\neq(i^{*},j^{*})$, where $(i^{*},j^{*})\in\operatorname{arg\,min}_{i\in[m],j\in[r],i\neq j}\underline{s}_{ij}$.
Since the inequalities in Eqs. 5 and 6 are tight, the proof is completed.
Appendix D Proof of Proposition 4.2
Proof D.1.
It suffices to prove $\Pi_{\mathrm{new},\mathbf{U}}({\bar{\eta}}_{1})={\bar{\eta}}_{1}-\mathbf{U}\mathbf{S}_{1}\mathbf{M}_{1,2}^{-1}$, and the others can be obtained in a same fashion. Recall the tangent space of $\mathcal{M}_{1}$ in Eq. 4. The orthogonal complement with regard to the metric Eq. 7 of the tangent space $\operatorname{T}_{\mathbf{U}}\!\mathcal{M}_{1}$ can be characterized by
$(\operatorname{T}_{\mathbf{U}}\!\mathcal{M}_{1})^{\perp}=\{\mathbf{U}\mathbf{S}_{1}\mathbf{M}_{1,2}^{-1}:\mathbf{S}_{1}\in\mathbb{R}^{m\times m},\ \mathbf{S}_{1}=\mathbf{S}_{1}^{\mathsf{T}}\},$
since the dimension of $\{\mathbf{U}\mathbf{S}_{1}\mathbf{M}_{1,2}^{-1}:\mathbf{S}_{1}\in\mathbb{R}^{m\times m},\ \mathbf{S}_{1}=\mathbf{S}_{1}^{\mathsf{T}}\}$ is $m(m+1)/2$ and $\mathrm{tr}((\mathbf{U}\mathbf{S}_{1}\mathbf{M}_{1,2}^{-1})^{\mathsf{T}}\Sigma_{xx}(\mathbf{U}\boldsymbol{\Omega}_{1}+\mathbf{U}_{\Sigma_{xx}\perp}\mathbf{K}_{1})\mathbf{M}_{1,2})=0$ holds for all $\mathbf{S}_{1},\boldsymbol{\Omega}_{1},\mathbf{K}_{1}$ satisfying that $\mathbf{S}_{1}=\mathbf{S}_{1}^{\mathsf{T}}$ and $\boldsymbol{\Omega}_{1}=-\boldsymbol{\Omega}_{1}^{\mathsf{T}}$. Moreover, in the light of $\operatorname{T}_{\mathbf{U}}\!\mathcal{M}_{1}\oplus(\operatorname{T}_{\mathbf{U}}\!\mathcal{M}_{1})^{\perp}=\operatorname{T}_{\mathbf{U}}\!\mathbb{R}^{d_{x}\times m}\simeq\mathbb{R}^{d_{x}\times m}$, there is a unique orthogonal decomposition for $\bar{\eta}_{1}\in\mathbb{R}^{d_{x}\times m}$
(1) ${\bar{\eta}}_{1}=\Pi_{\mathrm{new},\mathbf{U}}({\bar{\eta}}_{1})+\Pi_{\mathrm{new},\mathbf{U}}^{\perp}({\bar{\eta}}_{1})=(\mathbf{U}\boldsymbol{\Omega}_{1}+\mathbf{U}_{\Sigma_{xx}\perp}\mathbf{K}_{1})+\mathbf{U}\mathbf{S}_{1}\mathbf{M}_{1,2}^{-1},$
i.e., $\Pi_{\mathrm{new},\mathbf{U}}({\bar{\eta}}_{1})={\bar{\eta}}_{1}-\Pi_{\mathrm{new},\mathbf{U}}^{\perp}({\bar{\eta}}_{1})={\bar{\eta}}_{1}-\mathbf{U}\mathbf{S}_{1}\mathbf{M}_{1,2}^{-1}$. To characterize the symmetric matrix $\mathbf{S}_{1}$, we multiply Eq. 1 from the left by $\mathbf{U}^{\mathsf{T}}\Sigma_{xx}$, and yield $\mathbf{U}^{\mathsf{T}}\Sigma_{xx}{\bar{\eta}}_{1}=\boldsymbol{\Omega}_{1}+\mathbf{S}_{1}\mathbf{M}_{1,2}^{-1}$. Summing up $\mathbf{U}^{\mathsf{T}}\Sigma_{xx}{\bar{\eta}}_{1}$ and $(\mathbf{U}^{\mathsf{T}}\Sigma_{xx}{\bar{\eta}}_{1})^{\mathsf{T}}$, we obtain that $\mathbf{S}_{1}\mathbf{M}_{1,2}^{-1}+\mathbf{M}_{1,2}^{-1}\mathbf{S}_{1}=\mathbf{U}^{\mathsf{T}}\Sigma_{xx}{\bar{\eta}}_{1}+{\bar{\eta}}_{1}^{\mathsf{T}}\Sigma_{xx}\mathbf{U}$, which has a unique solution according to [14, Theorem 2.4.4.1].
Appendix E Implementation details of TR-GN
Recall that the search direction $\eta^{(t)}$ in Algorithm 7 is determined by the following least-squares problem
(1) $\underset{\eta\in\operatorname{T}_{{\vec{\mathbf{W}}}}\!\mathcal{M}}{\arg\min}\,\|\mathrm{D}F({\vec{\mathbf{W}}})[\eta]+F({\vec{\mathbf{W}}})\|_{\mathrm{F}}^{2}.$
Specifically, it follows from the multilinearity of $\tau$ that the directional derivative $\mathrm{D}F({\vec{\mathbf{W}}})[\eta]$ in Eq. 1 can be computed by
$\mathrm{D}F({\vec{\mathbf{W}}})[\eta]$ $=\lim_{h\to 0}\frac{\mathrm{Proj}_{\Omega}(\tau({\vec{\mathbf{W}}}+h\eta)-\mathcal{A})-\mathrm{Proj}_{\Omega}(\tau({\vec{\mathbf{W}}})-\mathcal{A})}{\sqrt{p}h}$
$=\frac{1}{\sqrt{p}}\sum_{k=1}^{d}\mathrm{Proj}_{\Omega}(\tau(\mathbf{W}_{1},\dots,\mathbf{W}_{k-1},\eta_{k},\mathbf{W}_{k+1},\dots,\mathbf{W}_{d})).$
Then, we yield
$\|\mathrm{D}F({\vec{\mathbf{W}}})[\eta]+F({\vec{\mathbf{W}}})\|_{\mathrm{F}}^{2}$
$=\frac{1}{p}\sum_{i=1}^{n_{1}n_{2}\cdots n_{d}}\langle\mathrm{Proj}_{\Omega}(\mathcal{B}_{i}),\sum_{k=1}^{d}\tau(\mathbf{W}_{1},\dots,\mathbf{W}_{k-1},\eta_{k},\mathbf{W}_{k+1},\dots,\mathbf{W}_{d})+\tau({\vec{\mathbf{W}}})-\mathcal{A}\rangle^{2}$
$=\frac{1}{p}\sum_{i=1}^{n_{1}n_{2}\cdots n_{d}}\Big{(}\sum_{k=1}^{d}\left\langle\mathrm{Proj}_{\Omega_{(k)}}\!((\mathcal{B}_{i})_{(k)})\mathbf{W}_{\neq k},\eta_{k}\right\rangle+\left\langle\mathrm{Proj}_{\Omega}(\mathcal{B}_{i}),\tau({\vec{\mathbf{W}}})-\mathcal{A}\right\rangle\Big{)}^{2}$
$=\frac{1}{p}\sum_{(i_{1},\dots,i_{d})\in\Omega}\Big{(}\sum_{k=1}^{d}\eta_{k}(i_{k},:)^{\mathsf{T}}\mathrm{vec}((\prod_{j=k+1}^{d}\mathbf{U}_{j}(i_{j})\prod_{j=1}^{k-1}\mathbf{U}_{j}(i_{j}))^{\mathsf{T}})+\mathcal{S}(i_{1},\dots,i_{d})\Big{)}^{2},$
where $\{\mathcal{B}_{i}\}_{i=1}^{n_{1}n_{2}\cdots n_{d}}$ is defined by $(\mathcal{B}_{i})(i_{1},i_{2},\dots,i_{d})=1$ if $i=\sum_{j=1}^{d}(i_{j}-1)\prod_{\ell=1}^{j-1}n_{\ell}$, otherwise $(\mathcal{B}_{i})(i_{1},i_{2},\dots,i_{d})=0$, and $\mathcal{S}:=\mathrm{Proj}_{\Omega}(\tau({\vec{\mathbf{W}}})-\mathcal{A})$ refers to the residual tensor. Note that for $i=\sum_{j=1}^{d}(i_{j}-1)\prod_{\ell=1}^{j-1}n_{\ell}$, the matrix $\mathrm{Proj}_{\Omega_{(k)}}\!((\mathcal{B}_{i})_{(k)})=0$ if $(i_{1},i_{2},\dots,i_{d})\notin\Omega$. Consequently, the problem Eq. 1 is a least-squares problem of $\sum_{k=1}^{d}n_{k}r_{k-1}r_{k}$ variables.

### v3

@@
 Since optimization on Riemannian manifolds relies on the chosen metric, it is appealing to know that how the performance of a Riemannian optimization method varies with different metrics and how to exquisitely construct a metric such that a method can be accelerated.
-To this end, we propose a general framework for optimization problems on product manifolds endowed with a preconditioned metric, and we develop the Riemannian gradient descent and Riemannian conjugate gradient methods under this metric.
+To this end, we propose a general framework for optimization problems on product manifolds endowed with a preconditioned metric, and we develop Riemannian methods under this metric.
 Generally, the metric is constructed by an operator that aims to approximate the diagonal blocks of the Riemannian Hessian of the cost function.
 We propose three specific approaches to design the operator: exact block diagonal preconditioning, left and right preconditioning, and Gauss–Newton type preconditioning.
-In addition, we tailor new preconditioned metrics and adapt the proposed Riemannian methods to the canonical correlation analysis and the truncated singular value decomposition problems, and we adopt the Gauss–Newton type preconditioning to solve the tensor ring completion problem.
+Specifically, we tailor new preconditioned metrics and adapt the proposed Riemannian methods to the canonical correlation analysis and the truncated singular value decomposition problems, which provably accelerate the Riemannian methods.
+Additionally, we adopt the Gauss–Newton type preconditioning to solve the tensor ring completion problem.
 Numerical results among these applications verify that a delicate metric does accelerate the Riemannian optimization methods.
@@
 Related works and motivation
-Recently, Riemannian optimization, designing algorithms based on the geometry of a Riemannian manifold $\mathcal{M}$, appears to be prosperous in many areas.
-Specifically, one can propose Riemannian optimization methods to solve problem Eq.
+Riemannian optimization, designing algorithms based on the geometry of a Riemannian manifold $\mathcal{M}$, appears to be prosperous in many areas.
+One can propose Riemannian optimization methods to solve problem Eq.
 1, e.g., Riemannian gradient descent and Riemannian conjugate gradient methods.
@@
 Consider the problem
-$\min\ f(\mathbf{x}):=-\mathbf{b}^{\mathsf{T}}\mathbf{x},\quad\mathrm{s.\,t.}\ \mathbf{x}\in\mathcal{M}_{\mathbf{B}}:=\{\mathbf{x}\in\mathbb{R}^{n}:\ \mathbf{x}^{\mathsf{T}}\mathbf{B}\mathbf{x}=1\},$
+$\min\ f(\mathbf{x}):=-\mathbf{b}^{\mathsf{T}}\mathbf{x},\quad\mathrm{s.\,t.}\ \mathbf{x}\in\mathcal{M}_{\mathbf{B}}:=\{\mathbf{x}\in\mathbb{R}^{n}:\ \mathbf{x}^{\mathsf{T}}\mathbf{B}\mathbf{x}=1\},=$
 where $\mathbf{B}\in\mathbb{R}^{n\times n}$ is symmetric positive definite and $\mathbf{b}\in\mathbb{R}^{n}$.
@@
 We explore the effect of a family of metrics,
-$g_{\lambda,\mathbf{x}}:=\langle\xi,(\lambda\mathbf{I}_{n}+(1-\lambda)\mathbf{B})\eta\rangle\quad\text{for tangent vectors }\xi\text{ and }\eta,$
-to the Riemannian gradient descent (RGD) method and the condition number of $\operatorname{Hess}_{g_{\lambda,\mathbf{x}}}\!f(\mathbf{x}^{*})$ in  Fig.
+$g_{\lambda,\mathbf{x}}(\xi,\eta):=\langle\xi,(\lambda\mathbf{I}_{n}+(1-\lambda)\mathbf{B})\eta\rangle\quad\text{for tangent vectors }\xi\text{ and }\eta,$
+to the Riemannian gradient descent (RGD) method and the condition number of $\operatorname{Hess}_{g_{\lambda}}\!f(\mathbf{x}^{*})$ in  Fig.
 1, where $\lambda\in\mathbb{R}$ such that $\lambda\mathbf{I}_{n}+(1-\lambda)\mathbf{B}$ is positive definite.
-The left figure depicts the sequences generated by RGD under the Euclidean metric $g_{1,\mathbf{x}}(\xi,\eta)=\langle\xi,\eta\rangle$ and the scaled metric $g_{0,\mathbf{x}}(\xi,\eta)=\langle\xi,\mathbf{B}\eta\rangle$, and it shows that RGD under the metric $g_{0,\mathbf{x}}$ converges faster than the one under the Euclidean metric.
-Furthermore, the right figure confirms that the condition number varies with the metrics and $g_{0,\mathbf{x}}$ leads to the smallest condition number.
+The left figure depicts the sequences generated by RGD under the Euclidean metric $g_{1,\mathbf{x}}(\xi,\eta)=\langle\xi,\eta\rangle$ and the scaled metric $g_{0,\mathbf{x}}(\xi,\eta)=\langle\xi,\mathbf{B}\eta\rangle$, and it shows that RGD under the metric $g_{0}$ converges faster than the one under the Euclidean metric.
+Furthermore, the right figure confirms that the condition number varies with the metrics and $g_{0}$ leads to the smallest condition number.
 The detailed computation can be found in Appendix A.
 Figure 1: Left: sequences generated by the Riemannian gradient descent method under two metrics for $\mathbf{B}=\operatorname{diag}(2^{2},3^{2},1)$ and $\mathbf{b}=(1,1,1)$.
-Right: the condition number of $\operatorname{Hess}_{g_{\lambda,\mathbf{x}}}\!f(\mathbf{x}^{*})$ for $\lambda\in(-1/8,1]$.
+Right: the condition number of $\operatorname{Hess}_{g_{\lambda}}\!f(\mathbf{x}^{*})$ for $\lambda\in(-1/8,1]$.
 Blue marker: the Euclidean metric; green marker: the scaled metric.
@@
 We propose three types of preconditioning approaches.
-The first is exact block diagonal preconditioning.
+The first type is exact block diagonal preconditioning.
 Instead of approximating the full Riemannian Hessian, which can be computationally unfavorable in practice, we benefit from the block structure of $\operatorname{Hess}_{\mathrm{e}}\!f(x)$ and construct a new metric by exploiting the diagonal blocks; see an illustration in Fig.
@@
 $g_{x_{k}}^{k}(\xi_{k},\eta_{k})=\langle\xi_{k},\mathbf{M}_{k,1}(x)\eta_{k}\mathbf{M}_{k,2}(x)\rangle\quad\text{for }\xi_{k},\eta_{k}\in\operatorname{T}_{x_{k}}\!\mathcal{M}_{k};$
-see Section 3.2.
-The Riemannian metric on $\mathcal{M}$ for both diagonal types is defined by the sum of the Riemannian metric on each component, i.e., $g_{x}(\xi,\eta)=\sum_{k=1}^{K}g_{x_{k}}^{k}(\xi_{k},\eta_{k})$.
+see Section 3.2. The Riemannian metric on $\mathcal{M}$ for both diagonal types is defined by the sum of the Riemannian metric on each component, i.e.,
+$g_{x}(\xi,\eta)=g_{x_{1}}^{1}(\xi_{1},\eta_{1})+g_{x_{2}}^{2}(\xi_{2},\eta_{2})+\cdots+g_{x_{K}}^{K}(\xi_{K},\eta_{K}).$
 Thirdly, we propose Gauss–Newton type preconditioning technique for minimization of $\frac{1}{2}\|F(x)\|_{2}^{2}$ by constructing the metric
@@
 see Section 3.3.
-By virtue of the new metric, we propose the Riemannian gradient descent and Riemannian conjugate gradient methods, and the condition number-related convergence results are presented.
+By virtue of the new metric, we propose the Riemannian gradient descent and Riemannian conjugate gradient methods, and the condition number-related convergence results are developed.
 The preconditioned metric expands the scope of Riemannian preconditioning in [19] since it facilitates flexible choices of the operator ¯${\bar{\mathcal{H}}}(x)$.
@@
 The existing preconditioning methods that can be interpreted by the proposed framework are listed in Table 1.
-Table 1: Existing and our works interpreted by preconditioned metrics. MC: matrix completion; TC: tensor completion; CP: canonical polyadic; TT: tensor train; TR: tensor ring.
+Table 1: Existing and our works interpreted by preconditioned metrics.
+MC: matrix completion; TC: tensor completion; CP: canonical polyadic; TT: tensor train; TR: tensor ring; “*”: non-sigular matrices or tensors.
 Problem Methods Search space $\mathcal{M}$ Category
@@
 In this section, we provide basic tools in Riemannian geometry on product manifolds and develop the Riemannian gradient descent and Riemannian conjugate gradient methods for optimization on product manifolds; see [2, 4] for an overview.
-Metric-based and condition number-related convergence properties are presented.
+Metric-based and condition number-related convergence properties are developed.
 2.1 Riemannian optimization on product manifolds
@@
 The tangent space of $\mathcal{M}$ at $x=(x_{1},x_{2},\dots,x_{K})$ is denoted by $\operatorname{T}_{x}\!\mathcal{M}=\operatorname{T}_{x_{1}}\!\mathcal{M}_{1}\times\operatorname{T}_{x_{2}}\!\mathcal{M}_{2}\times\cdots\times\operatorname{T}_{x_{K}}\!\mathcal{M}_{K}$, and a tangent vector is denoted by $\eta=(\eta_{1},\eta_{2},\dots,\eta_{K})$.
-Let each manifold $\mathcal{M}_{k}$ be endowed with a Riemannian metric $g^{k}$, and hence the Riemannian metric on the product manifold $\mathcal{M}$ can be defined by
+Let each manifold $\mathcal{M}_{k}$ be endowed with a Riemannian metric $g^{k}$.
+The Riemannian metric on the product manifold $\mathcal{M}$ can be defined by
 $g_{x}(\xi,\eta):=g_{x_{1}}^{1}(\xi_{1},\eta_{1})+g_{x_{2}}^{2}(\xi_{2},\eta_{2})+\cdots+g_{x_{K}}^{K}(\xi_{K},\eta_{K})$
@@
 Recall that a Riemannian metric on the product manifold $\mathcal{M}$ is defined by the sum of the metrics on each component, i.e., $g_{x}(\xi,\eta)=\sum_{k=1}^{K}g_{x_{k}}^{k}(\xi_{k},\eta_{k})$ for $\xi=(\xi_{1},\xi_{2},\dots,\xi_{K}),\eta=(\eta_{1},\eta_{2},\dots,\eta_{K})\in\operatorname{T}_{x}\!\mathcal{M}$.
-Consider the smooth extension $\bar{H}_{kk}(x):\operatorname{T}_{x_{k}}\!\mathcal{E}_{k}\to\operatorname{T}_{x_{k}}\!\mathcal{E}_{k}$ of the diagonal block $H_{kk}$ for $k=1,\dots,K$.
+Consider the smooth extension $\bar{H}_{kk}(x):\operatorname{T}_{x_{k}}\!\mathcal{E}_{k}\to\operatorname{T}_{x_{k}}\!\mathcal{E}_{k}$ of the diagonal block $H_{kk}$ for $k=1,2,\dots,K$.
 If $\bar{H}_{11},\bar{H}_{22},\dots,\bar{H}_{KK}$ are positive definite on the ambient space $\mathcal{E}$, one can immediately adopt these blocks to construct the operator ${\bar{\mathcal{H}}}$ by
@@
 Therefore, it follows from Eq.
-4 that the Riemannian gradient reads
+4 that the Riemannian gradient is
 (5) $\mathrm{grad}_{g}f(\mathbf{U},\mathbf{V})=($ $-\Sigma_{xx}^{-1}\Sigma_{xy}\mathbf{V}\mathbf{N}+\mathbf{U}\operatorname{sym}(\mathbf{U}^{\mathsf{T}}\Sigma_{xy}\mathbf{V}\mathbf{N}),$
@@
 A proof is given in [25, Theorem 5].
-Nevertheless, we provide a slightly different proof for Proposition 4.1 in Appendix C for the sake of brevity, since the proofs of Propositions 4.4 and 5.1 stem from the details of the proof of Proposition 4.1.
+Nevertheless, we provide a slightly different proof for Proposition 4.1 in Appendix C to facilitate the proofs of Propositions 4.4 and 5.1.
+Specifically, the proof sketch of Propositions 4.1, 4.4, and 5.1 follows from the same procedure: 1) compute the Riemannian Hessian under the given metric $\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*})[\eta]$; 2) compute the Rayleigh quotient $q(\eta)$; 3) compute the maximum and minimum of Rayleigh quotient by taking the parametrization of tangent spaces into $q(\eta)$.
 Note that Proposition 4.1 boils down to $\kappa_{g}(\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))=(\sigma_{1}+\sigma_{2})/(\sigma_{1}-\sigma_{2})$ for $m=1$, which coincides with the result in [26, Lemma 4.1].
@@
 Then, the condition number at the local minimizer $(\mathbf{U}^{*},\mathbf{V}^{*})$ is computed by
-(10) $\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))=\frac{\max\{\underset{i,j\in[m],i\neq j}{\max}\frac{(\mu_{i}+\mu_{j})(\sigma_{i}+\sigma_{j})}{\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta}},\frac{\mu_{1}(\sigma_{1}+\sigma_{m+1})}{\sqrt{\mu_{1}^{2}\sigma_{1}^{2}+\delta}}\}}{\min\{\underset{i,j\in[m],i\neq j}{\min}\frac{(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j})}{\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta}},\frac{\mu_{m}(\sigma_{m}-\sigma_{m+1})}{\sqrt{\mu_{m}^{2}\sigma_{m}^{2}+\delta}}\}}.$
+(10) $\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))=\frac{\max\{\underset{i,j\in[m],i\neq j}{\max}\frac{(\mu_{i}+\mu_{j})(\sigma_{i}+\sigma_{j})}{\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta}},\underset{i\in[m]}{\max}\frac{\mu_{i}(\sigma_{i}+\sigma_{m+1})}{\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}}\}}{\min\{\underset{i,j\in[m],i\neq j}{\min}\frac{(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j})}{\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta}},\underset{i\in[m]}{\min}\frac{\mu_{i}(\sigma_{i}-\sigma_{m+1})}{\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}}\}}.$
 Proof 4.5.
-We aim to compute the Rayleigh quotient Eq.
+By following the same procedure as Proposition 4.1, we compute the Rayleigh quotient Eq.
 4 under the proposed metric Eq.
 7, and evaluate its upper and lower bounds.
-To this end, we firstly simplify the Riemannian Hessian $\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*})[\eta]$ as follows.
+To this end, we firstly compute and simplify the Riemannian Hessian $\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*})[\eta]$ as follows.
 By solving the Lyapunov equations in Proposition 4.2, we obtain that $\mathbf{S}_{1}=\mathbf{S}_{2}=-\Sigma\mathbf{N}$, where $\Sigma=\operatorname{diag}(\sigma_{1},\sigma_{2},\dots,\sigma_{m})$.
@@
 $\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*})[\eta]$ $=\Pi_{\mathrm{new},(\mathbf{U}^{*},\mathbf{V}^{*})}(\mathrm{D}\bar{G}_{\mathrm{new}}(\mathbf{U}^{*},\mathbf{V}^{*})[\eta])$
-$=\Pi_{\mathrm{new},(\mathbf{U}^{*},\mathbf{V}^{*})}($ $-\Sigma_{xx}^{-1}\Sigma_{xy}\eta_{2}\mathbf{N}(\mathbf{M}_{1,2}^{*})^{-1}+\eta_{1}\Sigma\mathbf{N}(\mathbf{M}_{1,2}^{*})^{-1}-\mathbf{U}^{*}\dot{\mathbf{S}}_{1}(\mathbf{M}_{1,2}^{*})^{-1},$
-$-\Sigma_{yy}^{-1}\Sigma_{xy}^{\mathsf{T}}\eta_{1}\mathbf{N}(\mathbf{M}_{2,2}^{*})^{-1}+\eta_{2}\Sigma\mathbf{N}(\mathbf{M}_{2,2}^{*})^{-1}-\mathbf{V}^{*}\dot{\mathbf{S}}_{2}(\mathbf{M}_{2,2}^{*})^{-1}),$
-where $\mathbf{M}_{1,2}^{*}=\mathbf{M}_{2,2}^{*}=(\Sigma^{2}\mathbf{N}^{2}+\delta\mathbf{I}_{m})^{1/2}$.
+$=\Pi_{\mathrm{new},(\mathbf{U}^{*},\mathbf{V}^{*})}(-\Sigma_{xx}^{-1}\Sigma_{xy}\eta_{2}\mathbf{N}(\mathbf{M}_{1,2}^{*})^{-1}+\eta_{1}\Sigma\mathbf{N}(\mathbf{M}_{1,2}^{*})^{-1},$
+$\qquad\qquad\qquad~{}~{}~{}-\!\Sigma_{yy}^{-1}\Sigma_{xy}^{\mathsf{T}}\eta_{1}\mathbf{N}(\mathbf{M}_{2,2}^{*})^{-1}+\eta_{2}\Sigma\mathbf{N}(\mathbf{M}_{2,2}^{*})^{-1}),$
+where $\mathbf{M}_{1,2}^{*}=\mathbf{M}_{2,2}^{*}=(\Sigma^{2}\mathbf{N}^{2}+\delta\mathbf{I}_{m})^{1/2}$ are diagonal matrices, and we use the characterization of $(\operatorname{T}_{\mathbf{U}^{*}}\!\mathcal{M}_{1})^{\perp}$ and $(\operatorname{T}_{\mathbf{V}^{*}}\!\mathcal{M}_{2})^{\perp}$ in Eq.
+1.
 Subsequently, we compute the Rayleigh quotient. It follows from Eq. 4 that
@@
 4.
+We observe that only the denominator of $q(\eta)$ is different from the denominator $\langle\eta_{1},\Sigma_{xx}\eta_{1}\rangle+\langle\eta_{2},\Sigma_{yy}\eta_{2}\rangle$ in Eq.
+2, and $\mathbf{M}_{1,2}^{*},\mathbf{M}_{2,2}^{*}$ are diagonal matrices.
 Consequently, we can evaluate the upper and lower bounds of $q(\eta)$ in a similar fashion as Appendix C.
@@
 To this end, we first provide Lemma 4.6 to simplify Eq.
-9.
+10.
 Next, we prove that $\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))\leq\kappa_{g}(\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$ in Proposition 4.8.
 Lemma 4.6.
-Denote $\bar{v}_{ij}(\delta):=(\mu_{i}+\mu_{j})(\sigma_{i}+\sigma_{j})/(\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta})$ and $\underline{v}_{ij}(\delta):=(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j})/(\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta})$.
+Denote 1) $\bar{v}_{ij}(\delta):=(\mu_{i}+\mu_{j})(\sigma_{i}+\sigma_{j})/(\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta})$ and $\bar{v}_{i,m+1}(\delta)=\mu_{i}(\sigma_{i}+\sigma_{m+1})/\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}$; 2) $\underline{v}_{ij}(\delta):=(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j})/(\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta})$ and $\underline{v}_{i,m+1}(\delta):=\mu_{i}(\sigma_{i}-\sigma_{m+1})/\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}$ for $i,j\in[m]$.
 It holds that
-$\bar{v}_{ij}(0)>\bar{v}_{ik}(0)>\bar{v}_{jk}(0)\qquad\text{and}\qquad\underline{v}_{ij}(0)<\underline{v}_{ik}(0)<\underline{v}_{jk}(0)$
-for all $1\leq i<j<k\leq m$ with $m\geq 3$.
+$\bar{v}_{ij}(0)>\bar{v}_{ik}(0)\qquad\text{and}\qquad\underline{v}_{ij}(0)<\underline{v}_{ik}(0)$
+for all $11\leq i<j<k\leq m+$ 1 with $m\geq 3$.
 Proof 4.7.
-First, we prove that $\bar{v}_{ij}(0)>\bar{v}_{ik}(0)$. Since $\mu_{i}>\mu_{j}>\mu_{k}$ and $\sigma_{i}>\sigma_{j}>\sigma_{k}$, it holds that
+We observe that $\bar{v}_{ij}(0)=(\mu_{i}+\mu_{j})(\sigma_{i}+\sigma_{j})/(\mu_{i}\sigma_{i}+\mu_{j}\sigma_{j})$ and $\underline{v}_{ij}(0)=(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j})/(\mu_{i}\sigma_{i}+\mu_{j}\sigma_{j})$ for $i\in[m]$ and $j\in[m+1]$ with $\mu_{m+1}=0$.
+First, we prove that $\bar{v}_{ij}(0)>\bar{v}_{ik}(0)$.
+Since $\mu_{i}>\mu_{j}>\mu_{k}$ and $\sigma_{i}>\sigma_{j}>\sigma_{k}$, it holds that
 $\bar{v}_{ij}(0)-\bar{v}_{ik}(0)$ $=\frac{(\mu_{i}\sigma_{j}+\mu_{j}\sigma_{i})(\mu_{i}\sigma_{i}+\mu_{k}\sigma_{k})-(\mu_{i}\sigma_{k}+\mu_{k}\sigma_{i})(\mu_{i}\sigma_{i}+\mu_{j}\sigma_{j})}{(\mu_{i}\sigma_{i}+\mu_{j}\sigma_{j})(\mu_{i}\sigma_{i}+\mu_{k}\sigma_{k})}$
 $=\frac{(\mu_{i}^{2}-\mu_{j}\mu_{k})\sigma_{i}(\sigma_{j}-\sigma_{k})+\mu_{i}(\mu_{j}-\mu_{k})(\sigma_{i}^{2}-\sigma_{j}\sigma_{k})}{(\mu_{i}\sigma_{i}+\mu_{j}\sigma_{j})(\mu_{i}\sigma_{i}+\mu_{k}\sigma_{k})}>0.$
-Therefore, $\bar{v}_{ij}(0)>\bar{v}_{ik}(0)$ holds and $\bar{v}_{ik}(0)>\bar{v}_{jk}(0)$ is also true.
-The result $\underline{v}_{ij}(0)<\underline{v}_{ik}(0)<\underline{v}_{jk}(0)$ is ready by using $\bar{v}_{ij}(0)+\underline{v}_{ij}(0)=2$, $\bar{v}_{ik}(0)+\underline{v}_{ik}(0)=2$ and $\bar{v}_{jk}(0)+\underline{v}_{jk}(0)=2$.
-Then, we can prove that $\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))\leq\kappa_{g}(\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$ for $m\geq 2$.
+Therefore, $\bar{v}_{ij}(0)>\bar{v}_{ik}(0)$ holds.
+The results $\underline{v}_{ij}(0)<\underline{v}_{ik}(0)$ is ready by using $\bar{v}_{ij}(0)+\underline{v}_{ij}(0)=2$ and $\bar{v}_{ik}(0)+\underline{v}_{ik}(0)=2$.
+Then, it follows from Lemma 4.6 and the continuity of $\bar{v}_{ij}$ and $\underline{v}_{ij}$ with respect to $\delta\in[0,\infty)$ that there exists a constant $\bar{\delta}_{1}>0$, such that
+$\bar{v}_{ij}(\delta)>\bar{v}_{ik}(\delta)\qquad\text{and}\qquad\underline{v}_{ij}(\delta)<\underline{v}_{ik}(\delta)$
+hold for all $11\leq i<j<k\leq m+$ 1 and $\delta\in(0,\bar{\delta}_{1})$. Therefore, we can simplify the condition number
+(11) $\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$ $=\frac{\max_{i\in[m],j\in[m+1],i\neq j}\bar{v}_{ij}(\delta)}{\min_{i\in[m],j\in[m+1],i\neq j}\underline{v}_{ij}(\delta)}=\frac{\max_{i\in[m]}\bar{v}_{i,i+1}(\delta)}{\min_{i\in[m]}\underline{v}_{i,i+1}(\delta)}.$
+We aim to prove that $\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))\leq\kappa_{g}(\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$ for $m\geq 2$.
 Note that $\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))=\kappa_{g}(\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$ for $m=1$ since the right preconditioners in Eq.
@@
 Proof 4.9.
-First, it follows from Lemma 4.6 and the continuity of $\bar{v}_{ij}$ and $\underline{v}_{ij}$ with respect to $\delta\in[0,\infty)$ that there exists a constant $\bar{\delta}>0$, such that
-$\bar{v}_{ij}(\delta)>\bar{v}_{ik}(\delta)>\bar{v}_{jk}(\delta)\qquad\text{and}\qquad\underline{v}_{ij}(\delta)<\underline{v}_{ik}(\delta)<\underline{v}_{jk}(\delta)$
-hold for all $1\leq i<j<k\leq m$ and $\delta\in(0,\bar{\delta})$.
-Therefore, we have $\bar{v}_{12}(\delta)=\max_{i,j\in[m],i\neq j}\bar{v}_{ij}(\delta)$ and $\underline{v}_{12}(\delta)=\min_{i,j\in[m],i\neq j}\underline{v}_{ij}(\delta)$.
-Subsequently, we observe that
-$~{}~{}~{}~{}\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$
-$=\frac{\max\{\underset{i,j\in[m],i\neq j}{\max}\bar{v}_{ij}(\delta),\frac{\mu_{1}(\sigma_{1}+\sigma_{m+1})}{\sqrt{\mu_{1}^{2}\sigma_{1}^{2}+\delta}}\}}{\min\{\underset{i,j\in[m],i\neq j}{\min}\underline{v}_{ij}(\delta),\frac{\mu_{m}(\sigma_{m}-\sigma_{m+1})}{\sqrt{\mu_{m}^{2}\sigma_{m}^{2}+\delta}}\}}$
-$=\frac{\max\{\bar{v}_{12}(\delta),\frac{\mu_{1}(\sigma_{1}+\sigma_{m+1})}{\sqrt{\mu_{1}^{2}\sigma_{1}^{2}+\delta}}\}}{\min\{\underline{v}_{12}(\delta),\frac{\mu_{m}(\sigma_{m}-\sigma_{m+1})}{\sqrt{\mu_{m}^{2}\sigma_{m}^{2}+\delta}}\}}$
-(11) $=\frac{\max\{\frac{1}{2}(\mu_{1}+\mu_{2})(\sigma_{1}+\sigma_{2}),\frac{\mu_{1}(\sigma_{1}+\sigma_{m+1})}{2\sqrt{\mu_{1}^{2}\sigma_{1}^{2}+\delta}}(\sqrt{\mu_{1}^{2}\sigma_{1}^{2}+\delta}+\sqrt{\mu_{2}^{2}\sigma_{2}^{2}+\delta})\}}{\min\{\frac{1}{2}(\mu_{1}-\mu_{2})(\sigma_{1}-\sigma_{2}),\frac{\mu_{m}(\sigma_{m}-\sigma_{m+1})}{2\sqrt{\mu_{m}^{2}\sigma_{m}^{2}+\delta}}(\sqrt{\mu_{1}^{2}\sigma_{1}^{2}+\delta}+\sqrt{\mu_{2}^{2}\sigma_{2}^{2}+\delta})\}}$
-(12) $\leq\frac{\max\{\frac{1}{2}(\mu_{1}+\mu_{2})(\sigma_{1}+\sigma_{2}),\mu_{1}(\sigma_{1}+\sigma_{m+1})\}}{\min\{\underset{i,j\in[m],i\neq j}{\min}\frac{1}{2}(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j}),\mu_{m}(\sigma_{m}-\sigma_{m+1})\}}$
-$=\kappa_{g}(\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*})),$
-where we multiply both the numerator and the denominator by $\frac{1}{2}(\sqrt{\mu_{1}^{2}\sigma_{1}^{2}+\delta}+\sqrt{\mu_{2}^{2}\sigma_{2}^{2}+\delta})$ to yield Eq.
-11, and we use the fact that $2\sqrt{\mu_{m}^{2}\sigma_{m}^{2}+\delta}<\sqrt{\mu_{1}^{2}\sigma_{1}^{2}+\delta}+\sqrt{\mu_{2}^{2}\sigma_{2}^{2}+\delta}<2\sqrt{\mu_{1}^{2}\sigma_{1}^{2}+\delta}$ to yield Eq.
-12.
+If $\operatorname{arg\,max}_{i\in[m]}\bar{v}_{i,i+1}(0)=\{i^{*}\}$ for some $i^{*}\in[m]$, it follows from $\bar{v}_{i,i+1}(0)+\underline{v}_{i,i+1}(0)=2$ that $\operatorname{arg\,min}_{i\in[m]}\underline{v}_{i,i+1}(0)=\operatorname{arg\,max}_{i\in[m]}\bar{v}_{i,i+1}(0)=\{i^{*}\}$.
+Since $\bar{v}_{ij}(\delta)$ and $\underline{v}_{ij}(\delta)$ are continuous, there exists $\bar{\delta}>0$, such that $\{i^{*}\}=\operatorname{arg\,max}_{i\in[m]}\bar{v}_{i,i+1}(\delta)$ and $\{i^{*}\}=\operatorname{arg\,min}_{i\in[m]}\underline{v}_{i,i+1}(\delta)$ for all $\delta\in[0,\bar{\delta})$.
+Subsequently, we obtain that
+$\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$ $=\frac{\bar{v}_{i^{*},i^{*}+1}(\delta)}{\underline{v}_{i^{*},i^{*}+1}(\delta)}=\frac{\frac{1}{2}(\mu_{i^{*}}+\mu_{i^{*}+1})(\sigma_{i^{*}}+\sigma_{i^{*}+1})}{\frac{1}{2}(\mu_{i^{*}}-\mu_{i^{*}+1})(\sigma_{i^{*}}-\sigma_{i^{*}+1})}$
+$\leq\frac{\max\{\frac{1}{2}(\mu_{1}+\mu_{2})(\sigma_{1}+\sigma_{2}),\mu_{1}(\sigma_{1}+\sigma_{m+1})\}}{\min\{\underset{i,j\in[m],i\neq j}{\min}\frac{1}{2}(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j}),\mu_{m}(\sigma_{m}-\sigma_{m+1})\}}$
+$=\kappa_{g}(\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*})).$
+If $\{i_{1}^{*},i_{2}^{*}\}\subseteq\operatorname{arg\,max}_{i\in[m]}\bar{v}_{i,i+1}(0)$ for $i_{1}^{*}<i_{2}^{*}$, it follows from the continuity of $\bar{v}_{ij}$ and $\underline{v}_{ij}$ that there exists $\bar{\delta}_{i_{1}^{*},i_{2}^{*}}>0$, such that $\bar{v}_{i_{1}^{*},i_{1}^{*}+1}(\delta)>\bar{v}_{i_{2}^{*},i_{2}^{*}+1}(\delta)$ and $\underline{v}_{i_{1}^{*},i_{1}^{*}+1}(\delta)>\underline{v}_{i_{2}^{*},i_{2}^{*}+1}(\delta)$.
+The rationales behind these are $\bar{v}_{i_{1}^{*},i_{1}^{*}+1}(0)=\bar{v}_{i_{2}^{*},i_{2}^{*}+1}(0)$, $\underline{v}_{i_{1}^{*},i_{1}^{*}+1}(0)=\underline{v}_{i_{2}^{*},i_{2}^{*}+1}(0)$, and the derivatives satisfy
+$\bar{v}_{i_{1}^{*},i_{1}^{*}+1}^{\prime}(0)$ $=-\frac{\bar{v}_{i_{1}^{*},i_{1}^{*}+1}(0)}{2\mu_{i_{1}^{*}}\sigma_{i_{1}^{*}}\mu_{i_{1}^{*}+1}\sigma_{i_{1}^{*}+1}}>-\frac{\bar{v}_{i_{2}^{*},i_{2}^{*}+1}(0)}{2\mu_{i_{2}^{*}}\sigma_{i_{2}^{*}}\mu_{i_{2}^{*}+1}\sigma_{i_{2}^{*}+1}}=\bar{v}_{i_{2}^{*},i_{2}^{*}+1}^{\prime}(0),$
+$\underline{v}_{i_{1}^{*},i_{1}^{*}+1}^{\prime}(0)$ $=-\frac{\underline{v}_{i_{1}^{*},i_{1}^{*}+1}(0)}{2\mu_{i_{1}^{*}}\sigma_{i_{1}^{*}}\mu_{i_{1}^{*}+1}\sigma_{i_{1}^{*}+1}}>-\frac{\underline{v}_{i_{2}^{*},i_{2}^{*}+1}(0)}{2\mu_{i_{2}^{*}}\sigma_{i_{2}^{*}}\mu_{i_{2}^{*}+1}\sigma_{i_{2}^{*}+1}}=\underline{v}_{i_{2}^{*},i_{2}^{*}+1}^{\prime}(0)$
+if $i_{2}^{*}<m$ and
+$\bar{v}_{i_{1}^{*},i_{1}^{*}+1}^{\prime}(0)$ $=-\frac{\bar{v}_{i_{1}^{*},i_{1}^{*}+1}(0)}{2\mu_{i_{1}^{*}}\sigma_{i_{1}^{*}}\mu_{i_{1}^{*}+1}\sigma_{i_{1}^{*}+1}}>-\frac{\bar{v}_{m,m+1}(0)}{2\mu_{m}^{2}\sigma_{m}^{2}}=\bar{v}_{m,m+1}^{\prime}(0),$
+$\underline{v}_{i_{1}^{*},i_{1}^{*}+1}^{\prime}(0)$ $=-\frac{\underline{v}_{i_{1}^{*},i_{1}^{*}+1}(0)}{2\mu_{i_{1}^{*}}\sigma_{i_{1}^{*}}\mu_{i_{1}^{*}+1}\sigma_{i_{1}^{*}+1}}>-\frac{\underline{v}_{m,m+1}(0)}{2\mu_{m}^{2}\sigma_{m}^{2}}=\underline{v}_{m,m+1}^{\prime}(0)$
+if $i_{2}^{*}=m$.
+Therefore, there exists $\bar{\delta}\in(0,\min\{\bar{\delta}_{i,j}:i,j\in\operatorname{arg\,max}_{i\in[m]}\bar{v}_{i,i+1}(0)\})$ and $i^{*},j^{*}\in[m]$, such that: 1) $i^{*}=\operatorname{arg\,max}_{i\in[m]}\bar{v}_{i,i+1}(\delta)$; 2) $j^{*}=\operatorname{arg\,min}_{i\in[m]}\underline{v}_{i,i+1}(\delta)$ for all $\delta\in[0,\bar{\delta})$; 3) $i^{*}<j^{*}$.
+Consequently, we obtain that
+$~{}~{}~{}~{}\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))=\frac{\bar{v}_{i^{*},i^{*}+1}(\delta)}{\underline{v}_{j^{*},j^{*}+1}(\delta)}$
+$=\frac{(\mu_{i^{*}}+\mu_{i^{*}+1})(\sigma_{i^{*}}+\sigma_{i^{*}+1})}{(\mu_{j^{*}}-\mu_{j^{*}+1})(\sigma_{j^{*}}-\sigma_{j^{*}+1})}\cdot\frac{\sqrt{\mu_{j^{*}}^{2}\sigma_{j^{*}}^{2}+1}+\sqrt{\mu_{j^{*}+1}^{2}\sigma_{j^{*}+1}^{2}+1}}{\sqrt{\mu_{i^{*}}^{2}\sigma_{i^{*}}^{2}+1}+\sqrt{\mu_{i^{*}+1}^{2}\sigma_{i^{*}+1}^{2}+1}}$
+$<\frac{\max\{\frac{1}{2}(\mu_{1}+\mu_{2})(\sigma_{1}+\sigma_{2}),\mu_{1}(\sigma_{1}+\sigma_{m+1})\}}{\min\{\underset{i,j\in[m],i\neq j}{\min}\frac{1}{2}(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j}),\mu_{m}(\sigma_{m}-\sigma_{m+1})\}}$
+$=\kappa_{g}(\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*})).$
 It is worth noting that the parameter $\delta>0$ theoretically ensures that Eq. 7 is a Riemannian metric. In practice, one can choose a sufficiently small $\delta$, e.g., $\delta=10^{-15}$.
@@
 5:  $(\mathbf{U}^{(t)},\mathbf{V}^{(t)})\in\mathcal{M}$.
-Note that 1) the retraction mapping is the generalized QR decomposition [22] with respect to $\Sigma_{xx}$ and $\Sigma_{yy}$, i.e.,
+Note that 1) the retraction mapping is the generalized QR factorization [22] with respect to $\Sigma_{xx}$ and $\Sigma_{yy}$, i.e.,
 $\operatorname{R}_{(\mathbf{U},\mathbf{V})}(\eta):=(\Sigma_{xx}^{-\frac{1}{2}}\operatorname{qf}(\Sigma_{xx}^{\frac{1}{2}}(\mathbf{U}+\eta_{1})),\Sigma_{yy}^{-\frac{1}{2}}\operatorname{qf}(\Sigma_{yy}^{\frac{1}{2}}(\mathbf{V}+\eta_{2})))\quad\text{for }\eta\in\operatorname{T}_{(\mathbf{U},\mathbf{V})}\!\mathcal{M},$
-where $\operatorname{qf}(\mathbf{X})$ refers to the $\mathbf{Q}$ factor in the QR decomposition $\mathbf{Q}\mathbf{R}=\mathbf{X}$.
-In practice, the retraction can be efficiently computed [22] by $\operatorname{R}_{(\mathbf{U},\mathbf{V})}(\eta)=((\mathbf{U}+\eta_{1})\mathbf{R}_{1}^{-1},(\mathbf{V}+\eta_{2})\mathbf{R}_{2}^{-1})$ instead, where $\mathbf{R}_{1}^{\mathsf{T}}\mathbf{R}_{1}=(\mathbf{U}+\eta_{1})^{\mathsf{T}}\Sigma_{xx}(\mathbf{U}+\eta_{1})$ and $\mathbf{R}_{2}^{\mathsf{T}}\mathbf{R}_{2}=(\mathbf{V}+\eta_{2})^{\mathsf{T}}\Sigma_{yy}(\mathbf{V}+\eta_{2})$ are Cholesky decompositions; 2) the vector transport in Algorithm 4 is defined by the projection operator in Eq.
+where $\operatorname{qf}(\mathbf{X})$ refers to the $\mathbf{Q}$ factor in the QR factorization $\mathbf{Q}\mathbf{R}=\mathbf{X}$.
+In practice, the retraction can be efficiently computed [22] by $\operatorname{R}_{(\mathbf{U},\mathbf{V})}(\eta)=((\mathbf{U}+\eta_{1})\mathbf{R}_{1}^{-1},(\mathbf{V}+\eta_{2})\mathbf{R}_{2}^{-1})$ instead, where $\mathbf{R}_{1}^{\mathsf{T}}\mathbf{R}_{1}=(\mathbf{U}+\eta_{1})^{\mathsf{T}}\Sigma_{xx}(\mathbf{U}+\eta_{1})$ and $\mathbf{R}_{2}^{\mathsf{T}}\mathbf{R}_{2}=(\mathbf{V}+\eta_{2})^{\mathsf{T}}\Sigma_{yy}(\mathbf{V}+\eta_{2})$ are Cholesky factorization; 2) the vector transport in Algorithm 4 is defined by the projection operator in Eq.
 8, i.e., $\mathcal{T}_{t\leftarrow t-1}(\eta)=\Pi_{g,(\mathbf{U}^{(t)},\mathbf{V}^{(t)})}(\eta)$ for $\eta\in\operatorname{T}_{(\mathbf{U}^{(t-1)},\mathbf{V}^{(t-1)})}\!\mathcal{M}$.
@@
 We observe that the Riemannian Hessian under the proposed metric (LR12) has the smallest condition number among all choices, which is reflected in the observation that RGD(LR12) and RCG(LR12) outperform the others.
-5 Appplication to truncated singular value decomposition
+5 Application to truncated singular value decomposition
 In this section, the truncated singular value decomposition (SVD) problem is considered.
@@
 4 is comparable to one under the Euclidean metric since $\mathbf{M}_{1,2},\mathbf{M}_{2,2}\in\mathbb{R}^{p\times p}$ and $p\ll\min\{m,n\}$.
-The effect of the new metric Eq. 2 is illustrated by the following proposition, which can be proved in a same fashion as Propositions 4.1, 4.4, and 4.8.
+The effect of the new metric Eq.
+2 is illustrated by the following proposition, which can be proved in a same fashion by letting $\Sigma_{xx}=\mathbf{I}_{d_{x}}$, $\Sigma_{yy}=\mathbf{I}_{d_{y}}$, $\Sigma_{xy}=\mathbf{A}$, $d_{x}=m$ and $d_{y}=n$ in Propositions 4.1 and 4.4.
 Proposition 5.1.
@@
 $\kappa_{\mathrm{e}}(\operatorname{Hess}_{\mathrm{e}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$ $=\frac{\max\left\{\frac{1}{2}(\mu_{1}+\mu_{2})(\sigma_{1}+\sigma_{2}),\mu_{1}(\sigma_{1}+\sigma_{p+1})\right\}}{\min\{\min_{i,j\in[p],i\neq j}\frac{1}{2}(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j}),\mu_{p}(\sigma_{p}-\sigma_{p+1})\}},$
-$\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$ $=\frac{\max\{\underset{i,j\in[p],i\neq j}{\max}\frac{(\mu_{i}+\mu_{j})(\sigma_{i}+\sigma_{j})}{\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta}},\frac{\mu_{1}(\sigma_{1}+\sigma_{p+1})}{\sqrt{\sigma_{1}^{2}\mu_{1}^{2}+\delta}}\}}{\min\{\underset{i,j\in[p],i\neq j}{\min}\frac{(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j})}{\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta}},\frac{\mu_{p}(\sigma_{p}-\sigma_{p+1})}{\sqrt{\sigma_{p}^{2}\mu_{p}^{2}+\delta}}\}}.$
-Moreover, the new metric Eq.
-2 indeed improves the condition number of the Riemannian Hessian, i.e., $\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))\leq\kappa_{\mathrm{e}}(\operatorname{Hess}_{\mathrm{e}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$.
+$\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$ $=\frac{\max\{\underset{i,j\in[p],i\neq j}{\max}\frac{(\mu_{i}+\mu_{j})(\sigma_{i}+\sigma_{j})}{\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta}},\underset{i\in[p]}{\max}\frac{\mu_{i}(\sigma_{i}+\sigma_{p+1})}{\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}}\}}{\min\{\underset{i,j\in[p],i\neq j}{\min}\frac{(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j})}{\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta}},\underset{i\in[p]}{\min}\frac{\mu_{i}(\sigma_{i}-\sigma_{p+1})}{\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}}\}}.$
+Moreover, the new metric Eq. 2 indeed improves the condition number of the Riemannian Hessian, i.e.,
+$\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))\leq\kappa_{\mathrm{e}}(\operatorname{Hess}_{\mathrm{e}}\!f(\mathbf{U}^{*},\mathbf{V}^{*})).$
 5.2 RGD and RCG for truncated singular value decomposition
@@
 1 in Algorithms 5 and 6.
-Note that the retraction mapping is based on the QR decomposition, i.e., $\operatorname{R}_{(\mathbf{U},\mathbf{V})}(\eta):=(\operatorname{qf}(\mathbf{U}+\eta_{1}),\operatorname{qf}(\mathbf{V}+\eta_{2}))$ for $\eta\in\operatorname{T}_{(\mathbf{U},\mathbf{V})}\!\mathcal{M}$.
+Note that the retraction mapping is based on the QR factorization, i.e., $\operatorname{R}_{(\mathbf{U},\mathbf{V})}(\eta):=(\operatorname{qf}(\mathbf{U}+\eta_{1}),\operatorname{qf}(\mathbf{V}+\eta_{2}))$ for $\eta\in\operatorname{T}_{(\mathbf{U},\mathbf{V})}\!\mathcal{M}$.
 The vector transport in Algorithm 6 is defined by the projection operator Eq.
@@
 The matrix $\mathbf{A}$ is constructed by $\mathbf{A}=\mathbf{U}^{*}\Sigma(\mathbf{V}^{*})^{\mathsf{T}}$, where the entries of $\mathbf{U}^{*}\in\mathbb{R}^{m\times p}$ and $\mathbf{V}^{*}\in\mathbb{R}^{n\times p}$ are firstly sampled i.i.d.
-from the uniform distribution on $[0,1]$, and $\mathbf{U}^{*}$ and $\mathbf{V}^{*}$ are orthogonalized by QR decomposition.
+from the uniform distribution on $[0,1]$, and $\mathbf{U}^{*}$ and $\mathbf{V}^{*}$ are orthogonalized by QR factorization.
 We set $\Sigma:=\operatorname{diag}(1,\gamma,\gamma^{2},\dots,\gamma^{p-1})$ and $\gamma=1/1.5$.
@@
 We set $d=3$, $n_{1}=n_{2}=n_{3}=100$, $p=0.05$, TR ranks $\mathbf{r}^{*}=(1,1,1),(2,2,2),\dots,(8,8,8)$, and $\delta=10^{-15}$.
-Figure 7: Training and test errors for TR rank $\mathbf{r}^{*}=(5,5,5)$. Each method is tested for 10 runs
 We specify the default settings of all methods.
@@
 A method is terminated if one of the following stopping criteria is achieved: 1) training error $\varepsilon_{\Omega}({\vec{\mathbf{W}}}^{(t)})<10^{-14}$; 2) the maximum iteration $1000$; 3) the relative change $\lvert{(\varepsilon_{\Omega}({\vec{\mathbf{W}}}^{(t)})-\varepsilon_{\Omega}({\vec{\mathbf{W}}}^{(t-1)}))}/{\varepsilon_{\Omega}({\vec{\mathbf{W}}}^{(t-1)})}\rvert<\varepsilon$; 4) the stepsize $s^{(t)}<10^{-10}$.
-Figure 8: Computation time required for each method to reach the stopping criteria for TR rank $\mathbf{r}^{*}=(1,1,1),(2,2,2),\dots,(8,8,8)$
 Numerical results are illustrated in Figs.
@@
 In other words, there is a trade-off between exploiting second-order information and the computational efficiency.
+Figure 7: Training and test errors for TR rank $\mathbf{r}^{*}=(5,5,5)$. Each method is tested for 10 runs
+Figure 8: Computation time required for each method to reach the stopping criteria for TR rank $\mathbf{r}^{*}=(1,1,1),(2,2,2),\dots,(8,8,8)$
 7 Conclusions and future works
@@
 Numerical results verify that a delicate metric does improve the performance of the Riemannian optimization methods.
-In the future, we intend to specify the proposed framework to other problems, e.g., the distributed PCA.
+In the future, we intend to specify the proposed framework to other problems.
 Moreover, taking into account the block structure of product manifolds, parallel computing is capable of further accelerating the Riemannian optimization methods.
@@
 In Example 1.1, we investigate a class of preconditioned metrics defined by
-$g_{\lambda,\mathbf{x}}=\langle\xi,\mathbf{B}_{\lambda}\eta\rangle\quad\text{for }\xi,\eta\in\operatorname{T}_{\mathbf{x}}\!\mathcal{M}_{\mathbf{B}},$
+$g_{\lambda,\mathbf{x}}(\xi,\eta)=\langle\xi,\mathbf{B}_{\lambda}\eta\rangle\quad\text{for }\xi,\eta\in\operatorname{T}_{\mathbf{x}}\!\mathcal{M}_{\mathbf{B}},$
 where $\lambda\in\mathbb{R}$ such that the matrix $\mathbf{B}_{\lambda}:=\lambda\mathbf{I}_{n}+(1-\lambda)\mathbf{B}$ is positive definite.
 The Riemannian gradients at $x\in\mathcal{M}_{\mathbf{B}}$ under these metrics are
-$\mathrm{grad}_{g_{\lambda,\mathbf{x}}}f(x)=-(\mathbf{I}_{n}-\frac{\mathbf{B}_{\lambda}^{-1}\mathbf{B}\mathbf{x}\mathbf{x}^{\mathsf{T}}\mathbf{B}}{\mathbf{x}^{\mathsf{T}}\mathbf{B}\mathbf{B}_{\lambda}^{-1}\mathbf{B}\mathbf{x}})\mathbf{B}_{\lambda}^{-1}\mathbf{b}=-\mathbf{B}_{\lambda}^{-1}\mathbf{b}+\frac{\mathbf{x}^{\mathsf{T}}\mathbf{B}\mathbf{B}_{\lambda}^{-1}\mathbf{b}}{\mathbf{x}^{\mathsf{T}}\mathbf{B}\mathbf{B}_{\lambda}^{-1}\mathbf{B}\mathbf{x}}\mathbf{B}_{\lambda}^{-1}\mathbf{B}\mathbf{x},$
+$\mathrm{grad}_{g_{\lambda}}f(x)=-(\mathbf{I}_{n}-\frac{\mathbf{B}_{\lambda}^{-1}\mathbf{B}\mathbf{x}\mathbf{x}^{\mathsf{T}}\mathbf{B}}{\mathbf{x}^{\mathsf{T}}\mathbf{B}\mathbf{B}_{\lambda}^{-1}\mathbf{B}\mathbf{x}})\mathbf{B}_{\lambda}^{-1}\mathbf{b}=-\mathbf{B}_{\lambda}^{-1}\mathbf{b}+\frac{\mathbf{x}^{\mathsf{T}}\mathbf{B}\mathbf{B}_{\lambda}^{-1}\mathbf{b}}{\mathbf{x}^{\mathsf{T}}\mathbf{B}\mathbf{B}_{\lambda}^{-1}\mathbf{B}\mathbf{x}}\mathbf{B}_{\lambda}^{-1}\mathbf{B}\mathbf{x},$
 by following Proposition 4.2.
-Subsequently, the update rule for RGD under the metric $g_{\lambda,\mathbf{x}}$ is $\mathbf{x}^{(t+1)}={\bar{\mathbf{x}}^{(t)}}/{\|\bar{\mathbf{x}}^{(t)}\|_{\mathbf{B}}}$ with $\bar{\mathbf{x}}^{(t)}=\mathbf{x}^{(t)}-s^{(t)}\mathrm{grad}_{g_{\lambda,\mathbf{x}}}f(\mathbf{x}^{(t)})$, where we adopt the polar retraction [26, (3.3)].
+Subsequently, the update rule for RGD under the metric $g_{\lambda}$ is $\mathbf{x}^{(t+1)}={\bar{\mathbf{x}}^{(t)}}/{\|\bar{\mathbf{x}}^{(t)}\|_{\mathbf{B}}}$ with $\bar{\mathbf{x}}^{(t)}=\mathbf{x}^{(t)}-s^{(t)}\mathrm{grad}_{g_{\lambda}}f(\mathbf{x}^{(t)})$, where we adopt the polar retraction [26, (3.3)].
 The Riemannian Hessian of $f$ at $\mathbf{x}^{*}$ along $\eta\in\operatorname{T}_{\mathbf{x}^{*}}\!\mathcal{M}_{\mathbf{B}}$ can be computed by
-$\operatorname{Hess}_{g_{\lambda,\mathbf{x}}}\!f(\mathbf{x}^{*})[\eta]$ $=\Pi_{g_{\lambda,\mathbf{x}},\mathbf{x}^{*}}(\mathrm{D}\mathrm{grad}_{g_{\lambda,\mathbf{x}}}f(\mathbf{x}^{*})[\eta])=\Pi_{g_{\lambda,\mathbf{x}},\mathbf{x}^{*}}(\|\mathbf{B}^{-1}\mathbf{b}\|_{\mathbf{B}}{\mathbf{B}_{\lambda}}^{-1}\mathbf{B}\eta),$
-since $\mathrm{grad}_{g_{\lambda,\mathbf{x}}}f(\mathbf{x}^{*})=\Pi_{g_{\lambda,\mathbf{x}},\mathbf{x}^{*}}(-\mathbf{B}_{\lambda}^{-1}\mathbf{b})=0$. The Rayleigh quotient Eq. 4 is given by
+$\operatorname{Hess}_{g_{\lambda}}\!f(\mathbf{x}^{*})[\eta]$ $=\Pi_{g_{\lambda},\mathbf{x}^{*}}(\mathrm{D}\mathrm{grad}_{g_{\lambda}}f(\mathbf{x}^{*})[\eta])=\Pi_{g_{\lambda},\mathbf{x}^{*}}(\|\mathbf{B}^{-1}\mathbf{b}\|_{\mathbf{B}}{\mathbf{B}_{\lambda}}^{-1}\mathbf{B}\eta),$
+since $\mathrm{grad}_{g_{\lambda}}f(\mathbf{x}^{*})=\Pi_{g_{\lambda},\mathbf{x}^{*}}(-\mathbf{B}_{\lambda}^{-1}\mathbf{b})=0$. The Rayleigh quotient Eq. 4 is given by
 $q(\eta)=\|\mathbf{B}^{-1}\mathbf{b}\|_{\mathbf{B}}\cdot\frac{\langle\eta,\mathbf{B}\eta\rangle}{\langle\eta,{\mathbf{B}_{\lambda}}\eta\rangle}\quad\text{for }\eta\in\operatorname{T}_{\mathbf{x}^{*}}\!\mathcal{M}_{\mathbf{B}}.$
-Consequently, we can compute the condition number of $\operatorname{Hess}_{g_{\lambda,\mathbf{x}}}\!f(\mathbf{x}^{*})$ in a same fashion as Proposition 4.4.
-Note that if $\lambda=0$, the Rayleigh quotient boils down to a constant $\|\mathbf{B}^{-1}\mathbf{b}\|_{\mathbf{B}}$ and thus $\kappa_{g_{0,\mathbf{x}}}(\operatorname{Hess}_{g_{0,\mathbf{x}}}\!f(\mathbf{x}^{*}))=1$.
+Consequently, we can compute the condition number of $\operatorname{Hess}_{g_{\lambda}}\!f(\mathbf{x}^{*})$ in a same fashion as Proposition 4.4.
+Note that if $\lambda=0$, the Rayleigh quotient boils down to a constant $\|\mathbf{B}^{-1}\mathbf{b}\|_{\mathbf{B}}$ and thus $\kappa_{g_{0}}(\operatorname{Hess}_{g_{0}}\!f(\mathbf{x}^{*}))=1$.
 Appendix B Proof of Proposition 2.2
@@
 By using Eq.
-4, $\mathbf{U}^{*}=\Sigma_{xx}^{-1/2}\bar{\mathbf{U}}$ and $\mathbf{V}^{*}=\Sigma_{yy}^{-1/2}\bar{\mathbf{V}}$, we can represent $\tilde{\eta}$ by
-(3) $\tilde{\eta}=(\Sigma_{xx}^{\frac{1}{2}}\eta_{1},\Sigma_{yy}^{\frac{1}{2}}\eta_{2})=(\bar{\mathbf{U}}\boldsymbol{\Omega}_{1}+\bar{\mathbf{U}}_{\perp}\mathbf{K}_{1},\bar{\mathbf{V}}\boldsymbol{\Omega}_{2}+\bar{\mathbf{V}}_{\perp}\mathbf{K}_{2}),$
-where $\bar{\mathbf{U}}_{\perp}\in\mathrm{St}(d_{x}-m,d_{x})$ and $\bar{\mathbf{V}}_{\perp}\in\mathrm{St}(d_{y}-m,d_{y})$ satisfy $\bar{\mathbf{U}}^{\mathsf{T}}\bar{\mathbf{U}}_{\perp}=0$, $\bar{\mathbf{V}}^{\mathsf{T}}\bar{\mathbf{V}}_{\perp}=0$ and $\mathbf{M}=[\bar{\mathbf{U}}\ \bar{\mathbf{U}}_{\perp}]\operatorname{diag}(\sigma_{1},\dots,\sigma_{r},0,\dots,0)[\bar{\mathbf{V}}\ \bar{\mathbf{V}}_{\perp}]^{\mathsf{T}}$ with $r=\mathrm{rank}(\mathbf{M})$.
+4, we can represent $\tilde{\eta}$ by
+$\tilde{\eta}$ $=(\Sigma_{xx}^{\frac{1}{2}}\eta_{1},\Sigma_{yy}^{\frac{1}{2}}\eta_{2})=(\Sigma_{xx}^{\frac{1}{2}}\mathbf{U}^{*}\boldsymbol{\Omega}_{1}+\Sigma_{xx}^{\frac{1}{2}}\mathbf{U}_{\Sigma_{xx}\perp}^{*}\mathbf{K}_{1},\Sigma_{yy}^{\frac{1}{2}}\mathbf{V}^{*}\boldsymbol{\Omega}_{2}+\Sigma_{yy}^{\frac{1}{2}}\mathbf{V}_{\Sigma_{yy}\perp}^{*}\mathbf{K}_{2})$
+(3) $=(\bar{\mathbf{U}}\boldsymbol{\Omega}_{1}+\bar{\mathbf{U}}_{\perp}\mathbf{K}_{1},\bar{\mathbf{V}}\boldsymbol{\Omega}_{2}+\bar{\mathbf{V}}_{\perp}\mathbf{K}_{2}),$
+where $\bar{\mathbf{U}}=\Sigma_{xx}^{\frac{1}{2}}\mathbf{U}^{*}\in\mathrm{St}(m,d_{x})$, $\bar{\mathbf{V}}=\Sigma_{yy}^{\frac{1}{2}}\mathbf{V}^{*}\in\mathrm{St}(m,d_{y})$, $\bar{\mathbf{U}}_{\perp}=\mathbf{U}_{\Sigma_{xx}\perp}^{*}\in\mathrm{St}(d_{x}-m,d_{x})$ and $\bar{\mathbf{V}}_{\perp}=\Sigma_{yy}^{\frac{1}{2}}\mathbf{V}_{\Sigma_{yy}\perp}^{*}\in\mathrm{St}(d_{y}-m,d_{y})$ satisfy $\bar{\mathbf{U}}^{\mathsf{T}}\bar{\mathbf{U}}_{\perp}=0$, $\bar{\mathbf{V}}^{\mathsf{T}}\bar{\mathbf{V}}_{\perp}=0$ and $\mathbf{M}=[\bar{\mathbf{U}}\ \bar{\mathbf{U}}_{\perp}]\operatorname{diag}(\sigma_{1},\dots,\sigma_{r},0,\dots,0)[\bar{\mathbf{V}}\ \bar{\mathbf{V}}_{\perp}]^{\mathsf{T}}$ with $r=\mathrm{rank}(\mathbf{M})$ by Eq.
+2.
 Taking Eq.
@@
 7 of the tangent space $\operatorname{T}_{\mathbf{U}}\!\mathcal{M}_{1}$ can be characterized by
-$(\operatorname{T}_{\mathbf{U}}\!\mathcal{M}_{1})^{\perp}=\{\mathbf{U}\mathbf{S}_{1}\mathbf{M}_{1,2}^{-1}:\mathbf{S}_{1}\in\mathbb{R}^{m\times m},\ \mathbf{S}_{1}=\mathbf{S}_{1}^{\mathsf{T}}\},$
+(1) $(\operatorname{T}_{\mathbf{U}}\!\mathcal{M}_{1})^{\perp}=\{\mathbf{U}\mathbf{S}_{1}\mathbf{M}_{1,2}^{-1}:\mathbf{S}_{1}\in\mathbb{R}^{m\times m},\ \mathbf{S}_{1}=\mathbf{S}_{1}^{\mathsf{T}}\},$
 since the dimension of $\{\mathbf{U}\mathbf{S}_{1}\mathbf{M}_{1,2}^{-1}:\mathbf{S}_{1}\in\mathbb{R}^{m\times m},\ \mathbf{S}_{1}=\mathbf{S}_{1}^{\mathsf{T}}\}$ is $m(m+1)/2$ and $\mathrm{tr}((\mathbf{U}\mathbf{S}_{1}\mathbf{M}_{1,2}^{-1})^{\mathsf{T}}\Sigma_{xx}(\mathbf{U}\boldsymbol{\Omega}_{1}+\mathbf{U}_{\Sigma_{xx}\perp}\mathbf{K}_{1})\mathbf{M}_{1,2})=0$ holds for all $\mathbf{S}_{1},\boldsymbol{\Omega}_{1},\mathbf{K}_{1}$ satisfying that $\mathbf{S}_{1}=\mathbf{S}_{1}^{\mathsf{T}}$ and $\boldsymbol{\Omega}_{1}=-\boldsymbol{\Omega}_{1}^{\mathsf{T}}$.